GNU Make

GNU Make

A Program for Directing Recompilation Edition 0.42, for make Version 3.66 Beta.

May 1993

by Richard M. Stallman and Roland McGrath

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988, ’89, ’90, ’91, ’92, ’93 Free Software Foundation, Inc.

Published by the Free Software Foundation 675 Massachusetts Avenue, Cambridge, MA 02139 USA Printed copies are available for $20 each.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the section entitled “GNU General Public License” is included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the text of the translation of the section entitled “GNU General Public License” must be approved for accuracy by the Foundation.

Cover art by Etienne Suvasa.

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc.

675 Mass Ave, Cambridge, MA 02139, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

1. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

2. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

3. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you changed the

files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part contains

or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program.

In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

4. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following:

a. Accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

5. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

6. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

7. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

8. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

9. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

10. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

11. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

12. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WAR-RANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFOR-MANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DE-FECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

13. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CON-SEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSI-BILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

Copyright (C) 19yy name of author
This program is free software; you can redistribute it and/or
modify it under the terms of the GNU General Public License
as published by the Free Software Foundation; either version 2
of the License, or (at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright
interest in the program `Gnomovision'
(which makes passes at compilers) written
by James Hacker.

signature of Ty Coon, 1 April 1989

Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

1 Overview of make

The make utility automatically determines which pieces of a large program need to be recompiled, and issues commands to recompile them. This manual describes GNU make, which was implemented by Richard Stallman and Roland McGrath. GNU make conforms to section 6.2 of IEEE Standard 1003.2-1992 (POSIX.2).

Our examples show C programs, since they are most common, but you can use make with any programming language whose compiler can be run with a shell command. Indeed, make is not limited to programs. You can use it to describe any task where some files must be updated automatically from others whenever the others change.

To prepare to use make, you must write a file called the makefile that describes the relationships among files in your program and provides commands for updating each file. In a program, typically, the executable file is updated from object files, which are in turn made by compiling source files.

Once a suitable makefile exists, each time you change some source files, this simple shell command:

     make

suffices to perform all necessary recompilations. The make program uses the makefile data base and the last-modification times of the files to decide which of the files need to be updated. For each of those files, it issues the commands recorded in the data base.

You can provide command line arguments to make to control which files should be recompiled, or how. See Chapter 9 [How to Run make], page 97.

1.1 How to Read This Manual

If you are new to make, or are looking for a general introduction, read the first few sections of each chapter, skipping the later sections. In each chapter, the first few sections contain introductory or general information and the later sections contain specialized or technical information. The exception is Chapter 2 [An Introduction to Makefiles], page 13, all of which is introductory.

If you are familiar with other make programs, see Chapter 12 [Features of GNU make], page 131, which lists the enhancements GNU make has, and Chapter 13 [Incompatibilities and Missing Features], page 135, which explains the few things GNU make lacks that others have.

For a quick summary, see Section 9.7 [Options Summary], page 103, Appendix A [Quick Reference], page 145, and Section 4.7 [Special Targets], page 40.

1.2 Problems and Bugs

If you have problems with GNU make or think you’ve found a bug, please report it to the developers; we cannot promise to do anything but we might well want to fix it.

Before reporting a bug, make sure you’ve actually found a real bug. Carefully reread the documentation and see if it really says you can do what you’re trying to do. If it’s not clear whether you should be able to do something or not, report that too; it’s a bug in the documentation!

Before reporting a bug or trying to fix it yourself, try to isolate it to the smallest possible makefile that reproduces the problem. Then send us the makefile and the exact results make gave you. Also say what you expected to occur; this will help us decide whether the problem was really in the documentation.

Once you’ve got a precise problem, please send electronic mail either through the Internet or via UUCP:

Internet address:

         bug-gnu-utils@prep.ai.mit.edu

UUCP path:

         mit-eddie!prep.ai.mit.edu!bug-gnu-utils

Please include the version number of make you are using. You can get this information with the command ‘make --version’. Be sure also to include the type of machine and operating system you are using. If possible, include the contents of the file ‘config.h’ that is generated by the configuration process.

Non-bug suggestions are always welcome as well. If you have questions about things that are unclear in the documentation or are just obscure features, contact Roland McGrath; he will try to help you out, although he may not have time to fix the problem.

You can send electronic mail to Roland McGrath either through the Internet or via UUCP:

Internet address:

    roland@prep.ai.mit.edu

UUCP path:

    mit-eddie!prep.ai.mit.edu!roland

12 GNU make

2 An Introduction to Makefiles

You need a file called a makefile to tell make what to do. Most often, the makefile tells make how to compile and link a program.

In this chapter, we will discuss a simple makefile that describes how to compile and link a text editor which consists of eight C source files and three header files. The makefile can also tell make how to run miscellaneous commands when explicitly asked (for example, to remove certain files as a clean-up operation). To see a more complex example of a makefile, see Appendix B [Complex Makefile], page 151.

When make recompiles the editor, each changed C source file must be recompiled. If a header file has changed, each C source file that includes the header file must be recompiled to be safe. Each compilation produces an object file corresponding to the source file. Finally, if any source file has been recompiled, all the object files, whether newly made or saved from previous compilations, must be linked together to produce the new executable editor.

2.1 What a Rule Looks Like

A simple makefile consists of “rules” with the following shape:

target …: dependencies … command … …

A target is usually the name of a file that is generated by a program; examples of targets are executable or object files. A target can also be the name of an action to carry out, such as ‘clean’ (see Section 4.4 [Phony Targets], page 37).

A dependency is a file that is used as input to create the target. A target often depends on several files.

A command is an action that make carries out. A rule may have more than one command, each on its own line. Please note: you need to put a tab character at the beginning of every command line! This is an obscurity that catches the unwary.

Usually a command is in a rule with dependencies and serves to create a target file if any of the dependencies change. However, the rule that specifies commands for the target need not have dependencies. For example, the rule containing the delete command associated with the target ‘clean’ does not have dependencies.

A rule, then, explains how and when to remake certain files which are the targets of the particular rule. make carries out the commands on the dependencies to create or update the target. A rule can also explain how and when to carry out an action. See Chapter 4 [Writing Rules], page 29.

A makefile may contain other text besides rules, but a simple makefile need only contain rules. Rules may look somewhat more complicated than shown in this template, but all fit the pattern more or less.

2.2 A Simple Makefile

Here is a straightforward makefile that describes the way an executable file called edit depends on eight object files which, in turn, depend on eight C source and three header files.

In this example, all the C files include ‘defs.h’, but only those defining editing commands include ‘command.h’, and only low level files that change the editor buffer include ‘buffer.h’.

edit : main.o kbd.o command.o display.o \
       insert.o search.o files.o utils.o
        cc -o edit main.o kbd.o command.o display.o \
                   insert.o search.o files.o utils.o
main.o : main.c defs.h
        cc -c main.c
kbd.o : kbd.c defs.h command.h
        cc -c kbd.c
command.o : command.c defs.h command.h
        cc -c command.c
display.o : display.c defs.h buffer.h
        cc -c display.c
insert.o : insert.c defs.h buffer.h
        cc -c insert.c
search.o : search.c defs.h buffer.h
        cc -c search.c
files.o : files.c defs.h buffer.h command.h
        cc -c files.c
utils.o : utils.c defs.h
        cc -c utils.c
clean :
        rm edit main.o kbd.o command.o display.o \
           insert.o search.o files.o utils.o

We split each long line into two lines using backslash-newline; this is like using one long line, but is easier to read.

To use this makefile to create the executable file called ‘edit’, type:

make

To use this makefile to delete the executable file and all the object files from the directory, type:

make clean

In the example makefile, the targets include the executable file ‘edit’, and the object files ‘main.o’ and ‘kbd.o’. The dependencies are files such as ‘main.c’ and ‘defs.h’. In fact, each ‘.o’ file is both a target and a dependency. Commands include ‘cc -c main.c’ and ‘cc -c kbd.c’.

When a target is a file, it needs to be recompiled or relinked if any of its dependencies change. In addition, any dependencies that are themselves automatically generated should be updated first. In this example, ‘edit’ depends on each of the eight object files; the object file ‘main.o’ depends on the source file ‘main.c’ and on the header file ‘defs.h’.

A shell command follows each line that contains a target and dependencies. These shell commands say how to update the target file. A tab character must come at the beginning of every command line to distinguish commands lines from other lines in the makefile. (Bear in mind that make does not know anything about how the commands work. It is up to you to supply commands that will update the target file properly. All make does is execute the commands in the rule you have specified when the target file needs to be updated.)

The target ‘clean’ is not a file, but merely the name of an action. Since you normally do not want to carry out the actions in this rule, ‘clean’ is not a dependency of any other rule. Consequently, make never does anything with it unless you tell it specifically. Note that this rule not only is not a dependency, it also does not have any dependencies, so the only purpose of the rule is to run the specified commands. Targets that do not refer to files but are just actions are called phony targets. See Section 4.4 [Phony Targets], page 37, for information about this kind of target. See Section 5.4 [Errors in Commands], page 54, to see how to cause make to ignore errors from rm or any other command.

2.3 How make Processes a Makefile

By default, make starts with the first rule (not counting rules whose target names start with ‘.’). This is called the default goal. (Goals are the targets that make strives ultimately to update. See Section 9.2 [Arguments to Specify the Goals], page 97.)

In the simple example of the previous section, the default goal is to update the executable program ‘edit’; therefore, we put that rule first.

Thus, when you give the command:

     make

make reads the makefile in the current directory and begins by processing the first rule. In the example, this rule is for relinking ‘edit’; but before make can fully process this rule, it must process the rules for the files that ‘edit’ depends on, which in this case are the object files. Each of these files is processed according to its own rule. These rules say to update each ‘.o’ file by compiling its source file. The recompilation must be done if the source file, or any of the header files named as dependencies, is more recent than the object file, or if the object file does not exist.

The other rules are processed because their targets appear as dependencies of the goal. If some other rule is not depended on by the goal (or anything it depends on, etc.), that rule is not processed, unless you tell make to do so (with a command such as make clean).

Before recompiling an object file, make considers updating its dependencies, the source file and header files. This makefile does not specify anything to be done for them—the ‘.c’ and ‘.h’ files are not the targets of any rules—so make does nothing for these files. But make would update automatically generated C programs, such as those made by Bison or Yacc, by their own rules at this time.

After recompiling whichever object files need it, make decides whether to relink ‘edit’. This must be done if the file ‘edit’ does not exist, or if any of the object files are newer than it. If an object file was just recompiled, it is now newer than ‘edit’, so ‘edit’ is relinked.

Thus, if we change the file ‘insert.c’ and run make, make will compile that file to update ‘insert.o’, and then link ‘edit’. If we change the file ‘command.h’ and run make, make will recompile the object files ‘kbd.o’, ‘command.o’ and ‘files.o’ and then link the file ‘edit’.

2.4 Variables Make Makefiles Simpler

In our example, we had to list all the object files twice in the rule for ‘edit’ (repeated here):

     edit : main.o kbd.o command.o display.o \
                   insert.o search.o files.o utils.o
             cc -o edit main.o kbd.o command.o display.o \
                        insert.o search.o files.o utils.o

Such duplication is error-prone; if a new object file is added to the system, we might add it to one list and forget the other. We can eliminate the risk and simplify the makefile by using a variable. Variables allow a text string to be defined once and substituted in multiple places later (see Chapter 6 [How to Use Variables], page 65).

It is standard practice for every makefile to have a variable named objects, OBJECTS, objs, OBJS, obj, or OBJ which is a list of all object file names. We would define such a variable objects with a line like this in the makefile:

objects = main.o kbd.o command.o display.o \
          insert.o search.o files.o utils.o

Then, each place we want to put a list of the object file names, we can substitute the variable’s value by writing ‘$(objects)’ (see Chapter 6 [How to Use Variables], page 65).

Here is how the complete simple makefile looks when you use a variable for the object files:

objects = main.o kbd.o command.o display.o \
          insert.o search.o files.o utils.o
edit : $(objects)
        cc -o edit $(objects)
main.o : main.c defs.h
        cc -c main.c
kbd.o : kbd.c defs.h command.h
        cc -c kbd.c
command.o : command.c defs.h command.h
        cc -c command.c
display.o : display.c defs.h buffer.h
        cc -c display.c
insert.o : insert.c defs.h buffer.h
        cc -c insert.c
search.o : search.c defs.h buffer.h
        cc -c search.c
files.o : files.c defs.h buffer.h command.h
        cc -c files.c
utils.o : utils.c defs.h
        cc -c utils.c
clean :
        rm edit $(objects)

2.5 Letting make Deduce the Commands

It is not necessary to spell out the commands for compiling the individual C source files, because make can figure them out: it has an implicit rule for updating a ‘.o’ file from a correspondingly named ‘.c’ file using a ‘cc -c’ command. For example, it will use the command ‘cc -c main.c -o main.o’ to compile ‘main.c’ into ‘main.o’. We can therefore omit the commands from the rules for the object files. See Chapter 10 [Using Implicit Rules], page 107.

When a ‘.c’ file is used automatically in this way, it is also automatically added to the list of dependencies. We can therefore omit the ‘.c’ files from the dependencies, provided we omit the commands.

Here is the entire example, with both of these changes, and a variable objects as suggested above:

objects = main.o kbd.o command.o display.o \
          insert.o search.o files.o utils.o
edit : $(objects)
        cc -o edit $(objects)
main.o : defs.h
kbd.o : defs.h command.h
command.o : defs.h command.h
display.o : defs.h buffer.h
insert.o : defs.h buffer.h
search.o : defs.h buffer.h
files.o : defs.h buffer.h command.h
utils.o : defs.h
.PHONY : clean
clean :
        -rm edit $(objects)

This is how we would write the makefile in actual practice. (The complications associated with ‘clean’ are described elsewhere. See Section 4.4 [Phony Targets], page 37, and Section 5.4 [Errors in Commands], page 54.)

Because implicit rules are so convenient, they are important. You will see them used frequently.

2.6 Another Style of Makefile

When the objects of a makefile are created only by implicit rules, an alternative style of makefile is possible. In this style of makefile, you group entries by their dependencies instead of by their targets. Here is what one looks like:

     objects = main.o kbd.o command.o display.o \
               insert.o search.o files.o utils.o
     edit : $(objects)
             cc -o edit $(objects)
     $(objects) : defs.h
     kbd.o command.o files.o : command.h
     display.o insert.o search.o files.o : buffer.h

Here ‘defs.h’ is given as a dependency of all the object files; ‘command.h’ and ‘buffer.h’ are dependencies of the specific object files listed for them.

Whether this is better is a matter of taste: it is more compact, but some people dislike it because they find it clearer to put all the information about each target in one place.

2.7 Rules for Cleaning the Directory

Compiling a program is not the only thing you might want to write rules for. Makefiles commonly tell how to do a few other things besides compiling a program: for example, how to delete all the object files and executables so that the directory is ‘clean’.

Here is how we could write a make rule for cleaning our example editor:

     clean:
             rm edit $(objects)

In practice, we might want to write the rule in a somewhat more complicated manner to handle unanticipated situations. We would do this:

     .PHONY : clean
     clean :
             -rm edit $(objects)

This prevents make from getting confused by an actual file called ‘clean’ and causes it to continue in spite of errors from rm. (See Section 4.4 [Phony Targets], page 37, and Section 5.4 [Errors in Commands], page 54.)

A rule such as this should not be placed at the beginning of the makefile, because we do not want it to run by default! Thus, in the example makefile, we want the rule for edit, which recompiles the editor, to remain the default goal.

Since clean is not a dependency of edit, this rule will not run at all if we give the command ‘make’ with no arguments. In order to make the rule run, we have to type ‘make clean’. See Chapter 9 [How to Run make], page 97.

22 GNU make

3 Writing Makefiles

The information that tells make how to recompile a system comes from reading a data base called the makefile.

3.1 What Makefiles Contain

Makefiles contain five kinds of things: explicit rules, implicit rules, variable definitions, directives, and comments. Rules, variables, and directives are described at length in later chapters.

• An explicit rule says when and how to remake one or more files, called the rule’s targets. It lists the other files that the targets depend on, and may also give commands to use to create or update the targets. See Chapter 4 [Writing Rules], page 29.

• An implicit rule says when and how to remake a class of files based on their names. It describes how a target may depend on a file with a name similar to the target and gives commands to create or update such a target. See Chapter 10 [Using Implicit Rules], page 107.

• A variable definition is a line that specifies a text string value for a variable that can be substituted into the text later. The simple makefile example shows a variable definition for objects as a list of all object files (see Section 2.4 [Variables Make Makefiles Simpler], page 17).

• A directive is a command for make to do something special while reading the makefile. These include:

• Reading another makefile (see Section 3.3 [Including Other Makefiles], page 24).

• Deciding (based on the values of variables) whether to use or ignore a part of the makefile (see Chapter 7 [Conditional Parts of Makefiles], page 79).

• Defining a variable from a verbatim string containing multiple lines (see Section 6.8 [Defining Variables Verbatim], page 77).

• ‘#’ in a line of a makefile starts a comment. It and the rest of the line are ignored, except that a trailing backslash not escaped by another backslash will continue the comment across multiple lines. Comments may appear on any of the lines in the makefile, except within a define directive, and perhaps within commands (where the shell decides what is a comment). A line containing just a comment (with perhaps spaces before it) is effectively blank, and is ignored.

3.2 What Name to Give Your Makefile

By default, when make looks for the makefile, it tries the following names, in order: ‘GNUmakefile’,

‘makefile’ and ‘Makefile’.

Normally you should call your makefile either ‘makefile’ or ‘Makefile’. (We recommend ‘Makefile’ because it appears prominently near the beginning of a directory listing, right near other important files such as ‘README’.) The first name checked, ‘GNUmakefile’, is not recommended for most makefiles. You should use this name if you have a makefile that is specific to GNU make, and will not be understood by other versions of make. Other make programs look for

‘makefile’ and ‘Makefile’, but not ‘GNUmakefile’.

If make finds none of these names, it does not use any makefile. Then you must specify a goal with a command argument, and make will attempt to figure out how to remake it using only its built-in implicit rules. See Chapter 10 [Using Implicit Rules], page 107.

If you want to use a nonstandard name for your makefile, you can specify the makefile name with the ‘-f’ or ‘--file’ option. The arguments ‘-f name’ or ‘--file name’ tell make to read the file name as the makefile. If you use more than one ‘-f’ or ‘--file’ option, you can specify several makefiles. All the makefiles are effectively concatenated in the order specified. The default makefile names ‘GNUmakefile’, ‘makefile’ and ‘Makefile’ are not checked automatically if you specify ‘-f’

or ‘--file’.

3.3 Including Other Makefiles

The include directive tells make to suspend reading the current makefile and read one or more other makefiles before continuing. The directive is a line in the makefile that looks like this:

include filenames…

filenames can contain shell file name patterns.

Extra spaces are allowed and ignored at the beginning of the line, but a tab is not allowed. (If the line begins with a tab, it will be considered a command line.) Whitespace is required between include and the file names, and between file names; extra whitespace is ignored there and at the end of the directive. A comment starting with ‘#’ is allowed at the end of the line. If the file names contain any variable or function references, they are expanded. See Chapter 6 [How to Use Variables], page 65.

For example, if you have three ‘.mk’ files, ‘a.mk’, ‘b.mk’, and ‘c.mk’, and $(bar) expands to bish bash, then the following expression

     include foo *.mk $(bar)

is equivalent to

     include foo a.mk b.mk c.mk bish bash

When make processes an include directive, it suspends reading of the containing makefile and reads from each listed file in turn. When that is finished, make resumes reading the makefile in which the directive appears.

One occasion for using include directives is when several programs, handled by individual makefiles in various directories, need to use a common set of variable definitions (see Section 6.5 [Setting Variables], page 73) or pattern rules (see Section 10.5 [Defining and Redefining Pattern Rules], page 115).

Another such occasion is when you want to generate dependencies from source files automatically; the dependencies can be put in a file that is included by the main makefile. This practice is generally cleaner than that of somehow appending the dependencies to the end of the main makefile as has been traditionally done with other versions of make. See Section 4.12 [Automatic Dependencies], page 47.

If the specified name does not start with a slash, and the file is not found in the current directory, several other directories are searched. First, any directories you have specified with the ‘-I’ or ‘--include-dir’ option are searched (see Section 9.7 [Summary of Options], page 103). Then the following directories (if they exist) are searched, in this order: ‘prefix/include’ (normally

‘/usr/local/include’) ‘/usr/gnu/include’, ‘/usr/local/include’, ‘/usr/include’.

If an included makefile cannot be found in any of these directories, a warning message is generated, but it is not an immediately fatal error; processing of the makefile containing the include continues. Once it has finished reading makefiles, make will try to remake any that are out of date or don’t exist. See Section 3.5 [How Makefiles Are Remade], page 26. Only after it has tried to find a way to remake a makefile and failed, will make diagnose the missing makefile as a fatal error.

3.4 The Variable MAKEFILES

If the environment variable MAKEFILES is defined, make considers its value as a list of names (separated by whitespace) of additional makefiles to be read before the others. This works much like the include directive: various directories are searched for those files (see Section 3.3 [Including Other Makefiles], page 24). In addition, the default goal is never taken from one of these makefiles and it is not an error if the files listed in MAKEFILES are not found.

The main use of MAKEFILES is in communication between recursive invocations of make (see Section 5.6 [Recursive Use of make], page 56). It usually is not desirable to set the environment variable before a top-level invocation of make, because it is usually better not to mess with a makefile from outside. However, if you are running make without a specific makefile, a makefile in MAKEFILES can do useful things to help the built-in implicit rules work better, such as defining search paths (see Section 4.3 [Directory Search], page 33).

Some users are tempted to set MAKEFILES in the environment automatically on login, and program makefiles to expect this to be done. This is a very bad idea, because such makefiles will fail to work if run by anyone else. It is much better to write explicit include directives in the makefiles. See Section 3.3 [Including Other Makefiles], page 24.

3.5 How Makefiles Are Remade

Sometimes makefiles can be remade from other files, such as RCS or SCCS files. If a makefile can be remade from other files, you probably want make to get an up-to-date version of the makefile to read in.

To this end, after reading in all makefiles, make will consider each as a goal target and attempt to update it. If a makefile has a rule which says how to update it (found either in that very makefile or in another one) or if an implicit rule applies to it (see Chapter 10 [Using Implicit Rules], page 107), it will be updated if necessary. After all makefiles have been checked, if any have actually been changed, make starts with a clean slate and reads all the makefiles over again. (It will also attempt to update each of them over again, but normally this will not change them again, since they are already up to date.)

If the makefiles specify a double-colon rule to remake a file with commands but no dependencies, that file will always be remade (see Section 4.11 [Double-Colon], page 46). In the case of makefiles, a makefile that has a double-colon rule with commands but no dependencies will be remade every time make is run, and then again after make starts over and reads the makefiles in again. This would cause an infinite loop: make would constantly remake the makefile, and never do anything else. So, to avoid this, make will not attempt to remake makefiles which are specified as double-colon targets but have no dependencies.

If you do not specify any makefiles to be read with ‘-f’ or ‘--file’ options, make will try the default makefile names; see Section 3.2 [What Name to Give Your Makefile], page 24. Unlike makefiles explicitly requested with ‘-f’ or ‘--file’ options, make is not certain that these makefiles should exist. However, if a default makefile does not exist but can be created by running make rules, you probably want the rules to be run so that the makefile can be used.

Therefore, if none of the default makefiles exists, make will try to make each of them in the same order in which they are searched for (see Section 3.2 [What Name to Give Your Makefile], page 24) until it succeeds in making one, or it runs out of names to try. Note that it is not an error if make cannot find or make any makefile; a makefile is not always necessary.

When you use the ‘-t’ or ‘--touch’ option (see Section 9.3 [Instead of Executing the Commands], page 99), you would not want to use an out-of-date makefile to decide which targets to touch. So the ‘-t’ option has no effect on updating makefiles; they are really updated even if ‘-t’ is specified. Likewise, ‘-q’ (or ‘--question’) and ‘-n’ (or ‘--just-print’) do not prevent updating of makefiles, because an out-of-date makefile would result in the wrong output for other targets. Thus, ‘make -f mfile -n foo’ will update ‘mfile’, read it in, and then print the commands to update ‘foo’ and its dependencies without running them. The commands printed for ‘foo’ will be those specified in the updated contents of ‘mfile’.

However, on occasion you might actually wish to prevent updating of even the makefiles. You can do this by specifying the makefiles as goals in the command line as well as specifying them as makefiles. When the makefile name is specified explicitly as a goal, the options ‘-t’ and so on do apply to them.

Thus, ‘make -f mfile -n mfile foo’ would read the makefile ‘mfile’, print the commands needed to update it without actually running them, and then print the commands needed to update ‘foo’ without running them. The commands for ‘foo’ will be those specified by the existing contents of ‘mfile’.

3.6 Overriding Part of Another Makefile

Sometimes it is useful to have a makefile that is mostly just like another makefile. You can often use the ‘include’ directive to include one in the other, and add more targets or variable definitions.

However, if the two makefiles give different commands for the same target, make will not let you just do this. But there is another way.

In the containing makefile (the one that wants to include the other), you can use the .DEFAULT special target to say that to remake any target that cannot be made from the information in the containing makefile, make should look in another makefile. See Section 10.6 [Defining Last-Resort Default Rules], page 122, for more information on .DEFAULT.

For example, if you have a makefile called ‘Makefile’ that says how to make the target ‘foo’ (and other targets), you can write a makefile called ‘GNUmakefile’ that contains:

     foo:
             frobnicate > foo
     .DEFAULT:
             @$(MAKE) -f Makefile $@

If you say ‘make foo’, make will find ‘GNUmakefile’, read it, and see that to make ‘foo’, it needs to run the command ‘frobnicate > foo’. If you say ‘make bar’, make will find no way to make

‘bar’ in ‘GNUmakefile’, so it will use the commands from .DEFAULT: ‘make -f Makefile bar’. If

‘Makefile’ provides a rule for updating ‘bar’, make will apply the rule. And likewise for any other target that ‘GNUmakefile’ does not say how to make.

4 Writing Rules

A rule appears in the makefile and says when and how to remake certain files, called the rule’s targets (most often only one per rule). It lists the other files that are the dependencies of the target, and commands to use to create or update the target.

The order of rules is not significant, except for determining the default goal: the target for make to consider, if you do not otherwise specify one. The default goal is the target of the first rule in the first makefile. If the first rule has multiple targets, only the first target is taken as the default. There are two exceptions: a target starting with a period is not a default unless it contains one or more slashes, ‘/’, as well; and, a target that defines a pattern rule has no effect on the default goal. (See Section 10.5 [Defining and Redefining Pattern Rules], page 115.)

Therefore, we usually write the makefile so that the first rule is the one for compiling the entire program or all the programs described by the makefile (often with a target called ‘all’). See Section 9.2 [Arguments to Specify the Goals], page 97.

4.1 Rule Syntax

In general, a rule looks like this:

targets : dependencies command …

or like this:

targets : dependencies ; command command …

The targets are file names, separated by spaces. Wildcard characters may be used (see Section 4.2 [Using Wildcard Characters in File Names], page 30) and a name of the form ‘a(m)’ represents member m in archive file a (see Section 11.1 [Archive Members as Targets], page 127). Usually there is only one target per rule, but occasionally there is a reason to have more (see Section 4.8 [Multiple Targets in a Rule], page 42).

The command lines start with a tab character. The first command may appear on the line after the dependencies, with a tab character, or may appear on the same line, with a semicolon. Either way, the effect is the same. See Chapter 5 [Writing the Commands in Rules], page 51.

Because dollar signs are used to start variable references, if you really want a dollar sign in a rule you must write two of them, ‘$$’ (see Chapter 6 [How to Use Variables], page 65). You may split a long line by inserting a backslash followed by a newline, but this is not required, as make places no limit on the length of a line in a makefile.

A rule tells make two things: when the targets are out of date, and how to update them when necessary.

The criterion for being out of date is specified in terms of the dependencies, which consist of file names separated by spaces. (Wildcards and archive members (see Chapter 11 [Archives], page 127) are allowed here too.) A target is out of date if it does not exist or if it is older than any of the dependencies (by comparison of last-modification times). The idea is that the contents of the target file are computed based on information in the dependencies, so if any of the dependencies changes, the contents of the existing target file are no longer necessarily valid.

How to update is specified by commands. These are lines to be executed by the shell (normally ‘sh’), but with some extra features (see Chapter 5 [Writing the Commands in Rules], page 51).

4.2 Using Wildcard Characters in File Names

A single file name can specify many files using wildcard characters. The wildcard characters in make are ‘*’, ‘?’ and ‘[…]’, the same as in the Bourne shell. For example, ‘*.c’ specifies a list of all the files (in the working directory) whose names end in ‘.c’.

The character ‘~’ at the beginning of a file name also has special significance. If alone, or followed by a slash, it represents your home directory. For example ‘~/bin’ expands to ‘/home/you/bin’. If the ‘~’ is followed by a word, the string represents the home directory of the user named by that word. For example ‘~john/bin’ expands to ‘/home/john/bin’.

Wildcard expansion happens automatically in targets, in dependencies, and in commands (where the shell does the expansion). In other contexts, wildcard expansion happens only if you request it explicitly with the wildcard function.

The special significance of a wildcard character can be turned off by preceding it with a backslash. Thus, ‘foo\*bar’ would refer to a specific file whose name consists of ‘foo’, an asterisk, and ‘bar’.

4.2.1 Wildcard Examples

Wildcards can be used in the commands of a rule, where they are expanded by the shell. For example, here is a rule to delete all the object files:

     clean:  rm -f *.o

Wildcards are also useful in the dependencies of a rule. With the following rule in the makefile, ‘make print’ will print all the ‘.c’ files that have changed since the last time you printed them:

     print: *.c
             lpr -p $?
             touch print

This rule uses ‘print’ as an empty target file; see Section 4.6 [Empty Target Files to Record Events], page 40. (The automatic variable ‘$?’ is used to print only those files that have changed; see Section 10.5.3 [Automatic Variables], page 118.)

Wildcard expansion does not happen when you define a variable. Thus, if you write this:

     objects = *.o

then the value of the variable objects is the actual string ‘*.o’. However, if you use the value of objects in a target, dependency or command, wildcard expansion will take place at that time. To set objects to the expansion, instead use:

     objects := $(wildcard *.o)

See Section 4.2.3 [Wildcard Function], page 32.

4.2.2 Pitfalls of Using Wildcards

Now here is an example of a naive way of using wildcard expansion, that does not do what you would intend. Suppose you would like to say that the executable file ‘foo’ is made from all the object files in the directory, and you write this:

     objects = *.o
     foo : $(objects)
             cc -o foo $(CFLAGS) $(objects)

The value of objects is the actual string ‘*.o’. Wildcard expansion happens in the rule for ‘foo’, so that each existing ‘.o’ file becomes a dependency of ‘foo’ and will be recompiled if necessary.

But what if you delete all the ‘.o’ files? Then ‘*.o’ will expand into nothing. The target ‘foo’ will have no dependencies and would be remade by linking no object files. This is not what you want!

Actually it is possible to obtain the desired result with wildcard expansion, but you need more sophisticated techniques, including the wildcard function and string substitution. These are described in the following section.

4.2.3 The Function wildcard

Wildcard expansion happens automatically in rules. But wildcard expansion does not normally take place when a variable is set, or inside the arguments of a function. If you want to do wildcard expansion in such places, you need to use the wildcard function, like this:

     $(wildcard pattern)

This string, used anywhere in a makefile, is replaced by a space-separated list of names of existing files that match the pattern pattern.

One use of the wildcard function is to get a list of all the C source files in a directory, like this:

     $(wildcard *.c)

We can change the list of C source files into a list of object files by replacing the ‘.o’ suffix with ‘.c’ in the result, like this:

     $(patsubst %.c,%.o,$(wildcard *.c))

(Here we have used another function, patsubst. See Section 8.2 [Functions for String Substitution and Analysis], page 86.)

Thus, a makefile to compile all C source files in the directory and then link them together could be written as follows:

     objects := $(patsubst %.c,%.o,$(wildcard *.c))
     foo : $(objects)
             cc -o foo $(objects)

(This takes advantage of the implicit rule for compiling C programs, so there is no need to write explicit rules for compiling the files. See Section 6.2 [The Two Flavors of Variables], page 66, for an explanation of ‘:=’, which is a variant of ‘=’.)

4.3 Searching Directories for Dependencies

For large systems, it is often desirable to put sources in a separate directory from the binaries. The directory search features of make facilitate this by searching several directories automatically to find a dependency. When you redistribute the files among directories, you do not need to change the individual rules, just the search paths.

4.3.1 VPATH: Search Path for All Dependencies

The value of the make variable VPATH specifies a list of directories that make should search. Most often, the directories are expected to contain dependency files that are not in the current directory; however, VPATH specifies a search list that make applies for all files, including files which are targets of rules.

Thus, if a file that is listed as a target or dependency does not exist in the current directory, make searches the directories listed in VPATH for a file with that name. If a file is found in one of them, that file becomes the dependency. Rules may then specify the names of source files in the dependencies as if they all existed in the current directory. See Section 4.3.3 [Writing Shell Commands with Directory Search], page 36.

In the VPATH variable, directory names are separated by colons. The order in which directories are listed is the order followed by make in its search.

For example,

     VPATH = src:../headers

specifies a path containing two directories, ‘src’ and ‘../headers’, which make searches in that order.

With this value of VPATH, the following rule,

     foo.o : foo.c

is interpreted as if it were written like this:

     foo.o : src/foo.c

assuming the file ‘foo.c’ does not exist in the current directory but is found in the directory ‘src’.

4.3.2 The vpath Directive

Similar to the VPATH variable but more selective is the vpath directive (note lower case), which allows you to specify a search path for a particular class of file names, those that match a particular pattern. Thus you can supply certain search directories for one class of file names and other directories (or none) for other file names.

There are three forms of the vpath directive: vpath pattern directories

Specify the search path directories for file names that match pattern.

The search path, directories, is a colon-separated list of directories to be searched, just like the search path used in the VPATH variable. vpath pattern

Clear out the search path associated with pattern.

vpath

Clear all search paths previously specified with vpath directives.

A vpath pattern is a string containing a ‘%’ character. The string must match the file name of a dependency that is being searched for, the ‘%’ character matching any sequence of zero or more characters (as in pattern rules; see Section 10.5 [Defining and Redefining Pattern Rules], page 115). For example, %.h matches files that end in .h. (If there is no ‘%’, the pattern must match the dependency exactly, which is not useful very often.)

‘%’ characters in a vpath directive’s pattern can be quoted with preceding backslashes (‘\’). Backslashes that would otherwise quote ‘%’ characters can be quoted with more backslashes. Backslashes that quote ‘%’ characters or other backslashes are removed from the pattern before it is compared to file names. Backslashes that are not in danger of quoting ‘%’ characters go unmolested.

When a dependency fails to exist in the current directory, if the pattern in a vpath directive matches the name of the dependency file, then the directories in that directive are searched just like (and before) the directories in the VPATH variable.

For example,

     vpath %.h ../headers

tells make to look for any dependency whose name ends in ‘.h’ in the directory ‘../headers’ if the file is not found in the current directory.

If several vpath patterns match the dependency file’s name, then make processes each matching vpath directive one by one, searching all the directories mentioned in each directive. make handles multiple vpath directives in the order in which they appear in the makefile; multiple directives with the same pattern are independent of each other.

Thus,

     vpath %.c foo
     vpath %   blish
     vpath %.c bar

will look for a file ending in ‘.c’ in ‘foo’, then ‘blish’, then ‘bar’, while

     vpath %.c foo:bar
     vpath %   blish

will look for a file ending in ‘.c’ in ‘foo’, then ‘bar’, then ‘blish’.

4.3.3 Writing Shell Commands with Directory Search

When a dependency is found in another directory through directory search, this cannot change the commands of the rule; they will execute as written. Therefore, you must write the commands with care so that they will look for the dependency in the directory where make finds it.

This is done with the automatic variables such as ‘$^’ (see Section 10.5.3 [Automatic Variables], page 118). For instance, the value of ‘$^’ is a list of all the dependencies of the rule, including the names of the directories in which they were found, and the value of ‘$@’ is the target. Thus:

     foo.o : foo.c
             cc -c $(CFLAGS) $^-o $@

(The variable CFLAGS exists so you can specify flags for C compilation by implicit rules; we use it here for consistency so it will affect all C compilations uniformly; see Section 10.3 [Variables Used by Implicit Rules], page 112.)

Often the dependencies include header files as well, which you do not want to mention in the commands. The automatic variable ‘$<’ is just the first dependency:

     VPATH = src:../headers
     foo.o : foo.c defs.h hack.h
             cc -c $(CFLAGS) $< -o $@

4.3.4 Directory Search and Implicit Rules

The search through the directories specified in VPATH or with vpath also happens during consideration of implicit rules (see Chapter 10 [Using Implicit Rules], page 107).

For example, when a file ‘foo.o’ has no explicit rule, make considers implicit rules such as to compile ‘foo.c’ if that file exists. If such a file is lacking in the current directory, the appropriate directories are searched for it. If ‘foo.c’ exists (or is mentioned in the makefile) in any of the directories, the implicit rule for C compilation is applied.

The commands of implicit rules normally use automatic variables as a matter of necessity; consequently they will use the file names found by directory search with no extra effort.

4.3.5 Directory Search for Link Libraries

Directory search applies in a special way to libraries used with the linker. This special feature comes into play when you write a dependency whose name is of the form ‘-lname’. (You can tell something strange is going on here because the dependency is normally the name of a file, and the file name of the library looks like ‘libname.a’, not like ‘-lname’.)

When a dependency’s name has the form ‘-lname’, make handles it specially by searching for the file ‘libname.a’ in the current directory, in directories specified by matching vpath search paths and the VPATH search path, and then in the directories ‘/lib’, ‘/usr/lib’, and ‘prefix/lib’

(normally ‘/usr/local/lib’).

For example,

     foo : foo.c -lcurses
             cc $^-o $@

would cause the command ‘cc foo.c /usr/lib/libcurses.a -o foo’ to be executed when ‘foo’

is older than ‘foo.c’ or than ‘/usr/lib/libcurses.a’.

4.4 Phony Targets

A phony target is one that is not really the name of a file. It is just a name for some commands to be executed when you make an explicit request. There are two reasons to use a phony target: to avoid a conflict with a file of the same name, and to improve performance.

If you write a rule whose commands will not create the target file, the commands will be executed every time the target comes up for remaking. Here is an example:

     clean:  rm *.o temp

Because the rm command does not create a file named ‘clean’, probably no such file will ever exist. Therefore, the rm command will be executed every time you say ‘make clean’.

The phony target will cease to work if anything ever does create a file named ‘clean’ in this directory. Since it has no dependencies, the file ‘clean’ would inevitably be considered up to date, and its commands would not be executed. To avoid this problem, you can explicitly declare the target to be phony, using the special target .PHONY (see Section 4.7 [Special Built-in Target Names], page 40) as follows:

     .PHONY : clean

Once this is done, ‘make clean’ will run the commands regardless of whether there is a file named

‘clean’.

Since it knows that phony targets do not name actual files that could be remade from other files, make skips the implicit rule search for phony targets (see Chapter 10 [Implicit Rules], page 107). This is why declaring a target phony is good for performance, even if you are not worried about the actual file existing.

Thus, you first write the line that states that clean is a phony target, then you write the rule, like this:

     .PHONY: clean
     clean:
             rm *.o temp

A phony target should not be a dependency of a real target file; if it is, its commands are run every time make goes to update that file. As long as a phony target is never a dependency of a real target, the phony target commands will be executed only when the phony target is a specified goal (see Section 9.2 [Arguments to Specify the Goals], page 97).

Phony targets can have dependencies. When one directory contains multiple programs, it is most convenient to describe all of the programs in one makefile ‘./Makefile’. Since the target remade by default will be the first one in the makefile, it is common to make this a phony target named ‘all’ and give it, as dependencies, all the individual programs. For example:

all : prog1 prog2 prog3
.PHONY : all
prog1 : prog1.o utils.o
        cc -o prog1 prog1.o utils.o
prog2 : prog2.o
        cc -o prog2 prog2.o
prog3 : prog3.o sort.o utils.o
        cc -o prog3 prog3.o sort.o utils.o

Now you can say just ‘make’ to remake all three programs, or specify as arguments the ones to remake (as in ‘make prog1 prog3’).

When one phony target is a dependency of another, it serves as a subroutine of the other. For example, here ‘make cleanall’ will delete the object files, the difference files, and the file ‘program’:

.PHONY: cleanall cleanobj cleandiff
cleanall : cleanobj cleandiff
        rm program
cleanobj :
        rm *.o
cleandiff :
        rm *.diff

4.5 Rules without Commands or Dependencies

If a rule has no dependencies or commands, and the target of the rule is a nonexistent file, then make imagines this target to have been updated whenever its rule is run. This implies that all targets depending on this one will always have their commands run.

An example will illustrate this:

clean: FORCE
        rm $(objects)
FORCE:

Here the target ‘FORCE’ satisfies the special conditions, so the target ‘clean’ that depends on it is forced to run its commands. There is nothing special about the name ‘FORCE’, but that is one name commonly used this way.

As you can see, using ‘FORCE’ this way has the same results as using ‘.PHONY: clean’.

Using ‘.PHONY’ is more explicit and more efficient. However, other versions of make do not support ‘.PHONY’; thus ‘FORCE’ appears in many makefiles. See Section 4.4 [Phony Targets], page 37.

4.6 Empty Target Files to Record Events

The empty target is a variant of the phony target; it is used to hold commands for an action that you request explicitly from time to time. Unlike a phony target, this target file can really exist; but the file’s contents do not matter, and usually are empty.

The purpose of the empty target file is to record, with its last-modification time, when the rule’s commands were last executed. It does so because one of the commands is a touch command to update the target file.

The empty target file must have some dependencies. When you ask to remake the empty target, the commands are executed if any dependency is more recent than the target; in other words, if a dependency has changed since the last time you remade the target. Here is an example:

     print: foo.c bar.c
             lpr -p $?
             touch print

With this rule, ‘make print’ will execute the lpr command if either source file has changed since the last ‘make print’. The automatic variable ‘$?’ is used to print only those files that have changed (see Section 10.5.3 [Automatic Variables], page 118).

4.7 Special Built-in Target Names

Certain names have special meanings if they appear as targets.

.PHONY

The dependencies of the special target .PHONY are considered to be phony targets. When it is time to consider such a target, make will run its commands unconditionally, regardless of whether a file with that name exists or what its last-modification time is. See Section 4.4 [Phony Targets], page 37.

.SUFFIXES

The dependencies of the special target .SUFFIXES are the list of suffixes to be used in checking for suffix rules. See Section 10.7 [Old-Fashioned Suffix Rules], page 123.

.DEFAULT

The commands specified for .DEFAULT are used for any target for which no rules are found (either explicit rules or implicit rules). See Section 10.6 [Last Resort], page 122. If .DEFAULT commands are specified, every file mentioned as a dependency, but not as a target in a rule, will have these commands executed on its behalf. See Section 10.8 [Implicit Rule Search Algorithm], page 125.

.PRECIOUS

The targets which .PRECIOUS depends on are given the following special treatment: if make is killed or interrupted during the execution of their commands, the target is not deleted. See Section 5.5 [Interrupting or Killing make], page 55. Also, if the target is an intermediate file, it will not be deleted after it is no longer needed, as is normally done. See Section 10.4 [Chains of Implicit Rules], page 114.

You can also list the target pattern of an implicit rule (such as ‘%.o’) as a dependency file of the special target .PRECIOUS to preserve intermediate files whose target patterns match that file’s name.

.IGNORE

Simply by being mentioned as a target, .IGNORE says to ignore errors in execution of commands. The dependencies and commands for .IGNORE are not meaningful.

‘.IGNORE’ exists for historical compatibility. Since .IGNORE affects every command in the makefile, it is not very useful; we recommend you use the more selective ways to ignore errors in specific commands. See Section 5.4 [Errors in Commands], page 54.

.SILENT

Simply by being mentioned as a target, .SILENT says not to print commands before executing them. The dependencies and commands for .SILENT are not meaningful.

‘.SILENT’ exists for historical compatibility. We recommend you use the more selective ways to silence specific commands. See Section 5.1 [Command Echoing], page 51. If you want to silence all commands for a particular run of make, use the ‘-s’ or ‘--silent’ option (see Section 9.7 [Options Summary], page 103).

.EXPORT_ALL_VARIABLES

Simply by being mentioned as a target, this tells make to export all variables to child processes by default. See Section 5.6.2 [Communicating Variables to a Sub-make], page 57.

Any defined implicit rule suffix also counts as a special target if it appears as a target, and so does the concatenation of two suffixes, such as ‘.c.o’. These targets are suffix rules, an obsolete way of defining implicit rules (but a way still widely used). In principle, any target name could be special in this way if you break it in two and add both pieces to the suffix list. In practice, suffixes normally begin with ‘.’, so these special target names also begin with ‘.’. See Section 10.7 [Old-Fashioned Suffix Rules], page 123.

4.8 Multiple Targets in a Rule

A rule with multiple targets is equivalent to writing many rules, each with one target, and all identical aside from that. The same commands apply to all the targets, but their effects may vary because you can substitute the actual target name into the command using ‘$@’. The rule contributes the same dependencies to all the targets also.

This is useful in two cases.

• You want just dependencies, no commands. For example:

         kbd.o command.o files.o: command.h

gives an additional dependency to each of the three object files mentioned.

• Similar commands work for all the targets. The commands do not need to be absolutely identical, since the automatic variable ‘$@’ can be used to substitute the particular target to be remade into the commands (see Section 10.5.3 [Automatic Variables], page 118). For example:

         bigoutput littleoutput : text.g
                 generate text.g -$(subst output,,$@) > $@

is equivalent to

         bigoutput : text.g
                 generate text.g -big > bigoutput
         littleoutput : text.g
                 generate text.g -little > littleoutput

Here we assume the hypothetical program generate makes two types of output, one if given ‘-big’ and one if given ‘-little’. See Section 8.2 [Functions for String Substitution and Analysis], page 86, for an explanation of the subst function.

Suppose you would like to vary the dependencies according to the target, much as the variable ‘$@’ allows you to vary the commands. You cannot do this with multiple targets in an ordinary rule, but you can do it with a static pattern rule. See Section 4.10 [Static Pattern Rules], page 44.

4.9 Multiple Rules for One Target

One file can be the target of several rules. All the dependencies mentioned in all the rules are merged into one list of dependencies for the target. If the target is older than any dependency from any rule, the commands are executed.

There can only be one set of commands to be executed for a file. If more than one rule gives commands for the same file, make uses the last set given and prints an error message. (As a special case, if the file’s name begins with a dot, no error message is printed. This odd behavior is only for compatibility with other implementations of make.) There is no reason to write your makefiles this way; that is why make gives you an error message.

An extra rule with just dependencies can be used to give a few extra dependencies to many files at once. For example, one usually has a variable named objects containing a list of all the compiler output files in the system being made. An easy way to say that all of them must be recompiled if ‘config.h’ changes is to write the following:

     objects = foo.o bar.o
     foo.o : defs.h
     bar.o : defs.h test.h
     $(objects) : config.h

This could be inserted or taken out without changing the rules that really specify how to make the object files, making it a convenient form to use if you wish to add the additional dependency intermittently.

Another wrinkle is that the additional dependencies could be specified with a variable that you set with a command argument to make (see Section 9.5 [Overriding Variables], page 101). For example,

     extradeps=
     $(objects) : $(extradeps)

means that the command ‘make extradeps=foo.h’ will consider ‘foo.h’ as a dependency of each object file, but plain ‘make’ will not.

If none of the explicit rules for a target has commands, then make searches for an applicable implicit rule to find some commands see Chapter 10 [Using Implicit Rules], page 107).

4.10 Static Pattern Rules

Static pattern rules are rules which specify multiple targets and construct the dependency names for each target based on the target name. They are more general than ordinary rules with multiple targets because the targets do not have to have identical dependencies. Their dependencies must be analogous, but not necessarily identical.

4.10.1 Syntax of Static Pattern Rules

Here is the syntax of a static pattern rule:

targets …: target-pattern: dep-patterns … commands …

The targets list specifies the targets that the rule applies to. The targets can contain wildcard characters, just like the targets of ordinary rules (see Section 4.2 [Using Wildcard Characters in File Names], page 30).

The target-pattern and dep-patterns say how to compute the dependencies of each target. Each target is matched against the target-pattern to extract a part of the target name, called the stem. This stem is substituted into each of the dep-patterns to make the dependency names (one from each dep-pattern).

Each pattern normally contains the character ‘%’ just once. When the target-pattern matches a target, the ‘%’ can match any part of the target name; this part is called the stem. The rest of the pattern must match exactly. For example, the target ‘foo.o’ matches the pattern ‘%.o’, with ‘foo’ as the stem. The targets ‘foo.c’ and ‘foo.out’ do not match that pattern.

The dependency names for each target are made by substituting the stem for the ‘%’ in each dependency pattern. For example, if one dependency pattern is ‘%.c’, then substitution of the stem ‘foo’ gives the dependency name ‘foo.c’. It is legitimate to write a dependency pattern that does not contain ‘%’; then this dependency is the same for all targets.

‘%’ characters in pattern rules can be quoted with preceding backslashes (‘\’). Backslashes that would otherwise quote ‘%’ characters can be quoted with more backslashes. Backslashes that quote ‘%’ characters or other backslashes are removed from the pattern before it is compared to file names or has a stem substituted into it. Backslashes that are not in danger of quoting ‘%’ characters go unmolested. For example, the pattern ‘the\%weird\\%pattern\\’ has ‘the%weird\’ preceding the operative ‘%’ character, and ‘pattern\\’ following it. The final two backslashes are left alone because they cannot affect any ‘%’ character.

Here is an example, which compiles each of ‘foo.o’ and ‘bar.o’ from the corresponding ‘.c’ file:

     objects = foo.o bar.o
     $(objects): %.o: %.c
             $(CC) -c $(CFLAGS) $< -o $@

Here ‘$<’ is the automatic variable that holds the name of the dependency and ‘$@’ is the automatic variable that holds the name of the target; see Section 10.5.3 [Automatic Variables], page 118.

Each target specified must match the target pattern; a warning is issued for each target that does not. If you have a list of files, only some of which will match the pattern, you can use the filter function to remove nonmatching file names (see Section 8.2 [Functions for String Substitution and Analysis], page 86):

     files = foo.elc bar.o lose.o
     $(filter %.o,$(files)): %.o: %.c
             $(CC) -c $(CFLAGS) $< -o $@
     $(filter %.elc,$(files)): %.elc: %.el
             emacs -f batch-byte-compile $<

Here the result of ‘$(filter %.o,$(files))’ is ‘bar.o lose.o’, and the first static pattern rule causes each of these object files to be updated by compiling the corresponding C source file. The result of ‘$(filter %.elc,$(files))’ is ‘foo.elc’, so that file is made from ‘foo.el’.

Another example shows how to use $* in static pattern rules:

     bigoutput littleoutput : %output : text.g
             generate text.g -$* > $@

When the generate command is run, $* will expand to the stem, either ‘big’ or ‘little’.

4.10.2 Static Pattern Rules versus Implicit Rules

A static pattern rule has much in common with an implicit rule defined as a pattern rule (see Section 10.5 [Defining and Redefining Pattern Rules], page 115). Both have a pattern for the target and patterns for constructing the names of dependencies. The difference is in how make decides when the rule applies.

An implicit rule can apply to any target that matches its pattern, but it does apply only when the target has no commands otherwise specified, and only when the dependencies can be found. If more than one implicit rule appears applicable, only one applies; the choice depends on the order of rules.

By contrast, a static pattern rule applies to the precise list of targets that you specify in the rule. It cannot apply to any other target and it invariably does apply to each of the targets specified. If two conflicting rules apply, and both have commands, that’s an error.

The static pattern rule can be better than an implicit rule for these reasons:

• You may wish to override the usual implicit rule for a few files whose names cannot be categorized syntactically but can be given in an explicit list.

• If you cannot be sure of the precise contents of the directories you are using, you may not be sure which other irrelevant files might lead make to use the wrong implicit rule. The choice might depend on the order in which the implicit rule search is done. With static pattern rules, there is no uncertainty: each rule applies to precisely the targets specified.

4.11 Double-Colon Rules

Double-colon rules are rules written with ‘::’ instead of ‘:’ after the target names. They are handled differently from ordinary rules when the same target appears in more than one rule.

When a target appears in multiple rules, all the rules must be the same type: all ordinary, or all double-colon. If they are double-colon, each of them is independent of the others. Each doublecolon rule’s commands are executed if the target is older than any dependencies of that rule. This can result in executing none, any, or all of the double-colon rules.

Double-colon rules with the same target are in fact completely separate from one another. Each double-colon rule is processed individually, just as rules with different targets are processed.

The double-colon rules for a target are executed in the order they appear in the makefile. However, the cases where double-colon rules really make sense are those where the order of executing the commands would not matter.

Double-colon rules are somewhat obscure and not often very useful; they provide a mechanism for cases in which the method used to update a target differs depending on which dependency files caused the update, and such cases are rare.

Each double-colon rule should specify commands; if it does not, an implicit rule will be used if one applies. See Chapter 10 [Using Implicit Rules], page 107.

4.12 Generating Dependencies Automatically

In the makefile for a program, many of the rules you need to write often say only that some object file depends on some header file. For example, if ‘main.c’ uses ‘defs.h’ via an #include, you would write:

     main.o: defs.h

You need this rule so that make knows that it must remake ‘main.o’ whenever ‘defs.h’ changes. You can see that for a large program you would have to write dozens of such rules in your makefile. And, you must always be very careful to update the makefile every time you add or remove an

#include.

To avoid this hassle, most modern C compilers can write these rules for you, by looking at the #include lines in the source files. Usually this is done with the ‘-M’ option to the compiler. For example, the command:

     cc -M main.c

generates the output:

     main.o : main.c defs.h

Thus you no longer have to write all those rules yourself. The compiler will do it for you.

With old make programs, it was traditional practice to use this compiler feature to generate dependencies on demand with a command like ‘make depend’. That command would create a file ‘depend’ containing all the automatically-generated dependencies; then the makefile could use include to read them in (see Section 3.3 [Include], page 24).

In GNU make, the feature of remaking makefiles makes this practice obsolete—you need never tell make explicitly to regenerate the dependencies, because it always regenerates any makefile that is out of date. See Section 3.5 [Remaking Makefiles], page 26.

The practice we recommend for automatic dependency generation is to have one makefile corresponding to each source file. For each source file ‘name.c’ there is a makefile ‘name.d’ which lists what files the object file ‘name.o’ depends on. That way only the source files that have changed need to be rescanned to produce the new dependencies.

Here is the pattern rule to generate a file of dependencies (i.e., a makefile) called ‘name.d’ from a C source file called ‘name.c’:

     %.d: %.c
     $(CC) -M $(CPPFLAGS) $< | sed 's/$*.o/& $@/g' > $@

See Section 10.5 [Pattern Rules], page 115, for information on defining pattern rules. The purpose of the sed command is to translate (for example):

     main.o : main.c defs.h

into:

     main.o main.d : main.c defs.h

This makes each ‘.d’ file depend on all the source and header files that the corresponding ‘.o’ file depends on. make then knows it must regenerate the dependencies whenever any of the source or header files changes.

Once you’ve defined the rule to remake the ‘.d’ files, you then use the include directive to read them all in. See Section 3.3 [Include], page 24. For example:

     sources = foo.c bar.c
     include $(sources:.c=.d)

(This example uses a substitution variable reference to translate the list of source files ‘foo.c bar.c’ into a list of dependency makefiles, ‘foo.d bar.d’. See Section 6.3.1 [Substitution Refs], page 69, for full information on substitution references.) Since the ‘.d’ files are makefiles like any others, make will remake them as necessary with no further work from you. See Section 3.5 [Remaking Makefiles], page 26.

50 GNU make

5 Writing the Commands in Rules

The commands of a rule consist of shell command lines to be executed one by one. Each command line must start with a tab, except that the first command line may be attached to the target-and-dependencies line with a semicolon in between. Blank lines and lines of just comments may appear among the command lines; they are ignored.

Users use many different shell programs, but commands in makefiles are always interpreted by ‘/bin/sh’ unless the makefile specifies otherwise. See Section 5.2 [Command Execution], page 52.

The shell that is in use determines whether comments can be written on command lines, and what syntax they use. When the shell is ‘/bin/sh’, a ‘#’ starts a comment that extends to the end of the line. The ‘#’ does not have to be at the beginning of a line. Text on a line before a ‘#’ is not part of the comment.

5.1 Command Echoing

Normally make prints each command line before it is executed. We call this echoing because it gives the appearance that you are typing the commands yourself.

When a line starts with ‘@’, the echoing of that line is suppressed. The ‘@’ is discarded before the command is passed to the shell. Typically you would use this for a command whose only effect is to print something, such as an echo command to indicate progress through the makefile:

     @echo About to make distribution files

When make is given the flag ‘-n’ or ‘--just-print’, echoing is all that happens, no execution. See Section 9.7 [Summary of Options], page 103. In this case and only this case, even the commands starting with ‘@’ are printed. This flag is useful for finding out which commands make thinks are necessary without actually doing them.

The ‘-s’ or ‘--silent’ flag to make prevents all echoing, as if all commands started with ‘@’. A rule in the makefile for the special target .SILENT has the same effect (see Section 4.7 [Special Built-in Target Names], page 40). .SILENT is essentially obsolete since ‘@’ is more flexible.

5.2 Command Execution

When it is time to execute commands to update a target, they are executed by making a new subshell for each line. (In practice, make may take shortcuts that do not affect the results.)

Please note: this implies that shell commands such as cd that set variables local to each process will not affect the following command lines. If you want to use cd to affect the next command, put the two on a single line with a semicolon between them. Then make will consider them a single command and pass them, together, to a shell which will execute them in sequence. For example:

     foo : bar/lose
             cd bar; gobble lose > ../foo

If you would like to split a single shell command into multiple lines of text, you must use a backslash at the end of all but the last subline. Such a sequence of lines is combined into a single line, by deleting the backslash-newline sequences, before passing it to the shell. Thus, the following is equivalent to the preceding example:

     foo : bar/lose
             cd bar;  \
             gobble lose > ../foo

The program used as the shell is taken from the variable SHELL. By default, the program ‘/bin/sh’ is used.

Unlike most variables, the variable SHELL is never set from the environment. This is because the SHELL environment variable is used to specify your personal choice of shell program for interactive use. It would be very bad for personal choices like this to affect the functioning of makefiles. See Section 6.9 [Variables from the Environment], page 78.

5.3 Parallel Execution

GNU make knows how to execute several commands at once. Normally, make will execute only one command at a time, waiting for it to finish before executing the next. However, the ‘-j’ or ‘--jobs’ option tells make to execute many commands simultaneously.

If the ‘-j’ option is followed by an integer, this is the number of commands to execute at once; this is called the number of job slots. If there is nothing looking like an integer after the ‘-j’ option, there is no limit on the number of job slots. The default number of job slots is one, which means serial execution (one thing at a time).

One unpleasant consequence of running several commands simultaneously is that output from all of the commands comes when the commands send it, so messages from different commands may be interspersed.

Another problem is that two processes cannot both take input from the same device; so to make sure that only one command tries to take input from the terminal at once, make will invalidate the standard input streams of all but one running command. This means that attempting to read from standard input will usually be a fatal error (a ‘Broken pipe’ signal) for most child processes if there are several.

It is unpredictable which command will have a valid standard input stream (which will come from the terminal, or wherever you redirect the standard input of make). The first command run will always get it first, and the first command started after that one finishes will get it next, and so on.

We will change how this aspect of make works if we find a better alternative. In the mean time, you should not rely on any command using standard input at all if you are using the parallel execution feature; but if you are not using this feature, then standard input works normally in all commands.

If a command fails (is killed by a signal or exits with a nonzero status), and errors are not ignored for that command (see Section 5.4 [Errors in Commands], page 54), the remaining command lines to remake the same target will not be run. If a command fails and the ‘-k’ or ‘--keep-going’ option was not given (see Section 9.7 [Summary of Options], page 103), make aborts execution. If make terminates for any reason (including a signal) with child processes running, it waits for them to finish before actually exiting.

When the system is heavily loaded, you will probably want to run fewer jobs than when it is lightly loaded. You can use the ‘-l’ option to tell make to limit the number of jobs to run at once, based on the load average. The ‘-l’ or ‘--max-load’ option is followed by a floating-point number. For example,

     -l 2.5

will not let make start more than one job if the load average is above 2.5. The ‘-l’ option with no following number removes the load limit, if one was given with a previous ‘-l’ option.

More precisely, when make goes to start up a job, and it already has at least one job running, it checks the current load average; if it is not lower than the limit given with ‘-l’, make waits until the load average goes below that limit, or until all the other jobs finish.

By default, there is no load limit.

5.4 Errors in Commands

After each shell command returns, make looks at its exit status. If the command completed successfully, the next command line is executed in a new shell; after the last command line is finished, the rule is finished.

If there is an error (the exit status is nonzero), make gives up on the current rule, and perhaps on all rules.

Sometimes the failure of a certain command does not indicate a problem. For example, you may use the mkdir command to ensure that a directory exists. If the directory already exists, mkdir will report an error, but you probably want make to continue regardless.

To ignore errors in a command line, write a ‘-’ at the beginning of the line’s text (after the initial tab). The ‘-’ is discarded before the command is passed to the shell for execution.

For example,

     clean:  -rm -f *.o

This causes rm to continue even if it is unable to remove a file.

When you run make with the ‘-i’ or ‘--ignore-errors’ flag, errors are ignored in all commands of all rules. A rule in the makefile for the special target .IGNORE has the same effect. These ways of ignoring errors are obsolete because ‘-’ is more flexible.

When errors are to be ignored, because of either a ‘-’ or the ‘-i’ flag, make treats an error return just like success, except that it prints out a message that tells you the status code the command exited with, and says that the error has been ignored.

When an error happens that make has not been told to ignore, it implies that the current target cannot be correctly remade, and neither can any other that depends on it either directly or indirectly. No further commands will be executed for these targets, since their preconditions have not been achieved.

Normally make gives up immediately in this circumstance, returning a nonzero status. However, if the ‘-k’ or ‘--keep-going’ flag is specified, make continues to consider the other dependencies of the pending targets, remaking them if necessary, before it gives up and returns nonzero status. For example, after an error in compiling one object file, ‘make -k’ will continue compiling other object files even though it already knows that linking them will be impossible. See Section 9.7 [Summary of Options], page 103.

The usual behavior assumes that your purpose is to get the specified targets up to date; once make learns that this is impossible, it might as well report the failure immediately. The ‘-k’ option says that the real purpose is to test as many of the changes made in the program as possible, perhaps to find several independent problems so that you can correct them all before the next attempt to compile. This is why Emacs’ compile command passes the ‘-k’ flag by default.

5.5 Interrupting or Killing make

If make gets a fatal signal while a command is executing, it may delete the target file that the command was supposed to update. This is done if the target file’s last-modification time has changed since make first checked it.

The purpose of deleting the target is to make sure that it is remade from scratch when make is next run. Why is this? Suppose you type Ctrl-c while a compiler is running, and it has begun to write an object file ‘foo.o’. The Ctrl-c kills the compiler, resulting in an incomplete file whose last-modification time is newer than the source file ‘foo.c’. But make also receives the Ctrl-c signal and deletes this incomplete file. If make did not do this, the next invocation of make would think that ‘foo.o’ did not require updating—resulting in a strange error message from the linker when it tries to link an object file half of which is missing.

You can prevent the deletion of a target file in this way by making the special target .PRECIOUS depend on it. Before remaking a target, make checks to see whether it appears on the dependencies of .PRECIOUS, and thereby decides whether the target should be deleted if a signal happens. Some reasons why you might do this are that the target is updated in some atomic fashion, or exists only to record a modification-time (its contents do not matter), or must exist at all times to prevent other sorts of trouble.

5.6 Recursive Use of make

Recursive use of make means using make as a command in a makefile. This technique is useful when you want separate makefiles for various subsystems that compose a larger system. For example, suppose you have a subdirectory ‘subdir’ which has its own makefile, and you would like the containing directory’s makefile to run make on the subdirectory. You can do it by writing this:

     subsystem:
             cd subdir; $(MAKE)

or, equivalently, this (see Section 9.7 [Summary of Options], page 103):

     subsystem:
             $(MAKE) -C subdir

You can write recursive make commands just by copying this example, but there are many things to know about how they work and why, and about how the sub-make relates to the top-level make.

5.6.1 How the MAKE Variable Works

Recursive make commands should always use the variable MAKE, not the explicit command name ‘make’, as shown here:

     subsystem:
             cd subdir; $(MAKE)

The value of this variable is the file name with which make was invoked. If this file name was ‘/bin/make’, then the command executed is ‘cd subdir; /bin/make’. If you use a special version of make to run the top-level makefile, the same special version will be executed for recursive invocations.

Also, any arguments that define variable values are added to MAKE, so the sub-make gets them too. Thus, if you do ‘make CFLAGS=-O’, so that all C compilations will be optimized, the sub-make

is run with ‘cd subdir; /bin/make CFLAGS=-O’.

The MAKE variable actually just refers to two other variables which contain these special values.

In fact, MAKE is always defined as ‘$(MAKE_COMMAND) $(MAKEOVERRIDES)’. The variable MAKE_

COMMAND is the file name with which make was invoked (such as ‘/bin/make’, above). The variable MAKEOVERRIDES contains definitions for the variables defined on the command line; in the above example, its value is ‘CFLAGS=-O’. If you do not want these variable definitions done in all recursive make invocations, you can redefine the MAKEOVERRIDES variable to remove them. You do this in any of the normal ways for defining variables: in a makefile (see Section 6.5 [Setting Variables], page 73); on the command line with an argument like ‘MAKEOVERRIDES=’ (see Section 9.5 [Overriding Variables], page 101); or with an environment variable (see Section 6.9 [Variables from the Environment], page 78).

As a special feature, using the variable MAKE in the commands of a rule alters the effects of the ‘-t’ (‘--touch’), ‘-n’ (‘--just-print’), or ‘-q’ (‘--question’) option. Using the MAKE variable has the same effect as using a ‘+’ character at the beginning of the command line. See Section 9.3 [Instead of Executing the Commands], page 99.

Consider the command ‘make -t’ in the above example. (The ‘-t’ option marks targets as up to date without actually running any commands; see Section 9.3 [Instead of Execution], page 99.) Following the usual definition of ‘-t’, a ‘make -t’ command in the example would create a file named ‘subsystem’ and do nothing else. What you really want it to do is run ‘cd subdir; make -t’; but that would require executing the command, and ‘-t’ says not to execute commands.

The special feature makes this do what you want: whenever a command line of a rule contains the variable MAKE, the flags ‘-t’, ‘-n’ and ‘-q’ do not apply to that line. Command lines containing MAKE are executed normally despite the presence of a flag that causes most commands not to be run. The usual MAKEFLAGS mechanism passes the flags to the sub-make (see Section 5.6.3 [Communicating Options to a Sub-make], page 60), so your request to touch the files, or print the commands, is propagated to the subsystem.

5.6.2 Communicating Variables to a Sub-make

Variable values of the top-level make can be passed to the sub-make through the environment by explicit request. These variables are defined in the sub-make as defaults, but do not override what is specified in the sub-make’s makefile unless you use the ‘-e’ switch (see Section 9.7 [Summary of Options], page 103).

To pass down, or export, a variable, make adds the variable and its value to the environment for running each command. The sub-make, in turn, uses the environment to initialize its table of variable values. See Section 6.9 [Variables from the Environment], page 78.

Except by explicit request, make exports a variable only if it is either defined in the environment initially or set on the command line, and if its name consists only of letters, numbers, and underscores. Some shells cannot cope with environment variable names consisting of characters other than letters, numbers, and underscores.

The special variables SHELL and MAKEFLAGS are always exported. MAKEFILES is exported if you set it to anything.

Variables are not normally passed down if they were created by default by make (see Section 10.3 [Variables Used by Implicit Rules], page 112). The sub-make will define these for itself.

If you want to export specific variables to a sub-make, use the export directive, like this:

export variable … If you want to prevent a variable from being exported, use the unexport directive, like this:

unexport variable … As a convenience, you can define a variable and export it at the same time by doing:

     export variable = value

has the same result as:

variable = value export variable

and

     export variable := value

has the same result as:

variable := value export variable

Likewise,

     export variable += value

is just like:

variable += value export variable

See Section 6.6 [Appending More Text to Variables], page 74.

You may notice that the export and unexport directives work in make in the same way they work in the shell, sh.

If you want all variables to be exported by default, you can use export by itself:

     export

This tells make that variables which are not explicitly mentioned in an export or unexport directive should be exported. Any variable given in an unexport directive will still not be exported. If you use export by itself to export variables by default, variables whose names contain characters other than alphanumerics and underscores will not be exported unless specifically mentioned in an export directive.

The behavior elicited by an export directive by itself was the default in older versions of GNU make. If your makefiles depend on this behavior and you want to be compatible with old versions of make, you can write a rule for the special target .EXPORT_ALL_VARIABLES instead of using the export directive. This will be ignored by old makes, while the export directive will cause a syntax error.

Likewise, you can use unexport by itself to tell make not to export variables by default. Since this is the default behavior, you would only need to do this if export had been used by itself earlier (in an included makefile, perhaps). You cannot use export and unexport by themselves to have variables exported for some commands and not for others. The last export or unexport directive that appears by itself determines the behavior for the entire run of make.

As a special feature, the variable MAKELEVEL is changed when it is passed down from level to level. This variable’s value is a string which is the depth of the level as a decimal number. The value is ‘0’ for the top-level make; ‘1’ for a sub-make, ‘2’ for a sub-sub-make, and so on. The incrementation happens when make sets up the environment for a command.

The main use of MAKELEVEL is to test it in a conditional directive (see Chapter 7 [Conditional Parts of Makefiles], page 79); this way you can write a makefile that behaves one way if run recursively and another way if run directly by you.

You can use the variable MAKEFILES to cause all sub-make commands to use additional makefiles. The value of MAKEFILES is a whitespace-separated list of file names. This variable, if defined in the outer-level makefile, is passed down through the environment; then it serves as a list of extra makefiles for the sub-make to read before the usual or specified ones. See Section 3.4 [The Variable

MAKEFILES], page 26.

5.6.3 Communicating Options to a Sub-make

Flags such as ‘-s’ and ‘-k’ are passed automatically to the sub-make through the variable MAKEFLAGS. This variable is set up automatically by make to contain the flag letters that make received. Thus, if you do ‘make -ks’ then MAKEFLAGS gets the value ‘ks’.

As a consequence, every sub-make gets a value for MAKEFLAGS in its environment. In response, it takes the flags from that value and processes them as if they had been given as arguments. See Section 9.7 [Summary of Options], page 103.

The options ‘-C’, ‘-f’, ‘-I’, ‘-o’, and ‘-W’ are not put into MAKEFLAGS; these options are not passed down.

The ‘-j’ option is a special case (see Section 5.3 [Parallel Execution], page 52). If you set it to some numeric value, ‘-j 1’ is always put into MAKEFLAGS instead of the value you specified. This is because if the ‘-j’ option were passed down to sub-makes, you would get many more jobs running in parallel than you asked for. If you give ‘-j’ with no numeric argument, meaning to run as many jobs as possible in parallel, this is passed down, since multiple infinities are no more than one.

If you do not want to pass the other flags down, you must change the value of MAKEFLAGS, like this:

     MAKEFLAGS=
     subsystem:
             cd subdir; $(MAKE)

or like this:

     subsystem:
             cd subdir; $(MAKE) MAKEFLAGS=

A similar variable MFLAGS exists also, for historical compatibility. It has the same value as MAKEFLAGS except that a hyphen is added at the beginning if it is not empty. MFLAGS was traditionally used explicitly in the recursive make command, like this:

     subsystem:
             cd subdir; $(MAKE) $(MFLAGS)

but now MAKEFLAGS makes this usage redundant.

The MAKEFLAGS and MFLAGS variables can also be useful if you want to have certain options, such as ‘-k’ (see Section 9.7 [Summary of Options], page 103) set each time you run make. Just put ‘MAKEFLAGS=k’ or ‘MFLAGS=-k’ in your environment. These variables may also be set in makefiles, so a makefile can specify additional flags that should also be in effect for that makefile.

If you do put MAKEFLAGS or MFLAGS in your environment, you should be sure not to include any options that will drastically affect the actions of make and undermine the purpose of makefiles and of make itself. For instance, the ‘-t’, ‘-n’, and ‘-q’ options, if put in one of these variables, could have disastrous consequences and would certainly have at least surprising and probably annoying effects.

5.6.4 The ‘--print-directory’ Option

If you use several levels of recursive make invocations, the ‘-w’ or ‘--print-directory’ option can make the output a lot easier to understand by showing each directory as make starts processing it and as make finishes processing it. For example, if ‘make -w’ is run in the directory ‘/u/gnu/make’, make will print a line of the form:

     make: Entering directory `/u/gnu/make'.

before doing anything else, and a line of the form:

     make: Leaving directory `/u/gnu/make'.

when processing is completed.

Normally, you do not need to specify this option because ‘make’ does it for you: ‘-w’ is turned on automatically when you use the ‘-C’ option, and in sub-makes. make will not automatically turn on ‘-w’ if you also use ‘-s’, which says to be silent, or if you use ‘--no-print-directory’ to explicitly disable it.

5.7 Defining Canned Command Sequences

When the same sequence of commands is useful in making various targets, you can define it as a canned sequence with the define directive, and refer to the canned sequence from the rules for those targets. The canned sequence is actually a variable, so the name must not conflict with other variable names.

Here is an example of defining a canned sequence of commands:

     define run-yacc
     yacc $(firstword $^)
     mv y.tab.c $@
     endef

Here run-yacc is the name of the variable being defined; endef marks the end of the definition; the lines in between are the commands. The define directive does not expand variable references and function calls in the canned sequence; the ‘$’ characters, parentheses, variable names, and so on, all become part of the value of the variable you are defining. See Section 6.8 [Defining Variables Verbatim], page 77, for a complete explanation of define.

The first command in this example runs Yacc on the first dependency of whichever rule uses the canned sequence. The output file from Yacc is always named ‘y.tab.c’. The second command moves the output to the rule’s target file name.

To use the canned sequence, substitute the variable into the commands of a rule. You can substitute it like any other variable (see Section 6.1 [Basics of Variable References], page 65). Because variables defined by define are recursively expanded variables, all the variable references you wrote inside the define are expanded now. For example:

     foo.c : foo.y
             $(run-yacc)

‘foo.y’ will be substituted for the variable ‘$^’ when it occurs in run-yacc’s value, and ‘foo.c’ for ‘$@’.

This is a realistic example, but this particular one is not needed in practice because make has an implicit rule to figure out these commands based on the file names involved (see Chapter 10 [Using Implicit Rules], page 107).

In command execution, each line of a canned sequence is treated just as if the line appeared on its own in the rule, preceded by a tab. In particular, make invokes a separate subshell for each line. You can use the special prefix characters that affect command lines (‘@’, ‘-’, and ‘+’) on each line of a canned sequence. See Chapter 5 [Writing the Commands in Rules], page 51. For example, using this canned sequence:

     define frobnicate
     @echo "frobnicating target $@"
     frob-step-1 $< -o $@-step-1
     frob-step-2 $@-step-1 -o $@
     endef

make will not echo the first line, the echo command. But it will echo the following two command lines.

On the other hand, prefix characters on the command line that refers to a canned sequence apply to every line in the sequence. So the rule:

     frob.out: frob.in
     @$(frobnicate)

does not echo any commands. (See Section 5.1 [Command Echoing], page 51, for a full explanation of ‘@’.)

5.8 Using Empty Commands

It is sometimes useful to define commands which do nothing. This is done simply by giving a command that consists of nothing but whitespace. For example:

     target: ;

defines an empty command string for ‘target’. You could also use a line beginning with a tab character to define an empty command string, but this would be confusing because such a line looks empty.

You may be wondering why you would want to define a command string that does nothing. The only reason this is useful is to prevent a target from getting implicit commands (from implicit rules or the .DEFAULT special target; see Chapter 10 [Implicit Rules], page 107 and see Section 10.6 [Defining Last-Resort Default Rules], page 122).

You may be inclined to define empty command strings for targets that are not actual files, but only exist so that their dependencies can be remade. However, this is not the best way to do that, because the dependencies may not be remade properly if the target file actually does exist. See Section 4.4 [Phony Targets], page 37, for a better way to do this.

6 How to Use Variables

A variable is a name defined in a makefile to represent a string of text, called the variable’s value. These values are substituted by explicit request into targets, dependencies, commands, and other parts of the makefile. (In some other versions of make, variables are called macros.)

Variables and functions in all parts of a makefile are expanded when read, except for the shell commands in rules, the right-hand sides of variable definitions using ‘=’, and the bodies of variable definitions using the define directive.

Variables can represent lists of file names, options to pass to compilers, programs to run, directories to look in for source files, directories to write output in, or anything else you can imagine.

A variable name may be any sequence of characters not containing ‘:’, ‘#’, ‘=’, or leading or trailing whitespace. However, variable names containing characters other than letters, numbers, and underscores should be avoided, as they may be given special meanings in the future, and with some shells they cannot be passed through the environment to a sub-make (see Section 5.6.2 [Communicating Variables to a Sub-make], page 57).

It is traditional to use upper case letters in variable names, but we recommend using lower case letters for variable names that serve internal purposes in the makefile, and reserving upper case for parameters that control implicit rules or for parameters that the user should override with command options (see Section 9.5 [Overriding Variables], page 101).

6.1 Basics of Variable References

To substitute a variable’s value, write a dollar sign followed by the name of the variable in parentheses or braces: either ‘$(foo)’ or ‘${foo}’ is a valid reference to the variable foo. This special significance of ‘$’ is why you must write ‘$$’ to have the effect of a single dollar sign in a file name or command.

Variable references can be used in any context: targets, dependencies, commands, most directives, and new variable values. Here is an example of a common case, where a variable holds the names of all the object files in a program:

     objects = program.o foo.o utils.o
     program : $(objects)
             cc -o program $(objects)
     $(objects) : defs.h

Variable references work by strict textual substitution. Thus, the rule

     foo = c
     prog.o : prog.$(foo)
             $(foo)$(foo) -$(foo) prog.$(foo)

could be used to compile a C program ‘prog.c’. Since spaces before the variable value are ignored in variable assignments, the value of foo is precisely ‘c’. (Don’t actually write your makefiles this way!)

A dollar sign followed by a character other than a dollar sign, open-parenthesis or open-brace treats that single character as the variable name. Thus, you could reference the variable x with ‘$x’. However, this practice is strongly discouraged, except in the case of the automatic variables (see Section 10.5.3 [Automatic Variables], page 118).

6.2 The Two Flavors of Variables

There are two ways that a variable in GNU make can have a value; we call them the two flavors of variables. The two flavors are distinguished in how they are defined and in what they do when expanded.

The first flavor of variable is a recursively expanded variable. Variables of this sort are defined by lines using ‘=’ (see Section 6.5 [Setting Variables], page 73) or by the define directive (see Section 6.8 [Defining Variables Verbatim], page 77). The value you specify is installed verbatim; if it contains references to other variables, these references are expanded whenever this variable is substituted (in the course of expanding some other string). When this happens, it is called recursive expansion.

For example,

     foo = $(bar)
     bar = $(ugh)

     ugh = Huh?
     all:;echo $(foo)

will echo ‘Huh?’: ‘$(foo)’ expands to ‘$(bar)’ which expands to ‘$(ugh)’ which finally expands to

‘Huh?’.

This flavor of variable is the only sort supported by other versions of make. It has its advantages and its disadvantages. An advantage (most would say) is that:

     CFLAGS = $(include_dirs) -O
     include_dirs = -Ifoo -Ibar

will do what was intended: when ‘CFLAGS’ is expanded in a command, it will expand to ‘-Ifoo -Ibar -O’. A major disadvantage is that you cannot append something on the end of a variable, as in

     CFLAGS = $(CFLAGS) -O

because it will cause an infinite loop in the variable expansion. (Actually make detects the infinite loop and reports an error.)

Another disadvantage is that any functions (see Chapter 8 [Functions for Transforming Text], page 85) referenced in the definition will be executed every time the variable is expanded. This makes make run slower; worse, it causes the wildcard and shell functions to give unpredictable results because you cannot easily control when they are called, or even how many times.

To avoid all the problems and inconveniences of recursively expanded variables, there is another flavor: simply expanded variables.

Simply expanded variables are defined by lines using ‘:=’ (see Section 6.5 [Setting Variables], page 73). The value of a simply expanded variable is scanned once and for all, expanding any references to other variables and functions, when the variable is defined. The actual value of the simply expanded variable is the result of expanding the text that you write. It does not contain any references to other variables; it contains their values as of the time this variable was defined. Therefore,

     x := foo
     y := $(x) bar
     x := later

is equivalent to

     y := foo bar
     x := later

When a simply expanded variable is referenced, its value is substituted verbatim.

Here is a somewhat more complicated example, illustrating the use of ‘:=’ in conjunction with the shell function. (See Section 8.6 [The shell Function], page 95.) This example also shows use of the variable MAKELEVEL, which is changed when it is passed down from level to level. (See Section 5.6.2 [Communicating Variables to a Sub-make], page 57, for information about MAKELEVEL.)

     ifeq (0,${MAKELEVEL})
     cur-dir   := $(shell pwd)
     whoami    := $(shell whoami)
     host-type := $(shell arch)
     MAKE := ${MAKE} host-type=${host-type} whoami=${whoami}
     endif

An advantage of this use of ‘:=’ is that a typical ‘descend into a directory’ command then looks like this:

     ${subdirs}:
           ${MAKE} cur-dir=${cur-dir}/$@ -C $@ all

Simply expanded variables generally make complicated makefile programming more predictable because they work like variables in most programming languages. They allow you to redefine a variable using its own value (or its value processed in some way by one of the expansion functions) and to use the expansion functions much more efficiently (see Chapter 8 [Functions for Transforming Text], page 85).

You can also use them to introduce controlled leading or trailing spaces into variable values. Such spaces are discarded from your input before substitution of variable references and function calls; this means you can include leading or trailing spaces in a variable value by protecting them with variable references, like this:

     nullstring :=
     space := $(nullstring) $(nullstring)

Here the value of the variable space is precisely one space.

6.3 Advanced Features for Reference to Variables

This section describes some advanced features you can use to reference variables in more flexible ways.

6.3.1 Substitution References

A substitution reference substitutes the value of a variable with alterations that you specify. It has the form ‘$(var:a=b)’ (or ‘${var:a=b}’) and its meaning is to take the value of the variable var, replace every a at the end of a word with b in that value, and substitute the resulting string.

When we say “at the end of a word”, we mean that a must appear either followed by whitespace or at the end of the value in order to be replaced; other occurrences of a in the value are unaltered. For example:

     foo := a.o b.o c.o
     bar := $(foo:.o=.c)

sets ‘bar’ to ‘a.c b.c c.c’. See Section 6.5 [Setting Variables], page 73.

A substitution reference is actually an abbreviation for use of the patsubst expansion function (see Section 8.2 [Functions for String Substitution and Analysis], page 86). We provide substitution references as well as patsubst for compatibility with other implementations of make.

Another type of substitution reference lets you use the full power of the patsubst function. It has the same form ‘$(var:a=b)’ described above, except that now a must contain a single ‘%’ character. This case is equivalent to ‘$(patsubst a,b,$(var))’. See Section 8.2 [Functions for String Substitution and Analysis], page 86, for a description of the patsubst function.

For example:

     foo := a.o b.o c.o
     bar := $(foo:%.o=%.c)
sets ‘bar’ to ‘a.c b.c c.c’.

6.3.2 Computed Variable Names

Computed variable names are a complicated concept needed only for sophisticated makefile programming. For most purposes you need not consider them, except to know that making a variable with a dollar sign in its name might have strange results. However, if you are the type that wants to understand everything, or you are actually interested in what they do, read on.

Variables may be referenced inside the name of a variable. This is called a computed variable name or a nested variable reference. For example,

     x = y
     y = z
     a := $($(x))

defines a as ‘z’: the ‘$(x)’ inside ‘$($(x))’ expands to ‘y’, so ‘$($(x))’ expands to ‘$(y)’ which in turn expands to ‘z’. Here the name of the variable to reference is not stated explicitly; it is computed by expansion of ‘$(x)’. The reference ‘$(x)’ here is nested within the outer variable reference.

The previous example shows two levels of nesting, but any number of levels is possible. For example, here are three levels:

     x = y
     y = z
     z = u
     a := $($($(x)))

Here the innermost ‘$(x)’ expands to ‘y’, so ‘$($(x))’ expands to ‘$(y)’ which in turn expands to ‘z’; now we have ‘$(z)’, which becomes ‘u’.

References to recursively-expanded variables within a variable name are reexpanded in the usual fashion. For example:

     x = $(y)
     y = z
     z = Hello
     a := $($(x))

defines a as ‘Hello’: ‘$($(x))’ becomes ‘$($(y))’ which becomes ‘$(z)’ which becomes ‘Hello’.

Nested variable references can also contain modified references and function invocations (see Chapter 8 [Functions for Transforming Text], page 85), just like any other reference. For example, using the subst function (see Section 8.2 [Functions for String Substitution and Analysis], page 86):

x = variable1
variable2 := Hello
y = $(subst 1,2,$(x))
z = y
a := $($($(z)))

eventually defines a as ‘Hello’. It is doubtful that anyone would ever want to write a nested reference as convoluted as this one, but it works: ‘$($($(z)))’ expands to ‘$($(y))’ which becomes ‘$($(subst 1,2,$(x)))’. This gets the value ‘variable1’ from x and changes it by substitution to ‘variable2’, so that the entire string becomes ‘$(variable2)’, a simple variable reference whose value is ‘Hello’.

A computed variable name need not consist entirely of a single variable reference. It can contain several variable references, as well as some invariant text. For example,

a_dirs := dira dirb
1_dirs := dir1 dir2
a_files := filea fileb
1_files := file1 file2
ifeq "$(use_a)" "yes"
a1 := a
else
a1 := 1
endif
ifeq "$(use_dirs)" "yes"
df := dirs
else
df := files
endif
dirs := $($(a1)_$(df))

will give dirs the same value as a_dirs, 1_dirs, a_files or 1_files depending on the settings

of use_a and use_dirs.

Computed variable names can also be used in substitution references:

a_objects := a.o b.o c.o
1_objects := 1.o 2.o 3.o
sources := $($(a1)_objects:.o=.c)

defines sources as either ‘a.c b.c c.c’ or ‘1.c 2.c 3.c’, depending on the value of a1.

The only restriction on this sort of use of nested variable references is that they cannot specify part of the name of a function to be called. This is because the test for a recognized function name is done before the expansion of nested references. For example,

ifdef do_sort
func := sort
else
func := strip
endif
bar := a d b g q c
foo := $($(func) $(bar))

attempts to give ‘foo’ the value of the variable ‘sort a d b g q c’ or ‘strip a d b g q c’, rather than giving ‘a d b g q c’ as the argument to either the sort or the strip function. This restriction could be removed in the future if that change is shown to be a good idea.

You can also use computed variable names in the left-hand side of a variable assignment, or in a define directive, as in:

dir = foo
$(dir)_sources := $(wildcard $(dir)/*.c)
define $(dir)_print
lpr $($(dir)_sources)
endef

This example defines the variables ‘dir’, ‘foo_sources’, and ‘foo_print’.

Note that nested variable references are quite different from recursively expanded variables (see Section 6.2 [The Two Flavors of Variables], page 66), though both are used together in complex ways when doing makefile programming.

6.4 How Variables Get Their Values

Variables can get values in several different ways:

• You can specify an overriding value when you run make. See Section 9.5 [Overriding Variables], page 101.

• You can specify a value in the makefile, either with an assignment (see Section 6.5 [Setting Variables], page 73) or with a verbatim definition (see Section 6.8 [Defining Variables Verbatim], page 77).

• Variables in the environment become make variables. See Section 6.9 [Variables from the Environment], page 78.

• Several automatic variables are given new values for each rule. Each of these has a single conventional use. See Section 10.5.3 [Automatic Variables], page 118.

• Several variables have constant initial values. See Section 10.3 [Variables Used by Implicit Rules], page 112.

6.5 Setting Variables

To set a variable from the makefile, write a line starting with the variable name followed by ‘=’ or ‘:=’. Whatever follows the ‘=’ or ‘:=’ on the line becomes the value. For example,

     objects = main.o foo.o bar.o utils.o

defines a variable named objects. Whitespace around the variable name and immediately after the ‘=’ is ignored.

Variables defined with ‘=’ are recursively expanded variables. Variables defined with ‘:=’ are simply expanded variables; these definitions can contain variable references which will be expanded before the definition is made. See Section 6.2 [The Two Flavors of Variables], page 66.

The variable name may contain function and variable references, which are expanded when the line is read to find the actual variable name to use.

There is no limit on the length of the value of a variable except the amount of swapping space on the computer. When a variable definition is long, it is a good idea to break it into several lines by inserting backslash-newline at convenient places in the definition. This will not affect the functioning of make, but it will make the makefile easier to read.

Most variable names are considered to have the empty string as a value if you have never set them. Several variables have built-in initial values that are not empty, but you can set them in the usual ways (see Section 10.3 [Variables Used by Implicit Rules], page 112). Several special variables are set automatically to a new value for each rule; these are called the automatic variables (see Section 10.5.3 [Automatic Variables], page 118).

6.6 Appending More Text to Variables

Often it is useful to add more text to the value of a variable already defined. You do this with a line containing ‘+=’, like this:

     objects += another.o

This takes the value of the variable objects, and adds the text ‘another.o’ to it (preceded by a single space). Thus:

     objects = main.o foo.o bar.o utils.o
     objects += another.o
sets objects to ‘main.o foo.o bar.o utils.o another.o’.

Using ‘+=’ is similar to:

     objects = main.o foo.o bar.o utils.o
     objects := $(objects) another.o

but differs in ways that become important when you use more complex values.

When the variable in question has not been defined before, ‘+=’ acts just like normal ‘=’: it defines a recursively-expanded variable. However, when there is a previous definition, exactly what ‘+=’ does depends on what flavor of variable you defined originally. See Section 6.2 [The Two Flavors of Variables], page 66, for an explanation of the two flavors of variables.

When you add to a variable’s value with ‘+=’, make acts essentially as if you had included the extra text in the initial definition of the variable. If you defined it first with ‘:=’, making it a simply-expanded variable, ‘+=’ adds to that simply-expanded definition, and expands the new text before appending it to the old value just as ‘:=’ does (see Section 6.5 [Setting Variables], page 73, for a full explanation of ‘:=’). In fact,

     variable := value
     variable += more

is exactly equivalent to:

     variable := value
     variable := $(variable) more

On the other hand, when you use ‘+=’ with a variable that you defined first to be recursivelyexpanded using plain ‘=’, make does something a bit different. Recall that when you define a recursively-expanded variable, make does not expand the value you set for variable and function references immediately. Instead it stores the text verbatim, and saves these variable and function references to be expanded later, when you refer to the new variable (see Section 6.2 [The Two Flavors of Variables], page 66). When you use ‘+=’ on a recursively-expanded variable, it is this unexpanded text to which make appends the new text you specify.

     variable = value
     variable += more

is roughly equivalent to:

     temp = value
     variable = $(temp) more

except that of course it never defines a variable called temp. The importance of this comes when the variable’s old value contains variable references. Take this common example:

     CFLAGS = $(includes) -O

…

     CFLAGS += -pg # enable profiling

The first line defines the CFLAGS variable with a reference to another variable, includes. (CFLAGS is used by the rules for C compilation; see Section 10.2 [Catalogue of Implicit Rules], page 109.) Using ‘=’ for the definition makes CFLAGS a recursively-expanded variable, meaning ‘$(includes) -O’ is not expanded when make processes the definition of CFLAGS. Thus, includes need not be defined yet for its value to take effect. It only has to be defined before any reference to CFLAGS. If we tried to append to the value of CFLAGS without using ‘+=’, we might do it like this:

     CFLAGS := $(CFLAGS) -pg # enable profiling

This is close, but not quite what we want. Using ‘:=’ redefines CFLAGS as a simply-expanded variable; this means make expands the text ‘$(CFLAGS) -pg’ before setting the variable. If includes is not yet defined, we get ‘ -O -pg’, and a later definition of includes will have no effect. Conversely, by using ‘+=’ we set CFLAGS to the unexpanded value ‘$(includes) -O -pg’. Thus we preserve the reference to includes, so if that variable gets defined at any later point, a reference like ‘$(CFLAGS)’ still uses its value.

6.7 The override Directive

If a variable has been set with a command argument (see Section 9.5 [Overriding Variables], page 101), then ordinary assignments in the makefile are ignored. If you want to set the variable in the makefile even though it was set with a command argument, you can use an override directive, which is a line that looks like this:

override variable = value

or

override variable := value

To append more text to a variable defined on the command line, use:

override variable += more text

See Section 6.6 [Appending More Text to Variables], page 74.

The override directive was not invented for escalation in the war between makefiles and command arguments. It was invented so you can alter and add to values that the user specifies with command arguments.

For example, suppose you always want the ‘-g’ switch when you run the C compiler, but you would like to allow the user to specify the other switches with a command argument just as usual. You could use this override directive:

     override CFLAGS += -g

You can also use override directives with define directives. This is done as you might expect:

     override define foo
     bar
     endef

See the next section for information about define.

6.8 Defining Variables Verbatim

Another way to set the value of a variable is to use the define directive. This directive has an unusual syntax which allows newline characters to be included in the value, which is convenient for defining canned sequences of commands (see Section 5.7 [Defining Canned Command Sequences], page 62).

The define directive is followed on the same line by the name of the variable and nothing more. The value to give the variable appears on the following lines. The end of the value is marked by a line containing just the word endef. Aside from this difference in syntax, define works just like ‘=’: it creates a recursively-expanded variable (see Section 6.2 [The Two Flavors of Variables], page 66). The variable name may contain function and variable references, which are expanded when the directive is read to find the actual variable name to use.

     define two-lines
     echo foo
     echo $(bar)
     endef

The value in an ordinary assignment cannot contain a newline; but the newlines that separate the lines of the value in a define become part of the variable’s value (except for the final newline which precedes the endef and is not considered part of the value).

The previous example is functionally equivalent to this:

     two-lines = echo foo; echo $(bar)

since two commands separated by semicolon behave much like two separate shell commands. However, note that using two separate lines means make will invoke the shell twice, running an independent subshell for each line. See Section 5.2 [Command Execution], page 52.

If you want variable definitions made with define to take precedence over command-line variable definitions, you can use the override directive together with define:

     override define two-lines
     foo
     $(bar)
     endef

See Section 6.7 [The override Directive], page 76.

6.9 Variables from the Environment

Variables in make can come from the environment in which make is run. Every environment variable that make sees when it starts up is transformed into a make variable with the same name and value. But an explicit assignment in the makefile, or with a command argument, overrides the environment. (If the ‘-e’ flag is specified, then values from the environment override assignments in the makefile. See Section 9.7 [Summary of Options], page 103. But this is not recommended practice.)

Thus, by setting the variable CFLAGS in your environment, you can cause all C compilations in most makefiles to use the compiler switches you prefer. This is safe for variables with standard or conventional meanings because you know that no makefile will use them for other things. (But this is not totally reliable; some makefiles set CFLAGS explicitly and therefore are not affected by the value in the environment.)

When make is invoked recursively, variables defined in the outer invocation can be passed to inner invocations through the environment (see Section 5.6 [Recursive Use of make], page 56). By default, only variables that came from the environment or the command line are passed to recursive invocations. You can use the export directive to pass other variables. See Section 5.6.2 [Communicating Variables to a Sub-make], page 57, for full details.

Other use of variables from the environment is not recommended. It is not wise for makefiles to depend for their functioning on environment variables set up outside their control, since this would cause different users to get different results from the same makefile. This is against the whole purpose of most makefiles.

Such problems would be especially likely with the variable SHELL, which is normally present in the environment to specify the user’s choice of interactive shell. It would be very undesirable for this choice to affect make. So make ignores the environment value of SHELL.

7 Conditional Parts of Makefiles

A conditional causes part of a makefile to be obeyed or ignored depending on the values of variables. Conditionals can compare the value of one variable to another, or the value of a variable to a constant string. Conditionals control what make actually “sees” in the makefile, so they cannot be used to control shell commands at the time of execution.

7.1 Example of a Conditional

The following example of a conditional tells make to use one set of libraries if the CC variable is ‘gcc’, and a different set of libraries otherwise. It works by controlling which of two command lines will be used as the command for a rule. The result is that ‘CC=gcc’ as an argument to make changes not only which compiler is used but also which libraries are linked.

     libs_for_gcc = -lgnu
     normal_libs =
     foo: $(objects)
     ifeq ($(CC),gcc)
             $(CC) -o foo $(objects) $(libs_for_gcc)
     else
             $(CC) -o foo $(objects) $(normal_libs)
     endif

This conditional uses three directives: one ifeq, one else and one endif.

The ifeq directive begins the conditional, and specifies the condition. It contains two arguments, separated by a comma and surrounded by parentheses. Variable substitution is performed on both arguments and then they are compared. The lines of the makefile following the ifeq are obeyed if the two arguments match; otherwise they are ignored.

The else directive causes the following lines to be obeyed if the previous conditional failed. In the example above, this means that the second alternative linking command is used whenever the first alternative is not used. It is optional to have an else in a conditional.

The endif directive ends the conditional. Every conditional must end with an endif. Unconditional makefile text follows.

As this example illustrates, conditionals work at the textual level: the lines of the conditional are treated as part of the makefile, or ignored, according to the condition. This is why the larger syntactic units of the makefile, such as rules, may cross the beginning or the end of the conditional.

When the variable CC has the value ‘gcc’, the above example has this effect:

foo: $(objects)
        $(CC) -o foo $(objects) $(libs_for_gcc)

When the variable CC has any other value, the effect is this:

foo: $(objects)
        $(CC) -o foo $(objects) $(normal_libs)

Equivalent results can be obtained in another way by conditionalizing a variable assignment and then using the variable unconditionally:

libs_for_gcc = -lgnu
normal_libs =
ifeq ($(CC),gcc)
  libs=$(libs_for_gcc)
else
  libs=$(normal_libs)
endif
foo: $(objects)
        $(CC) -o foo $(objects) $(libs)

7.2 Syntax of Conditionals

The syntax of a simple conditional with no else is as follows:

conditional-directive text-if-true

endif

The text-if-true may be any lines of text, to be considered as part of the makefile if the condition is true. If the condition is false, no text is used instead.

The syntax of a complex conditional is as follows: conditional-directive text-if-true

     else

text-if-false

     endif

If the condition is true, text-if-true is used; otherwise, text-if-false is used instead. The text-if-false can be any number of lines of text.

The syntax of the conditional-directive is the same whether the conditional is simple or complex. There are four different directives that test different conditions. Here is a table of them:

ifeq (arg1, arg2)

ifeq 'arg1' 'arg2'
ifeq "arg1" "arg2"
ifeq "arg1" 'arg2'
ifeq 'arg1' "arg2"

Expand all variable references in arg1 and arg2 and compare them. If they are identical, the text-if-true is effective; otherwise, the text-if-false, if any, is effective. Often you want to test if a variable has a non-empty value. When the value results from complex expansions of variables and functions, expansions you would consider empty may actually contain whitespace characters and thus are not seen as empty. However, you can use the strip function (see Section 8.2 [Text Functions], page 86) to avoid interpreting whitespace as a non-empty value. For example:

               ifeq ($(strip $(foo)),)

text-if-empty

               endif

will evaluate text-if-empty even if the expansion of $(foo) contains whitespace characters.

ifneq (arg1, arg2)
ifneq 'arg1' 'arg2'
ifneq "arg1" "arg2"
ifneq "arg1" 'arg2'
ifneq 'arg1' "arg2"

Expand all variable references in arg1 and arg2 and compare them. If they are different, the text-if-true is effective; otherwise, the text-if-false, if any, is effective.

ifdef variable-name If the variable variable-name has a non-empty value, the text-if-true is effective; otherwise, the text-if-false, if any, is effective. Variables that have never been defined have an empty value.

Note that ifdef only tests whether a variable has a value. It does not expand the variable to see if that value is nonempty. Consequently, tests using ifdef return true for all definitions except those like foo =. To test for an empty value, use ifeq ($(foo),). For example,

               bar =
               foo = $(bar)
               ifdef foo
               frobozz = yes
               else
               frobozz = no
               endif

sets ‘frobozz’ to ‘yes’, while:

               foo =
               ifdef foo
               frobozz = yes
               else
               frobozz = no
               endif

sets ‘frobozz’ to ‘no’.

ifndef variable-name If the variable variable-name has an empty value, the text-if-true is effective; otherwise, the text-if-false, if any, is effective.

Extra spaces are allowed and ignored at the beginning of the conditional directive line, but a tab is not allowed. (If the line begins with a tab, it will be considered a command for a rule.) Aside from this, extra spaces or tabs may be inserted with no effect anywhere except within the directive name or within an argument. A comment starting with ‘#’ may appear at the end of the line.

The other two directives that play a part in a conditional are else and endif. Each of these directives is written as one word, with no arguments. Extra spaces are allowed and ignored at the beginning of the line, and spaces or tabs at the end. A comment starting with ‘#’ may appear at the end of the line.

Conditionals affect which lines of the makefile make uses. If the condition is true, make reads the lines of the text-if-true as part of the makefile; if the condition is false, make ignores those lines completely. It follows that syntactic units of the makefile, such as rules, may safely be split across the beginning or the end of the conditional.

make evaluates conditionals when it reads a makefile. Consequently, you cannot use automatic variables in the tests of conditionals because they are not defined until commands are run (see Section 10.5.3 [Automatic Variables], page 118).

To prevent intolerable confusion, it is not permitted to start a conditional in one makefile and end it in another. However, you may write an include directive within a conditional, provided you do not attempt to terminate the conditional inside the included file.

7.3 Conditionals that Test Flags

You can write a conditional that tests make command flags such as ‘-t’ by using the variable MAKEFLAGS together with the findstring function (see Section 8.2 [Functions for String Substitution and Analysis], page 86). This is useful when touch is not enough to make a file appear up to date.

The findstring function determines whether one string appears as a substring of another. If you want to test for the ‘-t’ flag, use ‘t’ as the first string and the value of MAKEFLAGS as the other.

For example, here is how to arrange to use ‘ranlib -t’ to finish marking an archive file up to date:

     archive.a: . . .
     ifneq (,$(findstring t,$(MAKEFLAGS)))
             +touch archive.a
             +ranlib -t archive.a
     else
             ranlib archive.a
     endif

The ‘+’ prefix marks those command lines as “recursive” so that they will be executed despite use of the ‘-t’ flag. See Section 5.6 [Recursive Use of make], page 56.

84 GNU make

8 Functions for Transforming Text

Functions allow you to do text processing in the makefile to compute the files to operate on or the commands to use. You use a function in a function call, where you give the name of the function and some text (the arguments) for the function to operate on. The result of the function’s processing is substituted into the makefile at the point of the call, just as a variable might be substituted.

8.1 Function Call Syntax

A function call resembles a variable reference. It looks like this:

$(function arguments)

or like this:

${function arguments}

Here function is a function name; one of a short list of names that are part of make. There is no provision for defining new functions.

The arguments are the arguments of the function. They are separated from the function name by one or more spaces or tabs, and if there is more than one argument, then they are separated by commas. Such whitespace and commas are not part of an argument’s value. The delimiters which you use to surround the function call, whether parentheses or braces, can appear in an argument only in matching pairs; the other kind of delimiters may appear singly. If the arguments themselves contain other function calls or variable references, it is wisest to use the same kind of delimiters for all the references; write ‘$(subst a,b,$(x))’, not ‘$(subst a,b,${x})’. This is because it is clearer, and because only one type of delimiter is matched to find the end of the reference.

The text written for each argument is processed by substitution of variables and function calls to produce the argument value, which is the text on which the function acts. The substitution is done in the order in which the arguments appear.

Commas and unmatched parentheses or braces cannot appear in the text of an argument as written; leading spaces cannot appear in the text of the first argument as written. These characters can be put into the argument value by variable substitution. First define variables comma and space whose values are isolated comma and space characters, then substitute these variables where such characters are wanted, like this:

comma:= ,
empty:=
space:= $(empty) $(empty)
foo:= a b c
bar:= $(subst $(space),$(comma),$(foo))

# bar is now ‘a,b,c’.

Here the subst function replaces each space with a comma, through the value of foo, and substitutes the result.

8.2 Functions for String Substitution and Analysis

Here are some functions that operate on strings:

$(subst from,to,text)

Performs a textual replacement on the text text: each occurrence of from is replaced by to. The result is substituted for the function call. For example,

     $(subst ee,EE,feet on the street)

substitutes the string ‘fEEt on the strEEt’.

$(patsubst pattern,replacement,text)

Finds whitespace-separated words in text that match pattern and replaces them with replacement. Here pattern may contain a ‘%’ which acts as a wildcard, matching any number of any characters within a word. If replacement also contains a ‘%’, the ‘%’ is replaced by the text that matched the ‘%’ in pattern. ‘%’ characters in patsubst function invocations can be quoted with preceding backslashes (‘\’). Backslashes that would otherwise quote ‘%’ characters can be quoted with more backslashes. Backslashes that quote ‘%’ characters or other backslashes are removed from the pattern before it is compared file names or has a stem substituted into it. Backslashes that are not in danger of quoting ‘%’ characters go unmolested. For example, the pattern ‘the\%weird\\%pattern\\’ has ‘the%weird\’ preceding the operative ‘%’ character, and ‘pattern\\’ following it. The final two backslashes are left alone because they cannot affect any ‘%’ character. Whitespace between words is folded into single space characters; leading and trailing whitespace is discarded. For example,

     $(patsubst %.c,%.o,x.c.c bar.c)

produces the value ‘x.c.o bar.o’. Substitution references (see Section 6.3.1 [Substitution References], page 69) are a simpler way to get the effect of the patsubst function:

$(var:pattern=replacement) is equivalent to

$(patsubst pattern,replacement,$(var)) The second shorthand simplifies one of the most common uses of patsubst: replacing the suffix at the end of file names.

$(var:suffix=replacement) is equivalent to

     $(patsubst %suffix,%replacement,$(var))

For example, you might have a list of object files:

     objects = foo.o bar.o baz.o

To get the list of corresponding source files, you could simply write:

     $(objects:.o=.c)

instead of using the general form:

     $(patsubst %.o,%.c,$(objects))
$(strip string)

Removes leading and trailing whitespace from string and replaces each internal sequence of one or more whitespace characters with a single space. Thus, ‘$(strip a b c )’ results in ‘a b c’. The function strip can be very useful when used in conjunction with conditionals. When comparing something with the null string ‘""’ using ifeq or ifneq, you usually want a string of just whitespace to match the null string (see Chapter 7 [Conditionals], page 79). Thus, the following may fail to have the desired results:

     .PHONY: all
     ifneq   "$(needs_made)" ""
     all: $(needs_made)
     else
     all:;@echo 'Nothing to make!'
     endif

Replacing the variable reference ‘$(needs_made)’ with the function call ‘$(strip $(needs_made))’ in the ifneq directive would make it more robust.

$(findstring find,in)

Searches in for an occurrence of find. If it occurs, the value is find; otherwise, the value is empty. You can use this function in a conditional to test for the presence of a specific substring in a given string. Thus, the two examples,

     $(findstring a,a b c)
     $(findstring a,b c)

produce the values ‘a’ and ‘’ (the empty string), respectively. See Section 7.3 [Testing Flags], page 83, for a practical application of findstring.

$(filter pattern…,text)

Removes all whitespace-separated words in text that do not match any of the pattern words, returning only matching words. The patterns are written using ‘%’, just like the patterns used in the patsubst function above. The filter function can be used to separate out different types of strings (such as file names) in a variable. For example:

     sources := foo.c bar.c baz.s ugh.h
     foo: $(sources)
             cc $(filter %.c %.s,$(sources)) -o foo

says that ‘foo’ depends of ‘foo.c’, ‘bar.c’, ‘baz.s’ and ‘ugh.h’ but only ‘foo.c’, ‘bar.c’ and ‘baz.s’ should be specified in the command to the compiler.

$(filter-out pattern. . .,text)

Removes all whitespace-separated words in text that do match the pattern words, returning only the words that do not match. This is the exact opposite of the filter function. For example, given:

     objects=main1.o foo.o main2.o bar.o
     mains=main1.o main2.o

the following generates a list which contains all the object files not in ‘mains’:

     $(filter-out $(mains),$(objects))
$(sort list)

Sorts the words of list in lexical order, removing duplicate words. The output is a list of words separated by single spaces. Thus,

     $(sort foo bar lose)

returns the value ‘bar foo lose’. Incidentally, since sort removes duplicate words, you can use it for this purpose even if you don’t care about the sort order.

Here is a realistic example of the use of subst and patsubst. Suppose that a makefile uses the VPATH variable to specify a list of directories that make should search for dependency files (see Section 4.3.1 [VPATH Search Path for All Dependencies], page 33). This example shows how to tell the C compiler to search for header files in the same list of directories.

The value of VPATH is a list of directories separated by colons, such as ‘src:../headers’. First, the subst function is used to change the colons to spaces:

     $(subst :, ,$(VPATH))

This produces ‘src ../headers’. Then patsubst is used to turn each directory name into a ‘-I’ flag. These can be added to the value of the variable CFLAGS, which is passed automatically to the C compiler, like this:

     override CFLAGS += $(patsubst %,-I%,$(subst :, ,$(VPATH)))

The effect is to append the text ‘-Isrc -I../headers’ to the previously given value of CFLAGS. The override directive is used so that the new value is assigned even if the previous value of CFLAGS was specified with a command argument (see Section 6.7 [The override Directive], page 76).

8.3 Functions for File Names

Several of the built-in expansion functions relate specifically to taking apart file names or lists of file names.

Each of the following functions performs a specific transformation on a file name. The argument of the function is regarded as a series of file names, separated by whitespace. (Leading and trailing whitespace is ignored.) Each file name in the series is transformed in the same way and the results are concatenated with single spaces between them.

$(dir names…)

Extracts the directory-part of each file name in names. The directory-part of the file name is everything up through (and including) the last slash in it. If the file name contains no slash, the directory part is the string ‘./’. For example,

               $(dir src/foo.c hacks)

produces the result ‘src/ ./’.

$(notdir names. . .)

Extracts all but the directory-part of each file name in names. If the file name contains no slash, it is left unchanged. Otherwise, everything through the last slash is removed from it. A file name that ends with a slash becomes an empty string. This is unfortunate, because it means that the result does not always have the same number of whitespaceseparated file names as the argument had; but we do not see any other valid alternative. For example,

               $(notdir src/foo.c hacks)

produces the result ‘foo.c hacks’.

$(suffix names. . .)

Extracts the suffix of each file name in names. If the file name contains a period, the suffix is everything starting with the last period. Otherwise, the suffix is the empty string. This frequently means that the result will be empty when names is not, and if names contains multiple file names, the result may contain fewer file names. For example,

     $(suffix src/foo.c hacks)

produces the result ‘.c’.

$(basename names. . .)

Extracts all but the suffix of each file name in names. If the file name contains a period, the basename is everything starting up to (and not including) the last period. Otherwise, the basename is the entire file name. For example,

     $(basename src/foo.c hacks)

produces the result ‘src/foo hacks’.

$(addsuffix suffix,names. . .)

The argument names is regarded as a series of names, separated by whitespace; suffix is used as a unit. The value of suffix is appended to the end of each individual name and the resulting larger names are concatenated with single spaces between them. For example,

     $(addsuffix .c,foo bar)

produces the result ‘foo.c bar.c’.

$(addprefix prefix,names. . .)

The argument names is regarded as a series of names, separated by whitespace; prefix is used as a unit. The value of prefix is prepended to the front of each individual name and the resulting larger names are concatenated with single spaces between them. For example,

     $(addprefix src/,foo bar)

produces the result ‘src/foo src/bar’.

$(join list1,list2)

Concatenates the two arguments word by word: the two first words (one from each argument) concatenated form the first word of the result, the two second words form the second word of the result, and so on. So the nth word of the result comes from the nth word of each argument. If one argument has more words that the other, the extra words are copied unchanged into the result. For example, ‘$(join a b,.c .o)’ produces ‘a.c b.o’. Whitespace between the words in the lists is not preserved; it is replaced with a single space.

This function can merge the results of the dir and notdir functions, to produce the original list of files which was given to those two functions.

$(word n,text)

Returns the nth word of text. The legitimate values of n start from 1. If n is bigger than the number of words in text, the value is empty. For example,

     $(word 2, foo bar baz)

returns ‘bar’.

$(words text)

Returns the number of words in text. Thus, the last word of text is $(word $(words text),text).

$(firstword names. . .)

The argument names is regarded as a series of names, separated by whitespace. The value is the first name in the series. The rest of the names are ignored. For example,

     $(firstword foo bar)

produces the result ‘foo’. Although $(firstword text) is the same as $(word 1,text), the firstword function is retained for its simplicity.

$(wildcard pattern)

The argument pattern is a file name pattern, typically containing wildcard characters (as in shell file name patterns). The result of wildcard is a space-separated list of the names of existing files that match the pattern. See Section 4.2 [Using Wildcard Characters in File Names], page 30.

8.4 The foreach Function

The foreach function is very different from other functions. It causes one piece of text to be used repeatedly, each time with a different substitution performed on it. It resembles the for command in the shell sh and the foreach command in the C-shell csh.

The syntax of the foreach function is:

$(foreach var,list,text)

The first two arguments, var and list, are expanded before anything else is done; note that the last argument, text, is not expanded at the same time. Then for each word of the expanded value of list, the variable named by the expanded value of var is set to that word, and text is expanded. Presumably text contains references to that variable, so its expansion will be different each time.

The result is that text is expanded as many times as there are whitespace-separated words in list. The multiple expansions of text are concatenated, with spaces between them, to make the result of foreach.

This simple example sets the variable ‘files’ to the list of all files in the directories in the list

‘dirs’:
     dirs := a b c d
     files := $(foreach dir,$(dirs),$(wildcard $(dir)/*))

Here text is ‘$(wildcard $(dir)/*)’. The first repetition finds the value ‘a’ for dir, so it produces the same result as ‘$(wildcard a/*)’; the second repetition produces the result of

‘$(wildcard b/*)’; and the third, that of ‘$(wildcard c/*)’.

This example has the same result (except for setting ‘dirs’) as the following example:

     files := $(wildcard a/* b/* c/* d/*)

When text is complicated, you can improve readability by giving it a name, with an additional variable:

     find_files = $(wildcard $(dir)/*)
     dirs := a b c d
     files := $(foreach dir,$(dirs),$(find_files))

Here we use the variable find_files this way. We use plain ‘=’ to define a recursively-expanding variable, so that its value contains an actual function call to be reexpanded under the control of foreach; a simply-expanded variable would not do, since wildcard would be called only once at the time of defining find_files.

The foreach function has no permanent effect on the variable var; its value and flavor after the foreach function call are the same as they were beforehand. The other values which are taken from list are in effect only temporarily, during the execution of foreach. The variable var is a simply-expanded variable during the execution of foreach. If var was undefined before the foreach function call, it is undefined after the call. See Section 6.2 [The Two Flavors of Variables], page 66.

You must take care when using complex variable expressions that result in variable names because many strange things are valid variable names, but are probably not what you intended. For example,

     files := $(foreach Es escrito en espanol!,b c ch,$(find_files))

might be useful if the value of find_files references the variable whose name is ‘Es escrito en espanol!’ (es un nombre bastante largo, no?), but it is more likely to be a mistake.

8.5 The origin Function

The origin function is unlike most other functions in that it does not operate on the values of variables; it tells you something about a variable. Specifically, it tells you where it came from.

The syntax of the origin function is:

     $(origin variable)

Note that variable is the name of a variable to inquire about; not a reference to that variable. Therefore you would not normally use a ‘$’ or parentheses when writing it. (You can, however, use a variable reference in the name if you want the name not to be a constant.)

The result of this function is a string telling you how the variable variable was defined:

‘undefined’

if variable was never defined.

‘default’

if variable has a default definition, as is usual with CC and so on. See Section 10.3 [Variables Used by Implicit Rules], page 112. Note that if you have redefined a default variable, the origin function will return the origin of the later definition.

‘environment’

if variable was defined as an environment variable and the ‘-e’ option is not turned on (see Section 9.7 [Summary of Options], page 103).

‘environment override’

if variable was defined as an environment variable and the ‘-e’ option is turned on (see Section 9.7 [Summary of Options], page 103).

‘file’

if variable was defined in a makefile.

‘command line’

if variable was defined on the command line.

‘override’

if variable was defined with an override directive in a makefile (see Section 6.7 [The override Directive], page 76).

‘automatic’

if variable is an automatic variable defined for the execution of the commands for each rule (see Section 10.5.3 [Automatic Variables], page 118).

This information is primarily useful (other than for your curiosity) to determine if you want to believe the value of a variable. For example, suppose you have a makefile ‘foo’ that includes another makefile ‘bar’. You want a variable bletch to be defined in ‘bar’ if you run the command ‘make -f bar’, even if the environment contains a definition of bletch. However, if ‘foo’ defined bletch before including ‘bar’, you do not want to override that definition. This could be done by using an override directive in ‘foo’, giving that definition precedence over the later definition in ‘bar’; unfortunately, the override directive would also override any command line definitions. So, ‘bar’ could include:

     ifdef bletch
     ifeq "$(origin bletch)" "environment"
     bletch = barf, gag, etc.
     endif
     endif

If bletch has been defined from the environment, this will redefine it.

If you want to override a previous definition of bletch if it came from the environment, even under ‘-e’, you could instead write:

     ifneq "$(findstring environment,$(origin bletch))" ""
     bletch = barf, gag, etc.
     endif

Here the redefinition takes place if ‘$(origin bletch)’ returns either ‘environment’ or ‘environment override’. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

8.6 The shell Function

The shell function is unlike any other function except the wildcard function (see Section 4.2.3 [The Function wildcard], page 32) in that it communicates with the world outside of make.

The shell function performs the same function that backquotes (‘`’) perform in most shells: it does command expansion. This means that it takes an argument that is a shell command and returns the output of the command. The only processing make does on the result, before substituting it into the surrounding text, is to convert newlines to spaces.

The commands run by calls to the shell function are run when the function calls are expanded. In most cases, this is when the makefile is read in. The exception is that function calls in the commands of the rules are expanded when the commands are run, and this applies to shell function calls like all others.

Here are some examples of the use of the shell function:

     contents := $(shell cat foo)

sets contents to the contents of the file ‘foo’, with a space (rather than a newline) separating each line.

     files := $(shell echo *.c)

sets files to the expansion of ‘*.c’. Unless make is using a very strange shell, this has the same

result as ‘$(wildcard *.c)’.

96 GNU make

9 How to Run make

A makefile that says how to recompile a program can be used in more than one way. The simplest use is to recompile every file that is out of date. Usually, makefiles are written so that if you run make with no arguments, it does just that.

But you might want to update only some of the files; you might want to use a different compiler or different compiler options; you might want just to find out which files are out of date without changing them.

By giving arguments when you run make, you can do any of these things and many others.

9.1 Arguments to Specify the Makefile

The way to specify the name of the makefile is with the ‘-f’ or ‘--file’ option (‘--makefile’ also works). For example, ‘-f altmake’ says to use the file ‘altmake’ as the makefile.

If you use the ‘-f’ flag several times and follow each ‘-f’ with an argument, all the specified files are used jointly as makefiles.

If you do not use the ‘-f’ or ‘--file’ flag, the default is to try ‘GNUmakefile’, ‘makefile’, and ‘Makefile’, in that order, and use the first of these three which exists or can be made (see Chapter 3 [Writing Makefiles], page 23).

9.2 Arguments to Specify the Goals

The goals are the targets that make should strive ultimately to update. Other targets are updated as well if they appear as dependencies of goals, or dependencies of dependencies of goals, etc.

By default, the goal is the first target in the makefile (not counting targets that start with a period). Therefore, makefiles are usually written so that the first target is for compiling the entire program or programs they describe.

You can specify a different goal or goals with arguments to make. Use the name of the goal as an argument. If you specify several goals, make processes each of them in turn, in the order you name them.

Any target in the makefile may be specified as a goal (unless it starts with ‘-’ or contains an ‘=’, in which case it will be parsed as a switch or variable definition, respectively). Even targets not in the makefile may be specified, if make can find implicit rules that say how to make them.

One use of specifying a goal is if you want to compile only a part of the program, or only one of several programs. Specify as a goal each file that you wish to remake. For example, consider a directory containing several programs, with a makefile that starts like this:

     .PHONY: all
     all: size nm ld ar as

If you are working on the program size, you might want to say ‘make size’ so that only the files of that program are recompiled.

Another use of specifying a goal is to make files that are not normally made. For example, there may be a file of debugging output, or a version of the program that is compiled specially for testing, which has a rule in the makefile but is not a dependency of the default goal.

Another use of specifying a goal is to run the commands associated with a phony target (see Section 4.4 [Phony Targets], page 37) or empty target (see Section 4.6 [Empty Target Files to Record Events], page 40). Many makefiles contain a phony target named ‘clean’ which deletes everything except source files. Naturally, this is done only if you request it explicitly with ‘make clean’. Here is a list of typical phony and empty target names:

‘all’ Make all the top-level targets the makefile knows about.

‘clean’ Delete all files that are normally created by running make.

‘mostlyclean’

Like ‘clean’, but may refrain from deleting a few files that people normally don’t want to recompile. For example, the ‘mostlyclean’ target for GCC does not delete ‘libgcc.a’, because recompiling it is rarely necessary and takes a lot of time.

‘distclean’
‘realclean’

‘clobber’ Any of these three might be defined to delete everything that would not be part of a standard distribution. For example, this would delete configuration files or links that you would normally create as preparation for compilation, even if the makefile itself cannot create these files.

‘install’ Copy the executable file into a directory that users typically search for commands; copy any auxiliary files that the executable uses into the directories where it will look for them.

‘print’ Print listings of the source files that have changed. ‘tar’ Create a tar file of the source files. ‘shar’ Create a shell archive (shar file) of the source files. ‘dist’ Create a distribution file of the source files. This might be a tar file, or a shar file, or

a compressed version of one of the above, or even more than one of the above.

‘TAGS’ Update a tags table for this program.

‘check’

‘test’ Perform self tests on the program this makefile builds.

9.3 Instead of Executing the Commands

The makefile tells make how to tell whether a target is up to date, and how to update each target. But updating the targets is not always what you want. Certain options specify other activities for

make.
‘-n’
‘--just-print’
‘--dry-run’
‘--recon’

“No-op”. The activity is to print what commands would be used to make the targets up to date, but not actually execute them.

‘-t’
‘--touch’

“Touch”. The activity is to mark the targets as up to date without actually changing them. In other words, make pretends to compile the targets but does not really change their contents.

‘-q’
‘--question’

“Question”. The activity is to find out silently whether the targets are up to date already; but execute no commands in either case. In other words, neither compilation nor output will occur.

‘-W’
‘--what-if’
‘--assume-new’
‘--new-file’

“What if”. Each ‘-W’ flag is followed by a file name. The given files’ modification times are recorded by make as being the present time, although the actual modification times remain the same. You can use the ‘-W’ flag in conjunction with the ‘-n’ flag to see what would happen if you were to modify specific files.

With the ‘-n’ flag, make prints the commands that it would normally execute but does not execute them.

With the ‘-t’ flag, make ignores the commands in the rules and uses (in effect) the command touch for each target that needs to be remade. The touch command is also printed, unless ‘-s’ or .SILENT is used. For speed, make does not actually invoke the program touch. It does the work directly.

With the ‘-q’ flag, make prints nothing and executes no commands, but the exit status code it returns is zero if and only if the targets to be considered are already up to date.

It is an error to use more than one of these three flags in the same invocation of make.

The ‘-n’, ‘-t’, and ‘-q’ options do not affect command lines that begin with ‘+’ characters or contain the strings ‘$(MAKE)’ or ‘${MAKE}’. Note that only the line containing the ‘+’ character or the strings ‘$(MAKE)’ or ‘${MAKE}’ is run regardless of these options. Other lines in the same rule are not run unless they too begin with ‘+’ or contain ‘$(MAKE)’ or ‘${MAKE}’ (See Section 5.6.1 [How the MAKE Variable Works], page 56.)

The ‘-W’ flag provides two features:

• If you also use the ‘-n’ or ‘-q’ flag, you can see what make would do if you were to modify some files.

• Without the ‘-n’ or ‘-q’ flag, when make is actually executing commands, the ‘-W’ flag can direct make to act as if some files had been modified, without actually modifying the files.

Note that the options ‘-p’ and ‘-v’ allow you to obtain other information about make or about the makefiles in use (see Section 9.7 [Summary of Options], page 103).

9.4 Avoiding Recompilation of Some Files

Sometimes you may have changed a source file but you do not want to recompile all the files that depend on it. For example, suppose you add a macro or a declaration to a header file that many other files depend on. Being conservative, make assumes that any change in the header file requires recompilation of all dependent files, but you know that they do not need to be recompiled and you would rather not waste the time waiting for them to compile.

If you anticipate the problem before changing the header file, you can use the ‘-t’ flag. This flag tells make not to run the commands in the rules, but rather to mark the target up to date by changing its last-modification date. You would follow this procedure:

1. Use the command ‘make’ to recompile the source files that really need recompilation.

2. Make the changes in the header files.

3. Use the command ‘make -t’ to mark all the object files as up to date. The next time you run make, the changes in the header files will not cause any recompilation.

If you have already changed the header file at a time when some files do need recompilation, it is too late to do this. Instead, you can use the ‘-o file’ flag, which marks a specified file as “old” (see Section 9.7 [Summary of Options], page 103). This means that the file itself will not be remade, and nothing else will be remade on its account. Follow this procedure:

1. Recompile the source files that need compilation for reasons independent of the particular header file, with ‘make -o headerfile’. If several header files are involved, use a separate ‘-o’ option for each header file.

2. Touch all the object files with ‘make -t’.

9.5 Overriding Variables

An argument that contains ‘=’ specifies the value of a variable: ‘v=x’ sets the value of the variable v to x. If you specify a value in this way, all ordinary assignments of the same variable in the makefile are ignored; we say they have been overridden by the command line argument.

The most common way to use this facility is to pass extra flags to compilers. For example, in a properly written makefile, the variable CFLAGS is included in each command that runs the C compiler, so a file ‘foo.c’ would be compiled something like this:

     cc -c $(CFLAGS) foo.c

Thus, whatever value you set for CFLAGS affects each compilation that occurs. The makefile probably specifies the usual value for CFLAGS, like this:

     CFLAGS=-g

Each time you run make, you can override this value if you wish. For example, if you say ‘make CFLAGS='-g -O'’, each C compilation will be done with ‘cc -c -g -O’. (This illustrates how you can use quoting in the shell to enclose spaces and other special characters in the value of a variable when you override it.)

The variable CFLAGS is only one of many standard variables that exist just so that you can change them this way. See Section 10.3 [Variables Used by Implicit Rules], page 112, for a complete list.

You can also program the makefile to look at additional variables of your own, giving the user the ability to control other aspects of how the makefile works by changing the variables.

When you override a variable with a command argument, you can define either a recursivelyexpanded variable or a simply-expanded variable. The examples shown above make a recursivelyexpanded variable; to make a simply-expanded variable, write ‘:=’ instead of ‘=’. But, unless you want to include a variable reference or function call in the value that you specify, it makes no difference which kind of variable you create.

There is one way that the makefile can change a variable that you have overridden. This is to use the override directive, which is a line that looks like this: ‘override variable = value’ (see Section 6.7 [The override Directive], page 76).

9.6 Testing the Compilation of a Program

Normally, when an error happens in executing a shell command, make gives up immediately, returning a nonzero status. No further commands are executed for any target. The error implies that the goal cannot be correctly remade, and make reports this as soon as it knows.

When you are compiling a program that you have just changed, this is not what you want. Instead, you would rather that make try compiling every file that can be tried, to show you as many compilation errors as possible.

On these occasions, you should use the ‘-k’ or ‘--keep-going’ flag. This tells make to continue to consider the other dependencies of the pending targets, remaking them if necessary, before it gives up and returns nonzero status. For example, after an error in compiling one object file, ‘make -k’ will continue compiling other object files even though it already knows that linking them will be impossible. In addition to continuing after failed shell commands, ‘make -k’ will continue as much as possible after discovering that it does not know how to make a target or dependency file. This will always cause an error message, but without ‘-k’, it is a fatal error (see Section 9.7 [Summary of Options], page 103).

The usual behavior of make assumes that your purpose is to get the goals up to date; once make learns that this is impossible, it might as well report the failure immediately. The ‘-k’ flag says that the real purpose is to test as much as possible of the changes made in the program, perhaps to find several independent problems so that you can correct them all before the next attempt to compile. This is why Emacs’ M-x compile command passes the ‘-k’ flag by default.

9.7 Summary of Options

Here is a table of all the options make understands:

‘-b’

‘-m’ These options are ignored for compatibility with other versions of make.

‘-C dir’

‘--directory dir’

Change to directory dir before reading the makefiles. If multiple ‘-C’ options are specified, each is interpreted relative to the previous one: ‘-C / -C etc’ is equivalent to ‘-C /etc’. This is typically used with recursive invocations of make (see Section 5.6 [Recursive Use of make], page 56).

‘-d’
‘--debug’

Print debugging information in addition to normal processing. The debugging information says which files are being considered for remaking, which file-times are being compared and with what results, which files actually need to be remade, which implicit rules are considered and which are applied—everything interesting about how make decides what to do.

‘-e’
‘--environment-overrides’

Give variables taken from the environment precedence over variables from makefiles. See Section 6.9 [Variables from the Environment], page 78.

‘-f file’

‘--file file’
‘--makefile file’

Read the file named file as a makefile. See Chapter 3 [Writing Makefiles], page 23.

‘-h’
‘--help’

Remind you of the options that make understands and then exit.

‘-i’
‘--ignore-errors’

Ignore all errors in commands executed to remake files. See Section 5.4 [Errors in Commands], page 54.

‘-I dir’

‘--include-dir dir’

Specifies a directory dir to search for included makefiles. See Section 3.3 [Including Other Makefiles], page 24. If several ‘-I’ options are used to specify several directories, the directories are searched in the order specified.

‘-j [jobs]’

‘--jobs [jobs]’

Specifies the number of jobs (commands) to run simultaneously. With no argument, make runs as many jobs simultaneously as possible. If there is more than one ‘-j’ option, the last one is effective. See Section 5.3 [Parallel Execution], page 52, for more information on how commands are run.

‘-k’
‘--keep-going’

Continue as much as possible after an error. While the target that failed, and those that depend on it, cannot be remade, the other dependencies of these targets can be processed all the same. See Section 9.6 [Testing the Compilation of a Program], page 102.

‘-l [load]’

‘--load-average [load]’
‘--max-load [load]’

Specifies that no new jobs (commands) should be started if there are others jobs running and the load average is at least load (a floating-point number). With no argument, removes a previous load limit. See Section 5.3 [Parallel Execution], page 52.

‘-n’
‘--just-print’
‘--dry-run’
‘--recon’

Print the commands that would be executed, but do not execute them. See Section 9.3 [Instead of Executing the Commands], page 99.

‘-o file’

‘--old-file file’
‘--assume-old file’

Do not remake the file file even if it is older than its dependencies, and do not remake anything on account of changes in file. Essentially the file is treated as very old and its rules are ignored. See Section 9.4 [Avoiding Recompilation of Some Files], page 101.

‘-p’
‘--print-data-base’

Print the data base (rules and variable values) that results from reading the makefiles; then execute as usual or as otherwise specified. This also prints the version information given by the ‘-v’ switch (see below). To print the data base without trying to remake

          any files, use ‘make -p -f /dev/null’.
‘-q’
‘--question’

“Question mode”. Do not run any commands, or print anything; just return an exit status that is zero if the specified targets are already up to date, nonzero otherwise. See Section 9.3 [Instead of Executing the Commands], page 99.

‘-r’
‘--no-builtin-rules’

Eliminate use of the built-in implicit rules (see Chapter 10 [Using Implicit Rules], page 107). You can still define your own by writing pattern rules (see Section 10.5 [Defining and Redefining Pattern Rules], page 115). The ‘-r’ option also clears out the default list of suffixes for suffix rules (see Section 10.7 [Old-Fashioned Suffix Rules], page 123). But you can still define your own suffixes with a rule for .SUFFIXES, and then define your own suffix rules.

‘-s’
‘--silent’
‘--quiet’

Silent operation; do not print the commands as they are executed. See Section 5.1 [Command Echoing], page 51.

‘-S’
‘--no-keep-going’
‘--stop’

Cancel the effect of the ‘-k’ option. This is never necessary except in a recursive make where ‘-k’ might be inherited from the top-level make via MAKEFLAGS (see Section 5.6 [Recursive Use of make], page 56) or if you set ‘-k’ in MAKEFLAGS in your environment.

‘-t’
‘--touch’

Touch files (mark them up to date without really changing them) instead of running their commands. This is used to pretend that the commands were done, in order to fool future invocations of make. See Section 9.3 [Instead of Executing the Commands], page 99.

‘-v’
‘--version’

Print the version of the make program plus a copyright, a list of authors, and a notice that there is no warranty; then exit.

‘-w’
‘--print-directory’

Print a message containing the working directory both before and after executing the makefile. This may be useful for tracking down errors from complicated nests of recursive make commands. See Section 5.6 [Recursive Use of make], page 56. (In practice, you rarely need to specify this option since ‘make’ does it for you; see Section 5.6.4 [The ‘--print-directory’ Option], page 61.)

‘--no-print-directory’

Disable printing of the working directory under -w. This option is useful when - w is turned on automatically, but you do not want to see the extra messages. See Section 5.6.4 [The ‘--print-directory’ Option], page 61.

‘-W file’

‘--what-if file’
‘--new-file file’
‘--assume-new file’

Pretend that the target file has just been modified. When used with the ‘-n’ flag, this shows you what would happen if you were to modify that file. Without ‘-n’, it is almost the same as running a touch command on the given file before running make, except that the modification time is changed only in the imagination of make. See Section 9.3 [Instead of Executing the Commands], page 99.

10 Using Implicit Rules

Certain standard ways of remaking target files are used very often. For example, one customary way to make an object file is from a C source file using the C compiler, cc.

Implicit rules tell make how to use customary techniques so that you do not have to specify them in detail when you want to use them. For example, there is an implicit rule for C compilation. File names determine which implicit rules are run. For example, C compilation typically takes a ‘.c’ file and makes a ‘.o’ file. So make applies the implicit rule for C compilation when it sees this combination of file name endings.

A chain of implicit rules can apply in sequence; for example, make will remake a ‘.o’ file from a ‘.y’ file by way of a ‘.c’ file. See Section 10.4 [Chains of Implicit Rules], page 114.

The built-in implicit rules use several variables in their commands so that, by changing the values of the variables, you can change the way the implicit rule works. For example, the variable CFLAGS controls the flags given to the C compiler by the implicit rule for C compilation. See Section 10.3 [Variables Used by Implicit Rules], page 112.

You can define your own implicit rules by writing pattern rules. See Section 10.5 [Defining and Redefining Pattern Rules], page 115.

Suffix rules are a more limited way to define implicit rules. Pattern rules are more general and clearer, but suffix rules are retained for compatibility. See Section 10.7 [Old-Fashioned Suffix Rules], page 123.

10.1 Using Implicit Rules

To allow make to find a customary method for updating a target file, all you have to do is refrain from specifying commands yourself. Either write a rule with no command lines, or don’t write a rule at all. Then make will figure out which implicit rule to use based on which kind of source file exists or can be made.

For example, suppose the makefile looks like this:

     foo : foo.o bar.o
             cc -o foo foo.o bar.o $(CFLAGS) $(LDFLAGS)

Because you mention ‘foo.o’ but do not give a rule for it, make will automatically look for an implicit rule that tells how to update it. This happens whether or not the file ‘foo.o’ currently exists.

If an implicit rule is found, it can supply both commands and one or more dependencies (the source files). You would want to write a rule for ‘foo.o’ with no command lines if you need to specify additional dependencies, such as header files, that the implicit rule cannot supply.

Each implicit rule has a target pattern and dependency patterns. There may be many implicit rules with the same target pattern. For example, numerous rules make ‘.o’ files: one, from a ‘.c’ file with the C compiler; another, from a ‘.p’ file with the Pascal compiler; and so on. The rule that actually applies is the one whose dependencies exist or can be made. So, if you have a file ‘foo.c’, make will run the C compiler; otherwise, if you have a file ‘foo.p’, make will run the Pascal compiler; and so on.

Of course, when you write the makefile, you know which implicit rule you want make to use, and you know it will choose that one because you know which possible dependency files are supposed to exist. See Section 10.2 [Catalogue of Implicit Rules], page 109, for a catalogue of all the predefined implicit rules.

Above, we said an implicit rule applies if the required dependencies “exist or can be made”. A file “can be made” if it is mentioned explicitly in the makefile as a target or a dependency, or if an implicit rule can be recursively found for how to make it. When an implicit dependency is the result of another implicit rule, we say that chaining is occurring. See Section 10.4 [Chains of Implicit Rules], page 114.

In general, make searches for an implicit rule for each target, and for each double-colon rule, that has no commands. A file that is mentioned only as a dependency is considered a target whose rule specifies nothing, so implicit rule search happens for it. See Section 10.8 [Implicit Rule Search Algorithm], page 125, for the details of how the search is done.

Note that explicit dependencies do not influence implicit rule search. For example, consider this explicit rule:

     foo.o: foo.p

The dependency on ‘foo.p’ does not necessarily mean that make will remake ‘foo.o’ according to the implicit rule to make an object file, a ‘.o’ file, from a Pascal source file, a ‘.p’ file. For example, if ‘foo.c’ also exists, the implicit rule to make an object file from a C source file is used instead, because it appears before the Pascal rule in the list of predefined implicit rules (see Section 10.2 [Catalogue of Implicit Rules], page 109).

If you do not want an implicit rule to be used for a target that has no commands, you can give that target empty commands by writing a semicolon (see Section 5.8 [Defining Empty Commands], page 63).

10.2 Catalogue of Implicit Rules

Here is a catalogue of predefined implicit rules which are always available unless the makefile explicitly overrides or cancels them. See Section 10.5.6 [Canceling Implicit Rules], page 122, for information on canceling or overriding an implicit rule. The ‘-r’ or ‘--no-builtin-rules’ option cancels all predefined rules.

Not all of these rules will always be defined, even when the ‘-r’ option is not given. Many of the predefined implicit rules are implemented in make as suffix rules, so which ones will be defined depends on the suffix list (the list of dependencies of the special target .SUFFIXES). The default

suffix list is: .out, .a, .ln, .o, .c, .cc, .C, .p, .f, .F, .r, .y, .l, .s, .S, .mod, .sym, .def, .h,
.info, .dvi, .tex, .texinfo, .texi, .txinfo, .cweb, .web, .sh, .elc, .el. All of the implicit

rules described below whose dependencies have one of these suffixes are actually suffix rules. If you modify the suffix list, the only predefined suffix rules in effect will be those named by one or two of the suffixes that are on the list you specify; rules whose suffixes fail to be on the list are disabled. See Section 10.7 [Old-Fashioned Suffix Rules], page 123, for full details on suffix rules.

Compiling C programs ‘n.o’ is made automatically from ‘n.c’ with a command of the form ‘$(CC) -c

          $(CPPFLAGS) $(CFLAGS)’.

Compiling C++ programs ‘n.o’ is made automatically from ‘n.cc’ or ‘n.C’ with a command of the form ‘$(CXX) -c $(CPPFLAGS) $(CXXFLAGS)’. We encourage you to use the suffix ‘.cc’ for C++ source files instead of ‘.C’.

Compiling Pascal programs

‘n.o’ is made automatically from ‘n.p’ with the command ‘$(PC) -c $(PFLAGS)’.

Compiling Fortran and Ratfor programs ‘n.o’ is made automatically from ‘n.r’, ‘n.F’ or ‘n.f’ by running the Fortran compiler. The precise command used is as follows:

          ‘.f’  ‘$(FC) -c $(FFLAGS)’.

‘.F’  ‘$(FC) -c $(FFLAGS) $(CPPFLAGS)’.
‘.r’  ‘$(FC) -c $(FFLAGS) $(RFLAGS)’.

Preprocessing Fortran and Ratfor programs ‘n.f’ is made automatically from ‘n.r’ or ‘n.F’. This rule runs just the preprocessor to convert a Ratfor or preprocessable Fortran program into a strict Fortran program. The precise command used is as follows:

‘.F’  ‘$(FC) -F $(CPPFLAGS) $(FFLAGS)’.
‘.r’  ‘$(FC) -F $(FFLAGS) $(RFLAGS)’.

Compiling Modula-2 programs ‘n.sym’ is made from ‘n.def’ with a command of the form ‘$(M2C) $(M2FLAGS)

$(DEFFLAGS)’. ‘n.o’ is made from ‘n.mod’; the form is: ‘$(M2C) $(M2FLAGS) $(MODFLAGS)’.

Assembling and preprocessing assembler programs ‘n.o’ is made automatically from ‘n.s’ by running the assembler, as. The precise

command is ‘$(AS) $(ASFLAGS)’.

‘n.s’ is made automatically from ‘n.S’ by running the C preprocessor, cpp. The precise

command is ‘$(CPP) $(CPPFLAGS)’.

Linking a single object file ‘n’ is made automatically from ‘n.o’ by running the linker ld via the C compiler. The

precise command used is ‘$(CC) $(LDFLAGS) n.o $(LOADLIBES)’.

This rule does the right thing for a simple program with only one source file. It will also do the right thing if there are multiple object files (presumably coming from various other source files), one of which has a name matching that of the executable file. Thus,

     x: y.o z.o

when ‘x.c’, ‘y.c’ and ‘z.c’ all exist will execute:

     cc -c x.c -o x.o
     cc -c y.c -o y.o
     cc -c z.c -o z.o
     cc x.o y.o z.o -o x
     rm -f x.o
     rm -f y.o
     rm -f z.o

In more complicated cases, such as when there is no object file whose name derives from the executable file name, you must write an explicit command for linking. Each kind of file automatically made into ‘.o’ object files will be automatically linked by using the compiler (‘$(CC)’, ‘$(FC)’ or ‘$(PC)’; the C compiler ‘$(CC)’ is used to assemble ‘.s’ files) without the ‘-c’ option. This could be done by using the ‘.o’ object files as intermediates, but it is faster to do the compiling and linking in one step, so that’s how it’s done.

Yacc for C programs ‘n.c’ is made automatically from ‘n.y’ by running Yacc with the command ‘$(YACC)

$(YFLAGS)’.

Lex for C programs ‘n.c’ is made automatically from ‘n.l’ by by running Lex. The actual command is

‘$(LEX) $(LFLAGS)’.

Lex for Ratfor programs ‘n.r’ is made automatically from ‘n.l’ by by running Lex. The actual command is

‘$(LEX) $(LFLAGS)’.

The convention of using the same suffix ‘.l’ for all Lex files regardless of whether they produce C code or Ratfor code makes it impossible for make to determine automatically which of the two languages you are using in any particular case. If make is called upon to remake an object file from a ‘.l’ file, it must guess which compiler to use. It will guess the C compiler, because that is more common. If you are using Ratfor, make sure make knows this by mentioning ‘n.r’ in the makefile. Or, if you are using Ratfor exclusively, with no C files, remove ‘.c’ from the list of implicit rule suffixes with:

     .SUFFIXES:
     .SUFFIXES: .o .r .f .l . . .

Making Lint Libraries from C, Yacc, or Lex programs ‘n.ln’ is made from ‘n.c’ with a command of the form ‘$(LINT) $(LINTFLAGS) $(CPPFLAGS) -i’. The same command is used on the C code produced from ‘n.y’ or ‘n.l’.

TEX and Web ‘n.dvi’ is made from ‘n.tex’ with the command ‘$(TEX)’. ‘n.tex’ is made from ‘n.web’ with ‘$(WEAVE)’, or from ‘n.cweb’ with ‘$(CWEAVE)’. ‘n.p’ is made from ‘n.web’ with ‘$(TANGLE)’ and ‘n.c’ is made from ‘n.cweb’ with ‘$(CTANGLE)’.

Texinfo and Info ‘n.dvi’ is made from ‘n.texinfo’, ‘n.texi’, or ‘n.txinfo’, with the ‘$(TEXI2DVI)’ command. ‘n.info’ is made from ‘n.texinfo’, ‘n.texi’, or ‘n.txinfo’, with the

‘$(MAKEINFO)’ command.

RCS Any file ‘n’ is extracted if necessary from an RCS file named either ‘n,v’ or ‘RCS/n,v’. The precise command used is ‘$(CO) $(COFLAGS)’. ‘n’ will not be extracted from RCS if it already exists, even if the RCS file is newer. The rules for RCS are terminal (see Section 10.5.5 [Match-Anything Pattern Rules], page 120), so RCS files cannot be generated from another source; they must actually exist.

SCCS Any file ‘n’ is extracted if necessary from an SCCS file named either ‘s.n’ or ‘SCCS/s.n’. The precise command used is ‘$(GET) $(GFLAGS)’. The rules for SCCS are terminal (see Section 10.5.5 [Match-Anything Pattern Rules], page 120), so SCCS files cannot be generated from another source; they must actually exist.

For the benefit of SCCS, a file ‘n’ is copied from ‘n.sh’ and made executable (by everyone). This is for shell scripts that are checked into SCCS. Since RCS preserves the execution permission of a file, you do not need to use this feature with RCS. We recommend that you avoid using of SCCS. RCS is widely held to be superior, and is also free. By choosing free software in place of comparable (or inferior) proprietary software, you support the free software movement.

Usually, you want to change only the variables listed in the table above, which are documented in the following section.

However, the commands in built-in implicit rules actually use variables such as COMPILE.c, LINK.p, and PREPROCESS.S, whose values contain the commands listed above.

make follows the convention that the rule to compile a ‘.x’ source file uses the variable COMPILE.x. Similarly, the rule to produce an executable from a ‘.x’ file uses LINK.x; and the rule to preprocess a ‘.x’ file uses PREPROCESS.x.

Every rule that produces an object file uses the variable OUTPUT_OPTION. make defines this variable either to contain ‘-o $@’, or to be empty, depending on a compile-time option. You need the ‘-o’ option to ensure that the output goes into the right file when the source file is in a different directory, as when using VPATH (see Section 4.3 [Directory Search], page 33). However, compilers on some systems do not accept a ‘-o’ switch for object files. If you use such a system, and use VPATH, some compilations will put their output in the wrong place. A possible workaround for this problem is to give OUTPUT_OPTION the value ‘; mv $*.o $@’.

10.3 Variables Used by Implicit Rules

The commands in built-in implicit rules make liberal use of certain predefined variables. You can alter these variables in the makefile, with arguments to make, or in the environment to alter how the implicit rules work without redefining the rules themselves.

For example, the command used to compile a C source file actually says ‘$(CC) -c $(CFLAGS) $(CPPFLAGS)’. The default values of the variables used are ‘cc’ and nothing, resulting in the command ‘cc -c’. By redefining ‘CC’ to ‘ncc’, you could cause ‘ncc’ to be used for all C compilations performed by the implicit rule. By redefining ‘CFLAGS’ to be ‘-g’, you could pass the ‘-g’ option to each compilation. All implicit rules that do C compilation use ‘$(CC)’ to get the program name for the compiler and all include ‘$(CFLAGS)’ among the arguments given to the compiler.

The variables used in implicit rules fall into two classes: those that are names of programs (like CC) and those that contain arguments for the programs (like CFLAGS). (The “name of a program” may also contain some command arguments, but it must start with an actual executable program name.) If a variable value contains more than one argument, separate them with spaces.

Here is a table of variables used as names of programs in built-in rules:

AR Archive-maintaining program; default ‘ar’. AS Program for doing assembly; default ‘as’. CC Program for compiling C programs; default ‘cc’. CXX Program for compiling C++ programs; default ‘g++’. CO Program for extracting a file from RCS; default ‘co’. CPP Program for running the C preprocessor, with results to standard output; default ‘$(CC)

          -E’.

FC Program for compiling or preprocessing Fortran and Ratfor programs; default ‘f77’. GET Program for extracting a file from SCCS; default ‘get’. LEX Program to use to turn Lex grammars into C programs or Ratfor programs; default

          ‘lex’.

PC Program for compiling Pascal programs; default ‘pc’. YACC Program to use to turn Yacc grammars into C programs; default ‘yacc’. YACCR Program to use to turn Yacc grammars into Ratfor programs; default ‘yacc -r’. MAKEINFO Program to convert a Texinfo source file into an Info file; default ‘makeinfo’. TEX Program to make TEX dvi files from TEX source; default ‘tex’. TEXI2DVI Program to make TEX dvi files from Texinfo source; default ‘texi2dvi’. WEAVE Program to translate Web into TEX; default ‘weave’. CWEAVE Program to translate C Web into TEX; default ‘cweave’. TANGLE Program to translate Web into Pascal; default ‘tangle’. CTANGLE Program to translate C Web into C; default ‘ctangle’. RM Command to remove a file; default ‘rm -f’.

Here is a table of variables whose values are additional arguments for the programs above. The default values for all of these is the empty string, unless otherwise noted.

ARFLAGS Flags to give the archive-maintaining program; default ‘rv’. ASFLAGS Extra flags to give to the assembler (when explicitly invoked on a ‘.s’ or ‘.S’ file). CFLAGS Extra flags to give to the C compiler. CXXFLAGS Extra flags to give to the C++ compiler. COFLAGS Extra flags to give to the RCS co program. CPPFLAGS Extra flags to give to the C preprocessor and programs that use it (the C and Fortran

compilers).

FFLAGS Extra flags to give to the Fortran compiler. GFLAGS Extra flags to give to the SCCS get program. LDFLAGS Extra flags to give to compilers when they are supposed to invoke the linker, ‘ld’. LFLAGS Extra flags to give to Lex. PFLAGS Extra flags to give to the Pascal compiler. RFLAGS Extra flags to give to the Fortran compiler for Ratfor programs. YFLAGS Extra flags to give to Yacc.

10.4 Chains of Implicit Rules

Sometimes a file can be made by a sequence of implicit rules. For example, a file ‘n.o’ could be made from ‘n.y’ by running first Yacc and then cc. Such a sequence is called a chain.

If the file ‘n.c’ exists, or is mentioned in the makefile, no special searching is required: make finds that the object file can be made by C compilation from ‘n.c’; later on, when considering how to make ‘n.c’, the rule for running Yacc is used. Ultimately both ‘n.c’ and ‘n.o’ are updated.

However, even if ‘n.c’ does not exist and is not mentioned, make knows how to envision it as the missing link between ‘n.o’ and ‘n.y’! In this case, ‘n.c’ is called an intermediate file. Once make has decided to use the intermediate file, it is entered in the data base as if it had been mentioned in the makefile, along with the implicit rule that says how to create it.

Intermediate files are remade using their rules just like all other files. The difference is that the intermediate file is deleted when make is finished. Therefore, the intermediate file which did not exist before make also does not exist after make. The deletion is reported to you by printing a ‘rm -f’ command that shows what make is doing. (You can list the target pattern of an implicit rule (such as ‘%.o’) as a dependency of the special target .PRECIOUS to preserve intermediate files made by implicit rules whose target patterns match that file’s name; see Section 5.5 [Interrupts], page 55.)

A chain can involve more than two implicit rules. For example, it is possible to make a file ‘foo’ from ‘RCS/foo.y,v’ by running RCS, Yacc and cc. Then both ‘foo.y’ and ‘foo.c’ are intermediate files that are deleted at the end.

No single implicit rule can appear more than once in a chain. This means that make will not even consider such a ridiculous thing as making ‘foo’ from ‘foo.o.o’ by running the linker twice. This constraint has the added benefit of preventing any infinite loop in the search for an implicit rule chain.

There are some special implicit rules to optimize certain cases that would otherwise by handled by rule chains. For example, making ‘foo’ from ‘foo.c’ could be handled by compiling and linking with separate chained rules, using ‘foo.o’ as an intermediate file. But what actually happens is that a special rule for this case does the compilation and linking with a single cc command. The optimized rule is used in preference to the step-by-step chain because it comes earlier in the ordering of rules.

10.5 Defining and Redefining Pattern Rules

You define an implicit rule by writing a pattern rule. A pattern rule looks like an ordinary rule, except that its target contains the character ‘%’ (exactly one of them). The target is considered a pattern for matching file names; the ‘%’ can match any nonempty substring, while other characters match only themselves. The dependencies likewise use ‘%’ to show how their names relate to the target name.

Thus, a pattern rule ‘%.o : %.c’ says how to make any file ‘stem.o’ from another file ‘stem.c’.

Note that expansion using ‘%’ in pattern rules occurs after any variable or function expansions, which take place when the makefile is read. See Chapter 6 [How to Use Variables], page 65, and Chapter 8 [Functions for Transforming Text], page 85.

10.5.1 Introduction to Pattern Rules

A pattern rule contains the character ‘%’ (exactly one of them) in the target; otherwise, it looks exactly like an ordinary rule. The target is a pattern for matching file names; the ‘%’ matches any nonempty substring, while other characters match only themselves.

For example, ‘%.c’ as a pattern matches any file name that ends in ‘.c’. ‘s.%.c’ as a pattern matches any file name that starts with ‘s.’, ends in ‘.c’ and is at least five characters long. (There must be at least one character to match the ‘%’.) The substring that the ‘%’ matches is called the stem.

‘%’ in a dependency of a pattern rule stands for the same stem that was matched by the ‘%’ in the target. In order for the pattern rule to apply, its target pattern must match the file name under consideration, and its dependency patterns must name files that exist or can be made. These files become dependencies of the target.

Thus, a rule of the form

%.o : %.c ; command…

specifies how to make a file ‘n.o’, with another file ‘n.c’ as its dependency, provided that ‘n.c’ exists or can be made.

There may also be dependencies that do not use ‘%’; such a dependency attaches to every file made by this pattern rule. These unvarying dependencies are useful occasionally.

A pattern rule need not have any dependencies that contain ‘%’, or in fact any dependencies at all. Such a rule is effectively a general wildcard. It provides a way to make any file that matches the target pattern. See Section 10.6 [Last Resort], page 122.

Pattern rules may have more than one target. Unlike normal rules, this does not act as many different rules with the same dependencies and commands. If a pattern rule has multiple targets, make knows that the rule’s commands are responsible for making all of the targets. The commands are executed only once to make all the targets. When searching for a pattern rule to match a target, the target patterns of a rule other than the one that matches the target in need of a rule are incidental: make worries only about giving commands and dependencies to the file presently in question. However, when this file’s commands are run, the other targets are marked as having been updated themselves.

The order in which pattern rules appear in the makefile is important since this is the order in which they are considered. Of equally applicable rules, only the first one found is used. The rules you write take precedence over those that are built in. Note however, that a rule whose dependencies actually exist or are mentioned always takes priority over a rule with dependencies that must be made by chaining other implicit rules.

10.5.2 Pattern Rule Examples

Here are some examples of pattern rules actually predefined in make. First, the rule that compiles ‘.c’ files into ‘.o’ files:

     %.o : %.c
             $(CC) -c $(CFLAGS) $(CPPFLAGS) $< -o $@

defines a rule that can make any file ‘x.o’ from ‘x.c’. The command uses the automatic variables ‘$@’ and ‘$<’ to substitute the names of the target file and the source file in each case where the rule applies (see Section 10.5.3 [Automatic Variables], page 118).

Here is a second built-in rule:

     % :: RCS/%,v
             $(CO) $(COFLAGS) $<

defines a rule that can make any file ‘x’ whatsoever from a corresponding file ‘x,v’ in the subdirectory ‘RCS’. Since the target is ‘%’, this rule will apply to any file whatever, provided the appropriate dependency file exists. The double colon makes the rule terminal, which means that its dependency may not be an intermediate file (see Section 10.5.5 [Match-Anything Pattern Rules], page 120).

This pattern rule has two targets:

     %.tab.c %.tab.h: %.y
             bison -d $<

This tells make that the command ‘bison -d x.y’ will make both ‘x.tab.c’ and ‘x.tab.h’. If the file ‘foo’ depends on the files ‘parse.tab.o’ and ‘scan.o’ and the file ‘scan.o’ depends on the file ‘parse.tab.h’, when ‘parse.y’ is changed, the command ‘bison -d parse.y’ will be executed only once, and the dependencies of both ‘parse.tab.o’ and ‘scan.o’ will be satisfied. (Presumably the file ‘parse.tab.o’ will be recompiled from ‘parse.tab.c’ and the file ‘scan.o’ from ‘scan.c’, while ‘foo’ is linked from ‘parse.tab.o’, ‘scan.o’, and its other dependencies, and it will execute happily ever after.)

10.5.3 Automatic Variables

Suppose you are writing a pattern rule to compile a ‘.c’ file into a ‘.o’ file: how do you write the ‘cc’ command so that it operates on the right source file name? You cannot write the name in the command, because the name is different each time the implicit rule is applied.

What you do is use a special feature of make, the automatic variables. These variables have values computed afresh for each rule that is executed, based on the target and dependencies of the rule. In this example, you would use ‘$@’ for the object file name and ‘$<’ for the source file name.

Here is a table of automatic variables:

$@ The file name of the target of the rule. If the target is an archive member, then ‘$@’ is the name of the archive file. In a pattern rule that has multiple targets (see Section 10.5.1 [Introduction to Pattern Rules], page 116), ‘$@’ is the name of whichever target caused the rule’s commands to be run.

$% The target member name, when the target is an archive member. See Chapter 11 [Archives], page 127. For example, if the target is ‘foo.a(bar.o)’ then ‘$%’ is ‘bar.o’ and ‘$@’ is ‘foo.a’. ‘$%’ is empty when the target is not an archive member.

$< The name of the first dependency. If the target got its commands from an implicit rule, this will be the first dependency added by the implicit rule (see Chapter 10 [Implicit Rules], page 107).

$? The names of all the dependencies that are newer than the target, with spaces between them. For dependencies which are archive members, only the member named is used (see Chapter 11 [Archives], page 127).

$^The names of all the dependencies, with spaces between them. For dependencies which are archive members, only the member named is used (see Chapter 11 [Archives], page 127).

$* The stem with which an implicit rule matches (see Section 10.5.4 [How Patterns Match], page 120). If the target is ‘dir/a.foo.b’ and the target pattern is ‘a.%.b’ then the stem is ‘dir/foo’. The stem is useful for constructing names of related files. In a static pattern rule, the stem is part of the file name that matched the ‘%’ in the target pattern.

In an explicit rule, there is no stem; so ‘$*’ cannot be determined in that way. Instead, if the target name ends with a recognized suffix (see Section 10.7 [Old-Fashioned Suffix Rules], page 123), ‘$*’ is set to the target name minus the suffix. For example, if the target name is ‘foo.c’, then ‘$*’ is set to ‘foo’, since ‘.c’ is a suffix. GNU make does this bizarre thing only for compatibility with other implementations of make. You should generally never use ‘$*’ except in implicit rules or static pattern rules. If the target name in an explicit rule does not end with a recognized suffix, ‘$*’ is set to the empty string for that rule.

‘$?’ is useful even in explicit rules when you wish to operate on only the dependencies that have changed. For example, suppose that an archive named ‘lib’ is supposed to contain copies of several object files. This rule copies just the changed object files into the archive:

     lib: foo.o bar.o lose.o win.o
             ar r lib $?

Of the variables listed above, four have values that are single file names, and two have values that are lists of file names. These six have variants that get just the file’s directory name or just the file name within the directory. The variant variables’ names are formed by appending ‘D’ or ‘F’, respectively. These variants are semi-obsolete in GNU make since the functions dir and notdir can be used to get an equivalent effect (see Section 8.3 [Functions for File Names], page 89). Here is a table of the variants:

‘$(@D)’ The directory part of the file name of the target. If the value of ‘$@’ is ‘dir/foo.o’ then ‘$(@D)’ is ‘dir/’. This value is ‘./’ if ‘$@’ does not contain a slash. ‘$(@D)’ is equivalent to ‘$(dir $@)’.

‘$(@F)’ The file-within-directory part of the file name of the target. If the value of ‘$@’ is

          ‘dir/foo.o’ then ‘$(@F)’ is ‘foo.o’. ‘$(@F)’ is equivalent to ‘$(notdir $@)’.
‘$(*D)’

‘$(*F)’ The directory part and the file-within-directory part of the stem; ‘dir/’ and ‘foo’ in this example.

‘$(%D)’

‘$(%F)’ The directory part and the file-within-directory part of the target archive member name. This makes sense only for archive member targets of the form ‘archive(member)’ and is useful only when member may contain a directory name. (See Section 11.1 [Archive Members as Targets], page 127.)

‘$(<D)’

‘$(<F)’ The directory part and the file-within-directory part of the first dependency.

‘$(^D)’

‘$(^F)’ Lists of the directory parts and the file-within-directory parts of all dependencies.

‘$(?D)’

‘$(?F)’ Lists of the directory parts and the file-within-directory parts of all dependencies that are newer than the target.

Note that we use a special stylistic convention when we talk about these automatic variables; we write “the value of ‘$<’”, rather than “the variable <” as we would write for ordinary variables such as objects and CFLAGS. We think this convention looks more natural in this special case. Please do not assume it has a deep significance; ‘$<’ refers to the variable named < just as ‘$(CFLAGS)’ refers to the variable named CFLAGS. You could just as well use ‘$(<)’ in place of ‘$<’.

10.5.4 How Patterns Match

A target pattern is composed of a ‘%’ between a prefix and a suffix, either or both of which may be empty. The pattern matches a file name only if the file name starts with the prefix and ends with the suffix, without overlap. The text between the prefix and the suffix is called the stem. Thus, when the pattern ‘%.o’ matches the file name ‘test.o’, the stem is ‘test’. The pattern rule dependencies are turned into actual file names by substituting the stem for the character ‘%’. Thus, if in the same example one of the dependencies is written as ‘%.c’, it expands to ‘test.c’.

When the target pattern does not contain a slash (and it usually does not), directory names in the file names are removed from the file name before it is compared with the target prefix and suffix. After the comparison of the file name to the target pattern, the directory names, along with the slash that ends them, are added on to the dependency file names generated from the pattern rule’s dependency patterns and the file name. The directories are ignored only for the purpose of finding an implicit rule to use, not in the application of that rule. Thus, ‘e%t’ matches the file name ‘src/eat’, with ‘src/a’ as the stem. When dependencies are turned into file names, the directories from the stem are added at the front, while the rest of the stem is substituted for the ‘%’. The stem ‘src/a’ with a dependency pattern ‘c%r’ gives the file name ‘src/car’.

10.5.5 Match-Anything Pattern Rules

When a pattern rule’s target is just ‘%’, it matches any file name whatever. We call these rules match-anything rules. They are very useful, but it can take a lot of time for make to think about them, because it must consider every such rule for each file name listed either as a target or as a dependency.

Suppose the makefile mentions ‘foo.c’. For this target, make would have to consider making it by linking an object file ‘foo.c.o’, or by C compilation-and-linking in one step from ‘foo.c.c’, or by Pascal compilation-and-linking from ‘foo.c.p’, and many other possibilities.

We know these possibilities are ridiculous since ‘foo.c’ is a C source file, not an executable. If make did consider these possibilities, it would ultimately reject them, because files such as ‘foo.c.o’ and ‘foo.c.p’ would not exist. But these possibilities are so numerous that make would run very slowly if it had to consider them.

To gain speed, we have put various constraints on the way make considers match-anything rules. There are two different constraints that can be applied, and each time you define a match-anything rule you must choose one or the other for that rule.

One choice is to mark the match-anything rule as terminal by defining it with a double colon. When a rule is terminal, it does not apply unless its dependencies actually exist. Dependencies that could be made with other implicit rules are not good enough. In other words, no further chaining is allowed beyond a terminal rule.

For example, the built-in implicit rules for extracting sources from RCS and SCCS files are terminal; as a result, if the file ‘foo.c,v’ does not exist, make will not even consider trying to make it as an intermediate file from ‘foo.c,v.o’ or from ‘RCS/SCCS/s.foo.c,v’. RCS and SCCS files are generally ultimate source files, which should not be remade from any other files; therefore, make can save time by not looking for ways to remake them.

If you do not mark the match-anything rule as terminal, then it is nonterminal. A nonterminal match-anything rule cannot apply to a file name that indicates a specific type of data. A file name indicates a specific type of data if some non-match-anything implicit rule target matches it.

For example, the file name ‘foo.c’ matches the target for the pattern rule ‘%.c : %.y’ (the rule to run Yacc). Regardless of whether this rule is actually applicable (which happens only if there is a file ‘foo.y’), the fact that its target matches is enough to prevent consideration of any nonterminal match-anything rules for the file ‘foo.c’. Thus, make will not even consider trying to make ‘foo.c’ as an executable file from ‘foo.c.o’, ‘foo.c.c’, ‘foo.c.p’, etc.

The motivation for this constraint is that nonterminal match-anything rules are used for making files containing specific types of data (such as executable files) and a file name with a recognized suffix indicates some other specific type of data (such as a C source file).

Special built-in dummy pattern rules are provided solely to recognize certain file names so that nonterminal match-anything rules will not be considered. These dummy rules have no dependencies and no commands, and they are ignored for all other purposes. For example, the built-in implicit rule

     %.p :

exists to make sure that Pascal source files such as ‘foo.p’ match a specific target pattern and thereby prevent time from being wasted looking for ‘foo.p.o’ or ‘foo.p.c’.

Dummy pattern rules such as the one for ‘%.p’ are made for every suffix listed as valid for use in suffix rules (see Section 10.7 [Old-Fashioned Suffix Rules], page 123).

10.5.6 Canceling Implicit Rules

You can override a built-in implicit rule (or one you have defined yourself) by defining a new pattern rule with the same target and dependencies, but different commands. When the new rule is defined, the built-in one is replaced. The new rule’s position in the sequence of implicit rules is determined by where you write the new rule.

You can cancel a built-in implicit rule by defining a pattern rule with the same target and dependencies, but no commands. For example, the following would cancel the rule that runs the assembler:

     %.o : %.s

10.6 Defining Last-Resort Default Rules

You can define a last-resort implicit rule by writing a terminal match-anything pattern rule with no dependencies (see Section 10.5.5 [Match-Anything Rules], page 120). This is just like any other pattern rule; the only thing special about it is that it will match any target. So such a rule’s commands are used for all targets and dependencies that have no commands of their own and for which no other implicit rule applies.

For example, when testing a makefile, you might not care if the source files contain real data, only that they exist. Then you might do this:

     %::
             touch $@

to cause all the source files needed (as dependencies) to be created automatically.

You can instead define commands to be used for targets for which there are no rules at all, even ones which don’t specify commands. You do this by writing a rule for the target .DEFAULT. Such a rule’s commands are used for all dependencies which do not appear as targets in any explicit rule, and for which no implicit rule applies. Naturally, there is no .DEFAULT rule unless you write one.

If you use .DEFAULT with no commands or dependencies:

     .DEFAULT:

the commands previously stored for .DEFAULT are cleared. Then make acts as if you had never defined .DEFAULT at all.

If you do not want a target to get the commands from a match-anything pattern rule or .DEFAULT, but you also do not want any commands to be run for the target, you can give it empty commands (see Section 5.8 [Defining Empty Commands], page 63).

You can use a last-resort rule to override part of another makefile. See Section 3.6 [Overriding Part of Another Makefile], page 27.

10.7 Old-Fashioned Suffix Rules

Suffix rules are the old-fashioned way of defining implicit rules for make. Suffix rules are obsolete because pattern rules are more general and clearer. They are supported in GNU make for compatibility with old makefiles. They come in two kinds: double-suffix and single-suffix.

A double-suffix rule is defined by a pair of suffixes: the target suffix and the source suffix. It matches any file whose name ends with the target suffix. The corresponding implicit dependency is made by replacing the target suffix with the source suffix in the file name. A two-suffix rule whose target and source suffixes are ‘.o’ and ‘.c’ is equivalent to the pattern rule ‘%.o : %.c’.

A single-suffix rule is defined by a single suffix, which is the source suffix. It matches any file name, and the corresponding implicit dependency name is made by appending the source suffix. A single-suffix rule whose source suffix is ‘.c’ is equivalent to the pattern rule ‘% : %.c’.

Suffix rule definitions are recognized by comparing each rule’s target against a defined list of known suffixes. When make sees a rule whose target is a known suffix, this rule is considered a single-suffix rule. When make sees a rule whose target is two known suffixes concatenated, this rule is taken as a double-suffix rule.

For example, ‘.c’ and ‘.o’ are both on the default list of known suffixes. Therefore, if you define a rule whose target is ‘.c.o’, make takes it to be a double-suffix rule with source suffix ‘.c’ and target suffix ‘.o’. Here is the old-fashioned way to define the rule for compiling a C source file:

     .c.o:
             $(CC) -c $(CFLAGS) $(CPPFLAGS) -o $@ $<

Suffix rules cannot have any dependencies of their own. If they have any, they are treated as normal files with funny names, not as suffix rules. Thus, the rule:

     .c.o: foo.h
             $(CC) -c $(CFLAGS) $(CPPFLAGS) -o $@ $<

tells how to make the file ‘.c.o’ from the dependency file ‘foo.h’, and is not at all like the pattern rule:

     %.o: %.c foo.h
             $(CC) -c $(CFLAGS) $(CPPFLAGS) -o $@ $<

which tells how to make ‘.o’ files from ‘.c’ files, and makes all ‘.o’ files using this pattern rule also depend on ‘foo.h’.

Suffix rules with no commands are also meaningless. They do not remove previous rules as do pattern rules with no commands (see Section 10.5.6 [Canceling Implicit Rules], page 122). They simply enter the suffix or pair of suffixes concatenated as a target in the data base.

The known suffixes are simply the names of the dependencies of the special target .SUFFIXES. You can add your own suffixes by writing a rule for .SUFFIXES that adds more dependencies, as in:

     .SUFFIXES: .hack .win

which adds ‘.hack’ and ‘.win’ to the end of the list of suffixes.

If you wish to eliminate the default known suffixes instead of just adding to them, write a rule for .SUFFIXES with no dependencies. By special dispensation, this eliminates all existing dependencies of .SUFFIXES. You can then write another rule to add the suffixes you want. For example,

.SUFFIXES: # Delete the default suffixes

     .SUFFIXES: .c .o .h   # Define our suffix list

The ‘-r’ or ‘--no-builtin-rules’ flag causes the default list of suffixes to be empty.

The variable SUFFIXES is defined to the default list of suffixes before make reads any makefiles. You can change the list of suffixes with a rule for the special target .SUFFIXES, but that does not alter this variable.

10.8 Implicit Rule Search Algorithm

Here is the procedure make uses for searching for an implicit rule for a target t. This procedure is followed for each double-colon rule with no commands, for each target of ordinary rules none of which have commands, and for each dependency that is not the target of any rule. It is also followed recursively for dependencies that come from implicit rules, in the search for a chain of rules.

Suffix rules are not mentioned in this algorithm because suffix rules are converted to equivalent pattern rules once the makefiles have been read in.

For an archive member target of the form ‘archive(member)’, the following algorithm is run twice, first using ‘(member)’ as the target t, and second using the entire target if the first run found no rule.

1. Split t into a directory part, called d, and the rest, called n. For example, if t is ‘src/foo.o’, then d is ‘src/’ and n is ‘foo.o’.

2. Make a list of all the pattern rules one of whose targets matches t or n. If the target pattern contains a slash, it is matched against t; otherwise, against n.

3. If any rule in that list is not a match-anything rule, then remove all nonterminal match-anything rules from the list.

4. Remove from the list all rules with no commands.

5. For each pattern rule in the list:

1. Find the stem s, which is the nonempty part of t or n matched by the ‘%’ in the target pattern.

2. Compute the dependency names by substituting s for ‘%’; if the target pattern does not contain a slash, append d to the front of each dependency name.

3. Test whether all the dependencies exist or ought to exist. (If a file name is mentioned in the makefile as a target or as an explicit dependency, then we say it ought to exist.) If all dependencies exist or ought to exist, or there are no dependencies, then this rule applies.

6. If no pattern rule has been found so far, try harder. For each pattern rule in the list: 1. If the rule is terminal, ignore it and go on to the next rule.

2. Compute the dependency names as before.

3. Test whether all the dependencies exist or ought to exist.

4. For each dependency that does not exist, follow this algorithm recursively to see if the dependency can be made by an implicit rule.

5. If all dependencies exist, ought to exist, or can be made by implicit rules, then this rule

applies.

7. If no implicit rule applies, the rule for .DEFAULT, if any, applies. In that case, give t the same commands that .DEFAULT has. Otherwise, there are no commands for t.

Once a rule that applies has been found, for each target pattern of the rule other than the one that matched t or n, the ‘%’ in the pattern is replaced with s and the resultant file name is stored until the commands to remake the target file t are executed. After these commands are executed, each of these stored file names are entered into the data base and marked as having been updated and having the same update status as the file t.

When the commands of a pattern rule are executed for t, the automatic variables are set corresponding to the target and dependencies. See Section 10.5.3 [Automatic Variables], page 118.

11 Using make to Update Archive Files

Archive files are files containing named subfiles called members; they are maintained with the program ar and their main use is as subroutine libraries for linking.

11.1 Archive Members as Targets

An individual member of an archive file can be used as a target or dependency in make. The archive file must already exist, but the member need not exist. You specify the member named member in archive file archive as follows:

archive(member)

This construct is available only in targets and dependencies, not in commands! Most programs that you might use in commands do not support this syntax and cannot act directly on archive members. Only ar and other programs specifically designed to operate on archives can do so. Therefore, valid commands to update an archive member target probably must use ar. For example, this rule says to create a member ‘hack.o’ in archive ‘foolib’ by copying the file ‘hack.o’:

     foolib(hack.o) : hack.o
             ar r foolib hack.o

In fact, nearly all archive member targets are updated in just this way and there is an implicit rule to do it for you.

11.2 Implicit Rule for Archive Member Targets

Recall that a target that looks like ‘a(m)’ stands for the member named m in the archive file a.

When make looks for an implicit rule for such a target, as a special feature it considers implicit rules that match ‘(m)’, as well as those that match the actual target ‘a(m)’.

This causes one special rule whose target is ‘(%)’ to match. This rule updates the target ‘a(m)’ by copying the file m into the archive. For example, it will update the archive member target ‘foo.a(bar.o)’ by copying the file ‘bar.o’ into the archive ‘foo.a’ as a member named ‘bar.o’.

When this rule is chained with others, the result is very powerful. Thus, ‘make "foo.a(bar.o)"’ (the quotes are needed to protect the ‘(’ and ‘)’ from being interpreted specially by the shell) in the presence of a file ‘bar.c’ is enough to cause the following commands to be run, even without a makefile:

     cc -c bar.c -o bar.o
     ar r foo.a bar.o
     rm -f bar.o

Here make has envisioned the file ‘bar.o’ as an intermediate file. See Section 10.4 [Chains of Implicit Rules], page 114.

Implicit rules such as this one are written using the automatic variable ‘$%’. See Section 10.5.3 [Automatic Variables], page 118.

An archive member name in an archive cannot contain a directory name, but it may be useful in a makefile to pretend that it does. If you write an archive member target ‘foo.a(dir/file.o)’, make will perform automatic updating with this command:

     ar r foo.a dir/file.o

which has the effect of copying the file ‘dir/foo.o’ into a member named ‘foo.o’. In connection with such usage, the automatic variables %D and %F may be useful.

11.2.1 Updating Archive Symbol Directories

An archive file that is used as a library usually contains a special member named ‘__.SYMDEF’ that contains a directory of the external symbol names defined by all the other members. After you update any other members, you need to update ‘__.SYMDEF’ so that it will summarize the other members properly. This is done by running the ranlib program:

ranlib archivefile

Normally you would put this command in the rule for the archive file, and make all the members of the archive file dependencies of that rule. For example,

     libfoo.a: libfoo.a(x.o) libfoo.a(y.o) . . .
             ranlib libfoo.a

The effect of this is to update archive members ‘x.o’, ‘y.o’, etc., and then update the symbol directory member ‘__.SYMDEF’ by running ranlib. The rules for updating the members are not shown here; most likely you can omit them and use the implicit rule which copies files into the archive, as described in the preceding section.

This is not necessary when using the GNU ar program, which updates the ‘__.SYMDEF’ member automatically.

11.3 Suffix Rules for Archive Files

You can write a special kind of suffix rule for dealing with archive files. See Section 10.7 [Suffix Rules], page 123, for a full explanation of suffix rules. Archive suffix rules are obsolete in GNU make, because pattern rules for archives are a more general mechanism (see Section 11.2 [Archive Update], page 127). But they are retained for compatibility with other makes.

To write a suffix rule for archives, you simply write a suffix rule using the target suffix ‘.a’ (the usual suffix for archive files). For example, here is the old-fashioned suffix rule to update a library archive from C source files:

     .c.a:
             $(CC) $(CFLAGS) $(CPPFLAGS) -c $< -o $*.o
             $(AR) r $@ $*.o
             $(RM) $*.o

This works just as if you had written the pattern rule:

     (%.o): %.c
             $(CC) $(CFLAGS) $(CPPFLAGS) -c $< -o $*.o
             $(AR) r $@ $*.o
             $(RM) $*.o

In fact, this is just what make does when it sees a suffix rule with ‘.a’ as the target suffix. Any double-suffix rule ‘.x.a’ is converted to a pattern rule with the target pattern ‘(%.o)’ and a dependency pattern of ‘%.x’.

130 GNU make

12 Features of GNU make

Here is a summary of the features of GNU make, for comparison with and credit to other versions of make. We consider the features of make in 4.2 BSD systems as a baseline. If you are concerned with writing portable makefiles, you should use only the features of make not listed here or in Chapter 13 [Missing], page 135.

Many features come from the version of make in System V.

• The VPATH variable and its special meaning. See Section 4.3 [Searching Directories for Dependencies], page 33. This feature exists in System V make, but is undocumented. It is documented in 4.3 BSD make (which says it mimics System V’s VPATH feature).

• Included makefiles. See Section 3.3 [Including Other Makefiles], page 24. Allowing multiple files to be included with a single directive is a GNU extension.

• Variables are read from and communicated via the environment. See Section 6.9 [Variables from the Environment], page 78.

• Options passed through the variable MAKEFLAGS to recursive invocations of make. See Section 5.6.3 [Communicating Options to a Sub-make], page 60.

• The automatic variable $% is set to the member name in an archive reference. See Section 10.5.3 [Automatic Variables], page 118.

• The automatic variables $@, $*, $<, $%, and $? have corresponding forms like $(@F) and $(@D). We have generalized this to $^as an obvious extension. See Section 10.5.3 [Automatic Variables], page 118.

• Substitution variable references. See Section 6.1 [Basics of Variable References], page 65.

• The command-line options ‘-b’ and ‘-m’, accepted and ignored. In System V make, these options actually do something.

• Execution of recursive commands to run make via the variable MAKE even if ‘-n’, ‘-q’ or ‘-t’ is specified. See Section 5.6 [Recursive Use of make], page 56.

• Support for suffix ‘.a’ in suffix rules. See Section 11.3 [Archive Suffix Rules], page 129. This feature is obsolete in GNU make, because the general feature of rule chaining (see Section 10.4 [Chains of Implicit Rules], page 114) allows one pattern rule for installing members in an archive (see Section 11.2 [Archive Update], page 127) to be sufficient.

• The arrangement of lines and backslash-newline combinations in commands is retained when the commands are printed, so they appear as they do in the makefile, except for the stripping of initial whitespace.

The following features were inspired by various other versions of make. In some cases it is unclear exactly which versions inspired which others.

• Pattern rules using ‘%’. This has been implemented in several versions of make. We’re not sure who invented it first, but it’s been spread around a bit. See Section 10.5 [Defining and Redefining Pattern Rules], page 115.

• Rule chaining and implicit intermediate files. This was implemented by Stu Feldman in his version of make for AT&T Eighth Edition Research Unix, and later by Andrew Hume of AT&T Bell Labs in his mk program (where he terms it “transitive closure”). We do not really know if we got this from either of them or thought it up ourselves at the same time. See Section 10.4 [Chains of Implicit Rules], page 114.

• The automatic variable $^containing a list of all dependencies of the current target. We did not invent this, but we have no idea who did. See Section 10.5.3 [Automatic Variables], page 118.

• The “what if” flag (‘-W’ in GNU make) was (as far as we know) invented by Andrew Hume in mk. See Section 9.3 [Instead of Executing the Commands], page 99.

• The concept of doing several things at once (parallelism) exists in many incarnations of make and similar programs, though not in the System V or BSD implementations. See Section 5.2 [Command Execution], page 52.

• Modified variable references using pattern substitution come from SunOS 4.0. See Section 6.1 [Basics of Variable References], page 65. This functionality was provided in GNU make by the patsubst function before the alternate syntax was implemented for compatibility with SunOS 4.0. It is not altogether clear who inspired whom, since GNU make had patsubst before SunOS 4.0 was released.

• The special significance of ‘+’ characters preceding command lines (see Section 9.3 [Instead of Executing the Commands], page 99) is mandated by draft 11.2 of IEEE Std 1003.2 (POSIX).

• The ‘+=’ syntax to append to the value of a variable comes from SunOS 4.0 make. See Section 6.6 [Appending More Text to Variables], page 74.

The remaining features are inventions new in GNU make:

• Use the ‘-v’ or ‘--version’ option to print version and copyright information.

• Use the ‘-h’ or ‘--help’ option to summarize the options to make.

• Simply-expand variables. See Section 6.2 [The Two Flavors of Variables], page 66.

• Pass command-line variable assignments automatically through the variable MAKE to recursive make invocations. See Section 5.6 [Recursive Use of make], page 56.

• Use the ‘-C’ or ‘--directory’ command option to change directory. See Section 9.7 [Summary of Options], page 103.

• Make verbatim variable definitions with define. See Section 6.8 [Defining Variables Verbatim], page 77.

• Declare phony targets with the special target .PHONY.

Andrew Hume of AT&T Bell Labs implemented a similar feature with a different syntax in his mk program. This seems to be a case of parallel discovery. See Section 4.4 [Phony Targets], page 37.

• Manipulate text by calling functions. See Chapter 8 [Functions for Transforming Text], page 85.

• Use the ‘-o’ or ‘--old-file’ option to pretend a file’s modification-time is old. See Section 9.4 [Avoiding Recompilation of Some Files], page 101.

• Conditional execution.

This feature has been implemented numerous times in various versions of make; it seems a natural extension derived from the features of the C preprocessor and similar macro languages and is not a revolutionary concept. See Chapter 7 [Conditional Parts of Makefiles], page 79.

• Specify the included makefile search path. See Section 3.3 [Including Other Makefiles], page 24.

• Specify extra makefiles to read. See Section 3.4 [The Variable MAKEFILES], page 26.

• Strip leading sequences of ‘./’ from file names, so that ‘./file’ and ‘file’ are considered to be the same file.

• Use a special search method for library dependencies written in the form ‘-lname’. See Section 4.3.5 [Directory Search for Link Libraries], page 37.

• Allow suffixes for suffix rules (see Section 10.7 [Old-Fashioned Suffix Rules], page 123) to contain any characters. In other version of make, they must begin with ‘.’ and not contain any ‘/’ characters.

• Keep track of the current level of make recursion using the variable MAKELEVEL. See Section 5.6 [Recursive Use of make], page 56.

• Specify static pattern rules. See Section 4.10 [Static Pattern Rules], page 44.

• Provide selective vpath search. See Section 4.3 [Searching Directories for Dependencies], page 33.

• Provide computed variable references. See Section 6.1 [Basics of Variable References], page 65.

• Update makefiles. See Section 3.5 [How Makefiles Are Remade], page 26. System V make has a very, very limited form of this functionality in that it will check out SCCS files for makefiles.

• Various new built-in implicit rules. See Section 10.2 [Catalogue of Implicit Rules], page 109.

134 GNU make

13 Incompatibilities and Missing Features

The make programs in various other systems support a few features that are not implemented in GNU make. Draft 11.2 of the POSIX.2 standard which specifies make does not require any of these features.

• A target of the form ‘file((entry))’ stands for a member of archive file file. The member is chosen, not by name, but by being an object file which defines the linker symbol entry. This feature was not put into GNU make because of the nonmodularity of putting knowledge into make of the internal format of archive file symbol tables. See Section 11.2.1 [Updating Archive Symbol Directories], page 128.

• Suffixes (used in suffix rules) that end with the character ‘~’ have a special meaning to System V make; they refer to the SCCS file that corresponds to the file one would get without the ‘~’. For example, the suffix rule ‘.c~.o’ would make the file ‘n.o’ file from the SCCS file ‘s.n.c’. For complete coverage, a whole series of such suffix rules is required. See Section 10.7 [Old-Fashioned Suffix Rules], page 123. In GNU make, this entire series of cases is handled by two pattern rules for extraction from SCCS, in combination with the general feature of rule chaining. See Section 10.4 [Chains of Implicit Rules], page 114.

• In System V make, the string ‘$$@’ has the strange meaning that, in the dependencies of a rule with multiple targets, it stands for the particular target that is being processed. This is not defined in GNU make because ‘$$’ should always stand for an ordinary ‘$’. It is possible to get this functionality through the use of static pattern rules (see Section 4.10 [Static Pattern Rules], page 44). The System V make rule:

     $(targets): $$@.o lib.a

can be replaced with the GNU make static pattern rule:

     $(targets): %: %.o lib.a

• In System V and 4.3 BSD make, files found by VPATH search (see Section 4.3 [Searching Directories for Dependencies], page 33) have their names changed inside command strings. We feel it is much cleaner to always use automatic variables and thus make this feature obsolete.

• In some Unix makes, implicit rule search (see Chapter 10 [Using Implicit Rules], page 107) is

apparently done for all targets, not just those without commands. This means you can do:

     foo.o:  cc -c foo.c

and Unix make will intuit that ‘foo.o’ depends on ‘foo.c’. We feel that such usage is broken. The dependency properties of make are well-defined (for GNU make, at least), and doing such a thing simply does not fit the model.

• GNU make does not include any built-in implicit rules for compiling or preprocessing EFL programs. If we hear of anyone who is using EFL, we will gladly add them.

• It appears that in SVR4 make, a suffix rule can be specified with no commands, and it is treated

as if it had empty commands (see Section 5.8 [Empty Commands], page 63). For example:

     .c.a:

will override the built-in ‘.c.a’ suffix rule. We feel that it is cleaner for a rule without commands to always simply add to the dependency list for the target. The above example can be easily rewritten to get the desired behavior in

GNU make:
     .c.a: ;

14 Makefile Conventions

This chapter describes conventions for writing the Makefiles for GNU programs.

14.1 General Conventions for Makefiles

Every Makefile should contain this line:

     SHELL = /bin/sh

to avoid trouble on systems where the SHELL variable might be inherited from the environment. (This is never a problem with GNU make.)

Don’t assume that ‘.’ is in the path for command execution. When you need to run programs that are a part of your package during the make, please make sure that it uses ‘./’ if the program is built as part of the make or ‘$(srcdir)/’ if the file is an unchanging part of the source code. Without one of these prefixes, the current search path is used.

The distinction between ‘./’ and ‘$(srcdir)/’ is important when using the ‘--srcdir’ option to ‘configure’. A rule of the form:

     foo.1 : foo.man sedscript
             sed -e sedscript foo.man > foo.1

will fail when the current directory is not the source directory, because ‘foo.man’ and ‘sedscript’ are not in the current directory.

When using GNU make, relying on ‘VPATH’ to find the source file will work in the case where there is a single dependency file, since the ‘make’ automatic variable ‘$<’ will represent the source file wherever it is. (Many versions of make set ‘$<’ only in implicit rules.) A makefile target like

     foo.o : bar.c
             $(CC) -I. -I$(srcdir) $(CFLAGS) -c bar.c -o foo.o

should instead be written as

     foo.o : bar.c
             $(CC) $(CFLAGS) $< -o $@

in order to allow ‘VPATH’ to work correctly. When the target has multiple dependencies, using an explicit ‘$(srcdir)’ is the easiest way to make the rule work well. For example, the target above for ‘foo.1’ is best written as:

     foo.1 : foo.man sedscript
             sed -s $(srcdir)/sedscript $(srcdir)/foo.man > foo.1

14.2 Utilities in Makefiles

Write the Makefile commands (and any shell scripts, such as configure) to run in sh, not in csh. Don’t use any special features of ksh or bash.

The configure script and the Makefile rules for building and installation should not use any utilities directly except these:

     cat cmp cp echo egrep expr grep
     ln mkdir mv pwd rm rmdir sed test touch

Stick to the generally supported options for these programs. For example, don’t use ‘mkdir -p’, convenient as it may be, because most systems don’t support it.

The Makefile rules for building and installation can also use compilers and related programs, but should do so via make variables so that the user can substitute alternatives. Here are some of the programs we mean:

     ar bison cc flex install ld lex
     make makeinfo ranlib texi2dvi yacc

When you use ranlib, you should test whether it exists, and run it only if it exists, so that the distribution will work on systems that don’t have ranlib.

If you use symbolic links, you should implement a fallback for systems that don’t have symbolic links.

It is ok to use other utilities in Makefile portions (or scripts) intended only for particular systems where you know those utilities to exist.

14.3 Standard Targets for Users

All GNU programs should have the following targets in their Makefiles:

‘all’ Compile the entire program. This should be the default target. This target need not rebuild any documentation files; info files should normally be included in the distribution, and DVI files should be made only when explicitly asked for.

‘install’ Compile the program and copy the executables, libraries, and so on to the file names where they should reside for actual use. If there is a simple test to verify that a program is properly installed then run that test. Use ‘-’ before any command for installing a man page, so that make will ignore any errors. This is in case there are systems that don’t have the Unix man page documentation system installed. In the future, when we have a standard way of installing info files, ‘install’ targets will be the proper place to do so.

‘uninstall’

Delete all the installed files that the ‘install’ target would create (but not the noninstalled files such as ‘make all’ would create).

‘clean’ Delete all files from the current directory that are normally created by building the program. Don’t delete the files that record the configuration. Also preserve files that could be made by building, but normally aren’t because the distribution comes with them. Delete ‘.dvi’ files here if they are not part of the distribution.

‘distclean’

Delete all files from the current directory that are created by configuring or building the program. If you have unpacked the source and built the program without creating any other files, ‘make distclean’ should leave only the files that were in the distribution.

‘mostlyclean’

Like ‘clean’, but may refrain from deleting a few files that people normally don’t want to recompile. For example, the ‘mostlyclean’ target for GCC does not delete ‘libgcc.a’, because recompiling it is rarely necessary and takes a lot of time.

‘realclean’

Delete everything from the current directory that can be reconstructed with this Makefile. This typically includes everything deleted by distclean, plus more: C source files produced by Bison, tags tables, info files, and so on. One exception, however: ‘make realclean’ should not delete ‘configure’ even if ‘configure’ can be remade using a rule in the Makefile. More generally, ‘make realclean’ should not delete anything that needs to exist in order to run ‘configure’ and then begin to build the program.

‘TAGS’ Update a tags table for this program.

‘info’ Generate any info files needed. The best way to write the rules is as follows:

     info:  foo.info
     foo.info: $(srcdir)/foo.texi $(srcdir)/chap1.texi $(srcdir)/chap2.texi
             $(MAKEINFO) $(srcdir)/foo.texi

You must define the variable MAKEINFO in the Makefile. It should run the Makeinfo program, which is part of the Texinfo2 distribution.

‘dvi’ Generate DVI files for all TeXinfo documentation. For example:

     dvi: foo.dvi
     foo.dvi: $(srcdir)/foo.texi $(srcdir)/chap1.texi $(srcdir)/chap2.texi
             $(TEXI2DVI) $(srcdir)/foo.texi

You must define the variable TEXI2DVI in the Makefile. It should run the program texi2dvi, which is part of the Texinfo2 distribution. Alternatively, write just the dependencies, and allow GNU Make to provide the command.

‘dist’ Create a distribution tar file for this program. The tar file should be set up so that the file names in the tar file start with a subdirectory name which is the name of the package it is a distribution for. This name can include the version number. For example, the distribution tar file of GCC version 1.40 unpacks into a subdirectory

named ‘gcc-1.40’.

The easiest way to do this is to create a subdirectory appropriately named, use ln or cp to install the proper files in it, and then tar that subdirectory. The dist target should explicitly depend on all non-source files that are in the distribution, to make sure they are up to date in the distribution. See section “Making Releases” in GNU Coding Standards.

‘check’ Perform self-tests (if any). The user must build the program before running the tests, but need not install the program; you should write the self-tests so that they work when the program is built but not installed.

The following target is suggested as a conventional name, for programs in which it is useful.

installcheck

Perform installation tests (if any). The user must build and install the program before running the tests. You should not assume that ‘$(bindir)’ is in the search path.

installdirs

It’s useful to add a target named ‘installdirs’ to create the directories where files are installed, and their parent directories. You can use a rule like this:

               # Make sure all installation directories, e.g. $(bindir) actually exist by
               # making them if necessary.
               installdirs:
                       for file in $(bindir) $(datadir) $(libdir) $(infodir) $(mandir) ; do \
                          oIFS="$${IFS}"; IFS='/'; set - $${file}; IFS="$${oIFS}"; \
                          pathcomp=''; test ".$${1}" = "." && shift; \
                          while test $$# -ne 0 ; do \
                            pathcomp="$${pathcomp}/$${1}"; shift; \
                            if test ! -d "$${pathcomp}"; then \
                               echo "making directory $$pathcomp" 1>&2 ; \
                               mkdir "$${pathcomp}"; \
                            fi; \
                          done; \
                       done

14.4 Variables for Specifying Commands

Makefiles should provide variables for overriding certain commands, options, and so on.

In particular, you should run most utility programs via variables. Thus, if you use Bison, have a variable named BISON whose default value is set with ‘BISON = bison’, and refer to it with $(BISON) whenever you need to use Bison.

File management utilities such as ln, rm, mv, and so on, need not be referred to through variables in this way, since users don’t need to replace them with other programs.

Each program-name variable should come with an options variable that is used to supply options to the program. Append ‘FLAGS’ to the program-name variable name to get the options variable name—for example, BISONFLAGS. (The name CFLAGS is an exception to this rule, but we keep it because it is standard.) Use CPPFLAGS in any compilation command that runs the preprocessor, and use LDFLAGS in any compilation command that does linking as well as in any direct use of ld.

If there are C compiler options that must be used for proper compilation of certain files, do not include them in CFLAGS. Users expect to be able to specify CFLAGS freely themselves. Instead, arrange to pass the necessary options to the C compiler independently of CFLAGS, by writing them explicitly in the compilation commands or by defining an implicit rule, like this:

     CFLAGS = -g

     ALL_CFLAGS = -I. $(CFLAGS)
     .c.o:
             $(CC) -c $(CPPFLAGS) $(ALL_CFLAGS) $<

Do include the ‘-g’ option in CFLAGS, because that is not required for proper compilation. You can consider it a default that is only recommended. If the package is set up so that it is compiled with GCC by default, then you might as well include ‘-O’ in the default value of CFLAGS as well.

Put CFLAGS last in the compilation command, after other variables containing compiler options, so the user can use CFLAGS to override the others.

Every Makefile should define the variable INSTALL, which is the basic command for installing a file into the system.

Every Makefile should also define variables INSTALL_PROGRAM and INSTALL_DATA. (The default for each of these should be $(INSTALL).) Then it should use those variables as the commands for actual installation, for executables and nonexecutables respectively. Use these variables as follows:

     $(INSTALL_PROGRAM) foo $(bindir)/foo
     $(INSTALL_DATA) libfoo.a $(libdir)/libfoo.a

Always use a file name, not a directory name, as the second argument of the installation commands. Use a separate command for each file to be installed.

14.5 Variables for Installation Directories

Installation directories should always be named by variables, so it is easy to install in a nonstandard place. The standard names for these variables are:

‘prefix’ A prefix used in constructing the default values of the variables listed below. The

default value of prefix should be ‘/usr/local’ (at least for now).

‘exec_prefix’

A prefix used in constructing the default values of the some of the variables listed below. The default value of exec_prefix should be $(prefix). Generally, $(exec_prefix) is used for directories that contain machine-specific files (such as executables and subroutine libraries), while $(prefix) is used directly for other directories.

‘bindir’ The directory for installing executable programs that users can run. This should nor-

mally be ‘/usr/local/bin’, but write it as ‘$(exec_prefix)/bin’.

‘libdir’ The directory for installing executable files to be run by the program rather than by users. Object files and libraries of object code should also go in this directory. The idea is that this directory is used for files that pertain to a specific machine architecture, but need not be in the path for commands. The value of libdir should normally be

‘/usr/local/lib’, but write it as ‘$(exec_prefix)/lib’.

‘datadir’ The directory for installing read-only data files which the programs refer to while they run. This directory is used for files which are independent of the type of machine being used. This should normally be ‘/usr/local/lib’, but write it as ‘$(prefix)/lib’.

‘statedir’

The directory for installing data files which the programs modify while they run. These files should be independent of the type of machine being used, and it should be possible to share them among machines at a network installation. This should normally be

‘/usr/local/lib’, but write it as ‘$(prefix)/lib’.
‘includedir’

The directory for installing header files to be included by user programs with the C ‘#include’ preprocessor directive. This should normally be ‘/usr/local/include’,

but write it as ‘$(prefix)/include’.

Most compilers other than GCC do not look for header files in ‘/usr/local/include’. So installing the header files this way is only useful with GCC. Sometimes this is not a problem because some libraries are only really intended to work with GCC. But some libraries are intended to work with other compilers. They should install their header files in two places, one specified by includedir and one specified by oldincludedir.

‘oldincludedir’

The directory for installing ‘#include’ header files for use with compilers other than GCC. This should normally be ‘/usr/include’. The Makefile commands should check whether the value of oldincludedir is empty. If it is, they should not try to use it; they should cancel the second installation of the header files. A package should not replace an existing header in this directory unless the header came from the same package. Thus, if your Foo package provides a header file ‘foo.h’, then it should install the header file in the oldincludedir directory if either (1) there is no ‘foo.h’ there or (2) the ‘foo.h’ that exists came from the Foo package. The way to tell whether ‘foo.h’ came from the Foo package is to put a magic string in the file—part of a comment—and grep for that string.

‘mandir’ The directory for installing the man pages (if any) for this package. It should include the suffix for the proper section of the manual—usually ‘1’ for a utility.

‘man1dir’ The directory for installing section 1 man pages.

‘man2dir’ The directory for installing section 2 man pages.

‘…’ Use these names instead of ‘mandir’ if the package needs to install man pages in more than one section of the manual. Don’t make the primary documentation for any GNU software be a man page. Write a manual in Texinfo instead. Man pages are just for the sake of people running GNU software on Unix, which is a secondary application only.

‘manext’ The file name extension for the installed man page. This should contain a period followed by the appropriate digit.

‘infodir’ The directory for installing the info files for this package. By default, it should be

          ‘/usr/local/info’, but it should be written as ‘$(prefix)/info’.

‘srcdir’ The directory for the sources being compiled. The value of this variable is normally inserted by the configure shell script.

For example:

     # Common prefix for installation directories.
     # NOTE: This directory must exist when you start the install.
     prefix = /usr/local
     exec_prefix = $(prefix)
     # Where to put the executable for the command `gcc'.
     bindir = $(exec_prefix)/bin
     # Where to put the directories used by the compiler.
     libdir = $(exec_prefix)/lib
     # Where to put the Info files.
     infodir = $(prefix)/info

If your program installs a large number of files into one of the standard user-specified directories, it might be useful to group them into a subdirectory particular to that program. If you do this, you should write the install rule to create these subdirectories.

Do not expect the user to include the subdirectory name in the value of any of the variables listed above. The idea of having a uniform set of variable names for installation directories is to enable the user to specify the exact same values for several different GNU packages. In order for this to be useful, all the packages must be designed so that they will work sensibly when the user does so.

Appendix A Quick Reference

This appendix summarizes the directives, text manipulation functions, and special variables which GNU make understands. See Section 4.7 [Special Targets], page 40, Section 10.2 [Catalogue of Implicit Rules], page 109, and Section 9.7 [Summary of Options], page 103. for other summaries.

Here is a summary of the directives GNU make recognizes:

define variable

endef

Define a multi-line, recursively-expanded variable.

See Section 5.7 [Sequences], page 62.

ifdef variable ifndef variable

ifeq (a,b)
ifeq "a" "b"
ifeq 'a' 'b'
ifneq (a,b)
ifneq "a" "b"
ifneq 'a' 'b'
else
endif

Conditionally evaluate part of the makefile. See Chapter 7 [Conditionals], page 79.

include file

Include another makefile.

See Section 3.3 [Including Other Makefiles], page 24.

override variable = value override variable := value override variable += value

override define variable
endef

Define a variable, overriding any previous definition, even one from the command line. See Section 6.7 [The override Directive], page 76.

export

Tell make to export all variables to child processes by default.

See Section 5.6.2 [Communicating Variables to a Sub-make], page 57. export variable export variable = value export variable := value export variable += value

unexport variable

Tell make whether or not to export a particular variable to child processes. See Section 5.6.2 [Communicating Variables to a Sub-make], page 57.

vpath pattern path Specify a search path for files matching a ‘%’ pattern. See Section 4.3.2 [The vpath Directive], page 34.

vpath pattern Remove all search paths previously specified for pattern.

vpath

Remove all search paths previously specified in any vpath directive.

Here is a summary of the text manipulation functions (see Chapter 8 [Functions], page 85):

$(subst from,to,text)

Replace from with to in text. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(patsubst pattern,replacement,text)

Replace words matching pattern with replacement in text. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(strip string)

Remove excess whitespace characters from string. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(findstring find,text)

Locate find in text. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(filter pattern…,text)

Select words in text that match one of the pattern words. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(filter-out pattern. . .,text)

Select words in text that do not match any of the pattern words. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(sort list)

Sort the words in list lexicographically, removing duplicates. See Section 8.2 [Functions for String Substitution and Analysis], page 86.

$(dir names…)

Extract the directory part of each file name. See Section 8.3 [Functions for File Names], page 89.

$(notdir names. . .)

Extract the non-directory part of each file name. See Section 8.3 [Functions for File Names], page 89.

$(suffix names. . .)

Extract the suffix (the last ‘.’ and following characters) of each file name. See Section 8.3 [Functions for File Names], page 89.

$(basename names. . .)

Extract the base name (name without suffix) of each file name. See Section 8.3 [Functions for File Names], page 89.

$(addsuffix suffix,names. . .)

Append suffix to each word in names. See Section 8.3 [Functions for File Names], page 89.

$(addprefix prefix,names. . .)

Prepend prefix to each word in names. See Section 8.3 [Functions for File Names], page 89.

$(join list1,list2)

Join two parallel lists of words. See Section 8.3 [Functions for File Names], page 89.

$(word n,text)

Extract the nth word (one-origin) of text. See Section 8.3 [Functions for File Names], page 89.

$(words text)

Count the number of words in text. See Section 8.3 [Functions for File Names], page 89.

$(firstword names. . .)

Extract the first word of names. See Section 8.3 [Functions for File Names], page 89.

$(wildcard pattern. . .)

Find file names matching a shell file name pattern (not a ‘%’ pattern). See Section 4.2.3 [The Function wildcard], page 32.

$(shell command)

Execute a shell command and return its output. See Section 8.6 [The shell Function], page 95.

$(origin variable)

Return a string describing how the make variable variable was defined. See Section 8.5 [The origin Function], page 93.

$(foreach var,words,text)

Evaluate text with var bound to each word in words, and concatenate the results. See Section 8.4 [The foreach Function], page 91.

Here is a summary of the automatic variables. See Section 10.5.3 [Automatic Variables], page 118, for full information.

$@ The file name of the target. $% The target member name, when the target is an archive member. $< The name of the first dependency. $? The names of all the dependencies that are newer than the target, with spaces between

them. For dependencies which are archive members, only the member named is used (see Chapter 11 [Archives], page 127).

$^The names of all the dependencies, with spaces between them. For dependencies which are archive members, only the member named is used (see Chapter 11 [Archives], page 127).

$* The stem with which an implicit rule matches (see Section 10.5.4 [How Patterns Match],

page 120).

$(@D)

$(@F) The directory part and the file-within-directory part of $@.

$(*D)

$(*F) The directory part and the file-within-directory part of $*.

$(%D)

$(%F) The directory part and the file-within-directory part of $%.

$(<D)

$(<F) The directory part and the file-within-directory part of $<.

$(^D)

$(^F) The directory part and the file-within-directory part of $^.

$(?D)

$(?F) The directory part and the file-within-directory part of $?.

These variables are used specially by GNU make:

MAKEFILES

Makefiles to be read on every invocation of make. See Section 3.4 [The Variable MAKEFILES], page 26.

VPATH

Directory search path for files not found in the current directory. See Section 4.3.1 [VPATH Search Path for All Dependencies], page 33.

SHELL

The name of the system default command interpreter, usually ‘/bin/sh’. You can set SHELL in the makefile to change the shell used to run commands. See Section 5.2 [Command Execution], page 52.

MAKE

The name with which make was invoked. Using this variable in commands has special meaning. See Section 5.6.1 [How the MAKE Variable Works], page 56.

MAKELEVEL

The number of levels of recursion (sub-makes). See Section 5.6.2 [Variables/Recursion], page 57.

MAKEFLAGS
MFLAGS

The flags given to make. You can set this in the environment or a makefile to set flags. See Section 5.6.3 [Communicating Options to a Sub-make], page 60.

SUFFIXES

The default list of suffixes before make reads any makefiles.

150 GNU make

Appendix B Complex Makefile Example

Here is the makefile for the GNU tar program. This is a moderately complex makefile.

Because it is the first target, the default goal is ‘all’. An interesting feature of this makefile is that ‘testpad.h’ is a source file automatically created by the testpad program, itself compiled

from ‘testpad.c’.

If you type ‘make’ or ‘make all’, then make creates the ‘tar’ executable, the ‘rmt’ daemon that provides remote tape access, and the ‘tar.info’ Info file.

If you type ‘make install’, then make not only creates ‘tar’, ‘rmt’, and ‘tar.info’, but also installs them.

If you type ‘make clean’, then make removes the ‘.o’ files, and the ‘tar’, ‘rmt’, ‘testpad’,

‘testpad.h’, and ‘core’ files.

If you type ‘make distclean’, then make not only removes the same files as does ‘make clean’ but also the ‘TAGS’, ‘Makefile’, and ‘config.status’ files. (Although it is not evident, this makefile (and ‘config.status’) is generated by the user with the configure program, which is provided in the tar distribution, but is not shown here.)

If you type ‘make realclean’, then make removes the same files as does ‘make distclean’ and also removes the Info files generated from ‘tar.texinfo’.

In addition, there are targets shar and dist that create distribution kits.

     # Generated automatically from Makefile.in by configure.
     # Un*x Makefile for GNU tar program.
     # Copyright (C) 1991 Free Software Foundation, Inc.
     # This program is free software; you can redistribute
     # it and/or modify it under the terms of the GNU
     # General Public License . . .

… …

     SHELL = /bin/sh
     #### Start of system configuration section. ####

srcdir = .
# If you use gcc, you should either run the
# fixincludes script that comes with it or else use
# gcc with the -traditional option.  Otherwise ioctl
# calls will be compiled incorrectly on some systems.
CC = gcc -O
YACC = bison -y
INSTALL = /usr/local/bin/install -c
INSTALLDATA = /usr/local/bin/install -c -m 644
# Things you might add to DEFS:
# -DSTDC_HEADERS        If you have ANSI C headers and
#                       libraries.
# -DPOSIX               If you have POSIX.1 headers and
#                       libraries.
# -DBSD42               If you have sys/dir.h (unless
#                       you use -DPOSIX), sys/file.h,
#                       and st_blocks in `struct stat'.
# -DUSG                 If you have System V/ANSI C
#                       string and memory functions
#                       and headers, sys/sysmacros.h,
#                       fcntl.h, getcwd, no valloc,
#                       and ndir.h (unless
#                       you use -DDIRENT).
# -DNO_MEMORY_H         If USG or STDC_HEADERS but do not
#                       include memory.h.
# -DDIRENT              If USG and you have dirent.h
#                       instead of ndir.h.
# -DSIGTYPE=int         If your signal handlers
#                       return int, not void.
# -DNO_MTIO             If you lack sys/mtio.h
#                       (magtape ioctls).
# -DNO_REMOTE           If you do not have a remote shell
#                       or rexec.
# -DUSE_REXEC           To use rexec for remote tape
#                       operations instead of
#                       forking rsh or remsh.
# -DVPRINTF_MISSING     If you lack vprintf function
#                       (but have _doprnt).
# -DDOPRNT_MISSING      If you lack _doprnt function.
#                       Also need to define
#                       -DVPRINTF_MISSING.
# -DFTIME_MISSING       If you lack ftime system call.
# -DSTRSTR_MISSING      If you lack strstr function.
# -DVALLOC_MISSING      If you lack valloc function.
# -DMKDIR_MISSING       If you lack mkdir and
#                       rmdir system calls.

# -DRENAME_MISSING      If you lack rename system call.
# -DFTRUNCATE_MISSING   If you lack frtruncate
#                       system call.
# -DV7                  On Version 7 Unix (not
#                       tested in a long time).
# -DEMUL_OPEN3          If you lack a 3-argument version
#                       of open, and want to emulate it
#                       with system calls you do have.
# -DNO_OPEN3            If you lack the 3-argument open
#                       and want to disable the tar -k
#                       option instead of emulating open.
# -DXENIX               If you have sys/inode.h
#                       and need it 94 to be included.
DEFS =  -DSIGTYPE=int -DDIRENT -DSTRSTR_MISSING \
        -DVPRINTF_MISSING -DBSD42
# Set this to rtapelib.o unless you defined NO_REMOTE,
# in which case make it empty.
RTAPELIB = rtapelib.o
LIBS =
DEF_AR_FILE = /dev/rmt8
DEFBLOCKING = 20
CDEBUG = -g
CFLAGS = $(CDEBUG) -I. -I$(srcdir) $(DEFS) \
        -DDEF_AR_FILE=\"$(DEF_AR_FILE)\" \
        -DDEFBLOCKING=$(DEFBLOCKING)
LDFLAGS = -g
prefix = /usr/local
# Prefix for each installed program,
# normally empty or `g'.
binprefix =
# The directory to install tar in.
bindir = $(prefix)/bin
# The directory to install the info files in.
infodir = $(prefix)/info
#### End of system configuration section. ####
SRC1 =  tar.c create.c extract.c buffer.c \
        getoldopt.c update.c gnu.c mangle.c
SRC2 =  version.c list.c names.c diffarch.c \
        port.c wildmat.c getopt.c
SRC3 =  getopt1.c regex.c getdate.y
SRCS =  $(SRC1) $(SRC2) $(SRC3)

OBJ1 =  tar.o create.o extract.o buffer.o \
getoldopt.o update.o gnu.o mangle.o
OBJ2 =  version.o list.o names.o diffarch.o \
port.o wildmat.o getopt.o
OBJ3 =  getopt1.o regex.o getdate.o $(RTAPELIB)
OBJS =  $(OBJ1) $(OBJ2) $(OBJ3)
AUX =   README COPYING ChangeLog Makefile.in  \
makefile.pc configure configure.in \
tar.texinfo tar.info* texinfo.tex \
tar.h port.h open3.h getopt.h regex.h \
rmt.h rmt.c rtapelib.c alloca.c \
msd_dir.h msd_dir.c tcexparg.c \
level-0 level-1 backup-specs testpad.c
all:    tar rmt tar.info
tar:    $(OBJS)
$(CC) $(LDFLAGS) -o $@ $(OBJS) $(LIBS)
rmt:    rmt.c
$(CC) $(CFLAGS) $(LDFLAGS) -o $@ rmt.c
tar.info: tar.texinfo
makeinfo tar.texinfo
install: all
$(INSTALL) tar $(bindir)/$(binprefix)tar
-test ! -f rmt || $(INSTALL) rmt /etc/rmt
$(INSTALLDATA) $(srcdir)/tar.info* $(infodir)
$(OBJS): tar.h port.h testpad.h
regex.o buffer.o tar.o: regex.h
# getdate.y has 8 shift/reduce conflicts.
testpad.h: testpad
./testpad
testpad: testpad.o
$(CC) -o $@ testpad.o
TAGS:   $(SRCS)
etags $(SRCS)
clean:
rm -f *.o tar rmt testpad testpad.h core
distclean: clean
rm -f TAGS Makefile config.status
realclean: distclean
rm -f tar.info*

shar: $(SRCS) $(AUX)
shar $(SRCS) $(AUX) | compress \
  > tar-`sed -e '/version_string/!d' \
             -e 's/[^0-9.]*\([0-9.]*\).*/\1/' \
             -e q
             version.c`.shar.Z
dist: $(SRCS) $(AUX)
echo tar-`sed \
     -e '/version_string/!d' \
     -e 's/[^0-9.]*\([0-9.]*\).*/\1/' \
     -e q
     version.c` > .fname
-rm -rf `cat .fname`
mkdir `cat .fname`
ln $(SRCS) $(AUX) `cat .fname`
-rm -rf `cat .fname` .fname
tar chZf `cat .fname`.tar.Z `cat .fname`
tar.zoo: $(SRCS) $(AUX)
-rm -rf tmp.dir
-mkdir tmp.dir
-rm tar.zoo
for X in $(SRCS) $(AUX) ; do \
    echo $$X ; \
    sed 's/$$/^M/' $$X \
    > tmp.dir/$$X ; done
cd tmp.dir ; zoo aM ../tar.zoo *
-rm -rf tmp.dir

156 GNU make

Index of Concepts 157

Index of Concepts

#

# (comments), in commands . . . . . . . . . . . . . . . . . . . . . . . 51 # (comments), in makefile . . . . . . . . . . . . . . . . . . . . . . . . . . 23 #include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

$

$, in function call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 $, in rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 $, in variable name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 $, in variable reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

%

%, in pattern rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 %, quoting in patsubst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 %, quoting in static pattern. . . . . . . . . . . . . . . . . . . . . . . . . 45 %, quoting in vpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 %, quoting with \ (backslash) . . . . . . . . . . . . . . . 35, 45, 86

*

* (wildcard character) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

,

,v (RCS file extension) . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

-

- (in commands). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 -, and define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --assume-new. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 106 --assume-new, and recursion . . . . . . . . . . . . . . . . . . . . . . . 60 --assume-old. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 --assume-old, and recursion . . . . . . . . . . . . . . . . . . . . . . . 60 --debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 --directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 103

--directory, and --print-directory . . . . . . . . . . . . . 61

--directory, and recursion . . . . . . . . . . . . . . . . . . . . . . . . 60 --dry-run. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 99, 105 --environment-overrides . . . . . . . . . . . . . . . . . . . . . . . . 104 --file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 97, 104 --file, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 --ignore-errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 104 --include-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 104 --include-dir, and recursion. . . . . . . . . . . . . . . . . . . . . . 60 --jobs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 104 --jobs, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 --just-print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 99, 105 --keep-going . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 102, 104 --load-average . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 104 --makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 97, 104 --max-load . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 104 --new-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 106 --new-file, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . 60 --no-builtin-rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 --no-keep-going . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 --no-print-directory. . . . . . . . . . . . . . . . . . . . . . . . 61, 106 --old-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 --old-file, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . 60 --print-data-base. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 --print-directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

--print-directory, and --directory . . . . . . . . . . . . . 61

--print-directory, and recursion . . . . . . . . . . . . . . . . . 61 --print-directory, disabling. . . . . . . . . . . . . . . . . . . . . . 61 --question . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 105 --quiet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 105 --recon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 99, 105 --silent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 105 --stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 --touch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 106 --touch, and recursion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 --what-if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 106 -b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 -C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 103 -C, and -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -C, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 -f. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 97, 104 -f, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 -i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 104 -I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 104 -I, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 104 -j, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 102, 104 -l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 -l (library search) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 -l (load average) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 -m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 -M (to compiler) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 -n. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 99, 105 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 105 -o, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 -q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 105 -r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 -s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51, 105 -S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 -t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99, 106 -t, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100, 106 -w, and -C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -w, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -W, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -w, disabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

.

.a (archives). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 .c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .cc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .cweb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 .def. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 .DEFAULT, used to override . . . . . . . . . . . . . . . . . . . . . . . . . 28 .dvi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .ln. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .mod. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 .o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .PRECIOUS intermediate files . . . . . . . . . . . . . . . . . . . . . . 115 .r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 .s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 .S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 .sh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .sym. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 .tex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .texi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .txinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .web. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 .y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

:

:: rules (double-colon). . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 := . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67, 73

=

= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66, 73

?

? (wildcard character) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

[

[…] (wildcard characters) . . . . . . . . . . . . . . . . . . . . . . . . 30

.SYMDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

                                @

@ (in commands). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 @, and define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

                                ~

~ (tilde) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

+

+, and define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 += . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

\

\ (backslash), for continuation lines . . . . . . . . . . . . . . . . 15 \ (backslash), in commands . . . . . . . . . . . . . . . . . . . . . . . . 52 \ (backslash), to quote %. . . . . . . . . . . . . . . . . . . . 35, 45, 86

A

all (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 appending to variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 ar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 archive member targets . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 archive symbol directory updating . . . . . . . . . . . . . . . . 128 archive, suffix rule for. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 arguments of functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110, 113 assembly, rule to compile . . . . . . . . . . . . . . . . . . . . . . . . . 110 automatic generation of dependencies. . . . . . . . . . . 25, 47 automatic variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

B

backquotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 backslash (\), for continuation lines . . . . . . . . . . . . . . . . 15 backslash (\), in commands . . . . . . . . . . . . . . . . . . . . . . . . 52 backslash (\), to quote %. . . . . . . . . . . . . . . . . . . . 35, 45, 86 basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 broken pipe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 built-in special targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

C

C, rule to compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 C++, rule to compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 cc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109, 113 cd (shell command) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 56 chains of rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 check (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 clean (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 clean target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 20 cleaning up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 clobber (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . 98 co . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 combining rules by dependency . . . . . . . . . . . . . . . . . . . . 20 command line variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 commands, backslash (\) in . . . . . . . . . . . . . . . . . . . . . . . . 52 commands, comments in . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 commands, echoing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 commands, empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 commands, errors in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 commands, execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 commands, execution in parallel. . . . . . . . . . . . . . . . . . . . 52 commands, expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 commands, how to write . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 commands, instead of executing . . . . . . . . . . . . . . . . . . . . 99 commands, introduction to. . . . . . . . . . . . . . . . . . . . . . . . . 13 commands, quoting newlines in . . . . . . . . . . . . . . . . . . . . 52 commands, sequences of . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 comments, in commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 comments, in makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 compatibility in exporting. . . . . . . . . . . . . . . . . . . . . . . . . . 59 compilation, testing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 computed variable name . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 continuation lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 conventions for makefiles. . . . . . . . . . . . . . . . . . . . . . . . . . 137 ctangle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 cweave . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113

D

deducing commands (implicit rules) . . . . . . . . . . . . . . . . 18 default goal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 29 default makefile name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 default rules, last-resort. . . . . . . . . . . . . . . . . . . . . . . . . . . 122 defining variables verbatim. . . . . . . . . . . . . . . . . . . . . . . . . 77 deletion of target files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 dependencies, automatic generation. . . . . . . . . . . . . 25, 47 dependencies, introduction to . . . . . . . . . . . . . . . . . . . . . . 13 dependencies, list of all . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 dependencies, list of changed. . . . . . . . . . . . . . . . . . . . . . 118 dependencies, varying (static pattern) . . . . . . . . . . . . . . 44 dependency . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 dependency pattern, implicit . . . . . . . . . . . . . . . . . . . . . . 116 dependency pattern, static (not implicit) . . . . . . . . . . . 44 directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 directories, printing them . . . . . . . . . . . . . . . . . . . . . . . . . . 61 directories, updating archive symbol . . . . . . . . . . . . . . 128 directory part. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 directory search (VPATH) . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 directory search (VPATH), and implicit rules . . . . . . . . . 36 directory search (VPATH), and link libraries . . . . . . . . . 37 directory search (VPATH), and shell commands . . . . . . 36 dist (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 distclean (standard target) . . . . . . . . . . . . . . . . . . . . . . . 98 dollar sign ($) in rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 dollar sign ($), in function call . . . . . . . . . . . . . . . . . . . . . 85 dollar sign ($), in variable name. . . . . . . . . . . . . . . . . . . . 70 dollar sign ($), in variable reference . . . . . . . . . . . . . . . . 65 double-colon rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 duplicate words, removing . . . . . . . . . . . . . . . . . . . . . . . . . 88

E

echoing of commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 editor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Emacs (M-x compile) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 empty commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 empty targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 environment, and recursion . . . . . . . . . . . . . . . . . . . . . . . . 57 environment, SHELL in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 errors (in commands) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 errors with wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 execution, in parallel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 execution, instead of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 execution, of commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 exit status (errors) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 explicit rule, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . 23 exporting variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

F

f77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109, 113 features of GNU make. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 features, missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 file name functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 file name of makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 file name of makefile, how to specify. . . . . . . . . . . . . . . . 24 file name prefix, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 file name suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 file name suffix, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 file name with wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 file name, basename of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 file name, directory part . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 file name, nondirectory part. . . . . . . . . . . . . . . . . . . . . . . . 89 files, assuming new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 files, assuming old . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 files, avoiding recompilation of . . . . . . . . . . . . . . . . . . . . 101 files, intermediate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 filtering out words . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 filtering words . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 finding strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 flags for compilers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 flavors of variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 FORCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 force targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Fortran, rule to compile. . . . . . . . . . . . . . . . . . . . . . . . . . . 109 function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 functions, for file names . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 functions, for text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 functions, syntax of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

G

g++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109, 113 gcc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 generating dependencies automatically. . . . . . . . . . 25, 47 get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 globbing (wildcards) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 goal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 goal, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 29 goal, how to specify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

H

home directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

I

IEEE Standard 1003.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 implicit rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 implicit rule, and directory search . . . . . . . . . . . . . . . . . . 36 implicit rule, and VPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 implicit rule, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . 23 implicit rule, how to use . . . . . . . . . . . . . . . . . . . . . . . . . . 107 implicit rule, introduction to . . . . . . . . . . . . . . . . . . . . . . . 18 implicit rule, predefined. . . . . . . . . . . . . . . . . . . . . . . . . . . 109 implicit rule, search algorithm . . . . . . . . . . . . . . . . . . . . 125 including (MAKEFILES variable) . . . . . . . . . . . . . . . . . . . . . 26 including other makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 incompatibilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Info, rule to format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 install (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . 99 intermediate file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 intermediate files, preserving . . . . . . . . . . . . . . . . . . . . . . 115 interrupt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

J

job slots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 job slots, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 jobs, limiting based on load . . . . . . . . . . . . . . . . . . . . . . . . 53 joining lists of words . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

K

killing (interruption) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

L

last-resort default rules . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 lex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 Lex, rule to run. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 libraries for linking, directory search . . . . . . . . . . . . . . . 37 library archive, suffix rule for . . . . . . . . . . . . . . . . . . . . . 129 limiting jobs based on load. . . . . . . . . . . . . . . . . . . . . . . . . 53 link libraries, and directory search. . . . . . . . . . . . . . . . . . 37 linking, predefined rule for . . . . . . . . . . . . . . . . . . . . . . . . 110 lint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 lint, rule to run. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 list of all dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 list of changed dependencies . . . . . . . . . . . . . . . . . . . . . . 118 load average . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 loops in variable expansion. . . . . . . . . . . . . . . . . . . . . . . . . 67 lpr (shell command) . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 40

M

m2c. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 make depend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 makefile name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 makefile name, how to specify . . . . . . . . . . . . . . . . . . . . . . 24 makefile rule parts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 makefile, and MAKEFILES variable . . . . . . . . . . . . . . . . . . . 26 makefile, conventions for . . . . . . . . . . . . . . . . . . . . . . . . . . 137 makefile, how make processes . . . . . . . . . . . . . . . . . . . . . . . 16 makefile, how to write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 makefile, including . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 makefile, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 makefile, remaking of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 makefile, simple. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 makeinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 match-anything rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 missing features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 mistakes with wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 modified variable reference . . . . . . . . . . . . . . . . . . . . . . . . . 69 Modula-2, rule to compile. . . . . . . . . . . . . . . . . . . . . . . . . 110 mostlyclean (standard target) . . . . . . . . . . . . . . . . . . . . . 98 multiple rules for one target. . . . . . . . . . . . . . . . . . . . . . . . 43 multiple rules for one target (::) . . . . . . . . . . . . . . . . . . . 46 multiple targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 multiple targets, in pattern rule . . . . . . . . . . . . . . . . . . . 116

N

name of makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 name of makefile, how to specify . . . . . . . . . . . . . . . . . . . 24 nested variable reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 newline, quoting, in commands. . . . . . . . . . . . . . . . . . . . . 52 newline, quoting, in makefile . . . . . . . . . . . . . . . . . . . . . . . 15 nondirectory part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

O

obj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 OBJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 OBJECTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 objs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 OBJS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 old-fashioned suffix rules . . . . . . . . . . . . . . . . . . . . . . . . . . 123 options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 options, and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 options, setting from environment . . . . . . . . . . . . . . . . . . 61 options, setting in makefiles . . . . . . . . . . . . . . . . . . . . . . . . 61 order of pattern rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 origin of variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 overriding makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 overriding variables with arguments . . . . . . . . . . . . . . . 101 overriding with override . . . . . . . . . . . . . . . . . . . . . . . . . . 76

P

parallel execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 parts of makefile rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Pascal, rule to compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 pattern rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 pattern rules, order of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 pattern rules, static (not implicit) . . . . . . . . . . . . . . . . . . 44 pattern rules, static, syntax of . . . . . . . . . . . . . . . . . . . . . 44 pc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109, 113 phony targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 pitfalls of wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 portability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 precious targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 prefix, adding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 preserving intermediate files . . . . . . . . . . . . . . . . . . . . . . 115 preserving with .PRECIOUS . . . . . . . . . . . . . . . . . . . . . . . . . 41 preserving with .PRECIOUS . . . . . . . . . . . . . . . . . . . . . . 115 print (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 print target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 40 printing directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 printing of commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 problems and bugs, reporting . . . . . . . . . . . . . . . . . . . . . . 10 problems with wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 processing a makefile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

Q

question mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 quoting %, in patsubst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 quoting %, in static pattern. . . . . . . . . . . . . . . . . . . . . . . . . 45 quoting %, in vpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 quoting newline, in commands . . . . . . . . . . . . . . . . . . . . . 52 quoting newline, in makefile. . . . . . . . . . . . . . . . . . . . . . . . 15

R

Ratfor, rule to compile. . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 RCS, rule to extract from . . . . . . . . . . . . . . . . . . . . . . . . . 111 README . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 realclean (standard target) . . . . . . . . . . . . . . . . . . . . . . . 98 recompilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 recompilation, avoiding . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 recording events with empty targets . . . . . . . . . . . . . . . . 40 recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 recursion, and -C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and -I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and -j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and -t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 recursion, and -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 recursion, and -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and environment . . . . . . . . . . . . . . . . . . . . . . . . 57 recursion, and MAKE variable. . . . . . . . . . . . . . . . . . . . . . . . 56 recursion, and MAKEFILES variable . . . . . . . . . . . . . . . . . . 26 recursion, and options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 recursion, and printing directories . . . . . . . . . . . . . . . . . . 61 recursion, and variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 recursion, level of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 recursive variable expansion . . . . . . . . . . . . . . . . . . . . 65, 66 recursively expanded variables . . . . . . . . . . . . . . . . . . . . . 66 reference to variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 69 relinking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 remaking makefiles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 removing duplicate words . . . . . . . . . . . . . . . . . . . . . . . . . . 88 removing, to clean up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 rm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 rm (shell command) . . . . . . . . . . . . . . . . . . . . . 16, 31, 38, 54 rule commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 rule dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 rule syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 rule targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 rule, and $. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 rule, double-colon (::) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 rule, explicit, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . 23 rule, how to write. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 rule, implicit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 rule, implicit, and directory search . . . . . . . . . . . . . . . . . 36 rule, implicit, and VPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 rule, implicit, chains of . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 rule, implicit, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 23 rule, implicit, how to use. . . . . . . . . . . . . . . . . . . . . . . . . . 107 rule, implicit, introduction to . . . . . . . . . . . . . . . . . . . . . . 18 rule, implicit, predefined . . . . . . . . . . . . . . . . . . . . . . . . . . 109 rule, introduction to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 rule, multiple for one target . . . . . . . . . . . . . . . . . . . . . . . . 43 rule, no commands or dependencies . . . . . . . . . . . . . . . . 39 rule, pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 rule, static pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 rule, static pattern versus implicit . . . . . . . . . . . . . . . . . . 46 rule, with multiple targets . . . . . . . . . . . . . . . . . . . . . . . . . 42

S

s. (SCCS file prefix). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 SCCS, rule to extract from . . . . . . . . . . . . . . . . . . . . . . . 111 search algorithm, implicit rule . . . . . . . . . . . . . . . . . . . . 125 search path for dependencies (VPATH). . . . . . . . . . . . . . . 33 search path for dependencies (VPATH), and implicit

rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 search path for dependencies (VPATH), and link libraries

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 searching for strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 sed (shell command). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 selecting words . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 sequences of commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 setting options from environment . . . . . . . . . . . . . . . . . . 61 setting options in makefiles . . . . . . . . . . . . . . . . . . . . . . . . 61 setting variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 several rules for one target . . . . . . . . . . . . . . . . . . . . . . . . . 43 several targets in a rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 shar (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 shell command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 shell command, and directory search . . . . . . . . . . . . . . . 36 shell command, execution . . . . . . . . . . . . . . . . . . . . . . . . . . 52 shell command, function for . . . . . . . . . . . . . . . . . . . . . . . . 95 shell file name pattern (in include) . . . . . . . . . . . . . . . . 24 shell wildcards (in include). . . . . . . . . . . . . . . . . . . . . . . . 24 signal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 silent operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 simple makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 simple variable expansion . . . . . . . . . . . . . . . . . . . . . . . . . . 65 simplifying with variables . . . . . . . . . . . . . . . . . . . . . . . . . . 17 simply expanded variables . . . . . . . . . . . . . . . . . . . . . . . . . 67 sorting words . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 spaces, in variable values . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 spaces, stripping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 special targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 specifying makefile name . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 standard input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 standards conformance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 standards for makefiles. . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 static pattern rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 static pattern rule, syntax of . . . . . . . . . . . . . . . . . . . . . . . 44 static pattern rule, versus implicit . . . . . . . . . . . . . . . . . . 46 stem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 120 stem, variable for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 strings, searching for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 stripping whitespace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 sub-make. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 subdirectories, recursion for . . . . . . . . . . . . . . . . . . . . . . . . 56 substitution variable reference. . . . . . . . . . . . . . . . . . . . . . 69 suffix rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 suffix rule, for archive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 suffix, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 suffix, function to find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 suffix, substituting in variables . . . . . . . . . . . . . . . . . . . . . 69 switches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 symbol directories, updating archive . . . . . . . . . . . . . . 128 syntax of rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

T

tab character (in commands). . . . . . . . . . . . . . . . . . . . . . . 29 tabs in rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 TAGS (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 tangle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 tar (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 target pattern, implicit . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 target pattern, static (not implicit). . . . . . . . . . . . . . . . . 44 target, deleting on interrupt. . . . . . . . . . . . . . . . . . . . . . . . 55 target, multiple in pattern rule. . . . . . . . . . . . . . . . . . . . 116 target, multiple rules for one . . . . . . . . . . . . . . . . . . . . . . . 43 target, touching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 targets without a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 targets, built-in special. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 targets, empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 targets, force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 targets, introduction to . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 targets, multiple . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 targets, phony . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 terminal rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 test (standard target) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 testing compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 TEX, rule to run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 texi2dvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 Texinfo, rule to format. . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 tilde (~) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 touch (shell command) . . . . . . . . . . . . . . . . . . . . . . . . . 31, 40 touching files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

U

updating archive symbol directories . . . . . . . . . . . . . . . 128 updating makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

V

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 value, how a variable gets it. . . . . . . . . . . . . . . . . . . . . . . . 73 variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 variable definition. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 variables, ‘$’ in name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 variables, and implicit rule . . . . . . . . . . . . . . . . . . . . . . . . 118 variables, appending to . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 variables, automatic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 variables, command line . . . . . . . . . . . . . . . . . . . . . . . . . . 101 variables, computed names . . . . . . . . . . . . . . . . . . . . . . . . . 70 variables, defining verbatim . . . . . . . . . . . . . . . . . . . . . . . . 77 variables, environment . . . . . . . . . . . . . . . . . . . . . . . . . 57, 78 variables, exporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 variables, flavors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 variables, how they get their values . . . . . . . . . . . . . . . . 73 variables, how to reference . . . . . . . . . . . . . . . . . . . . . . . . . 65 variables, loops in expansion . . . . . . . . . . . . . . . . . . . . . . . 67 variables, modified reference . . . . . . . . . . . . . . . . . . . . . . . 69 variables, nested references . . . . . . . . . . . . . . . . . . . . . . . . . 70 variables, origin of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 variables, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 variables, overriding with arguments . . . . . . . . . . . . . . 101 variables, recursively expanded . . . . . . . . . . . . . . . . . . . . . 66 variables, setting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 variables, simply expanded . . . . . . . . . . . . . . . . . . . . . . . . . 67 variables, spaces in values . . . . . . . . . . . . . . . . . . . . . . . . . . 68 variables, substituting suffix in . . . . . . . . . . . . . . . . . . . . . 69 variables, substitution reference . . . . . . . . . . . . . . . . . . . . 69 varying dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 verbatim variable definition . . . . . . . . . . . . . . . . . . . . . . . . 77 vpath. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 VPATH, and implicit rules . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 VPATH, and link libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

W

weave. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 113 Web, rule to run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 what if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 whitespace, stripping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 wildcard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 wildcard pitfalls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 wildcard, function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 wildcard, in include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 words, extracting first. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 words, filtering. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 words, filtering out. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 words, finding number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 words, iterating over . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 words, joining lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 words, removing duplicates. . . . . . . . . . . . . . . . . . . . . . . . . 88 words, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 writing rule commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 writing rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Y

yacc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 111, 113 Yacc, rule to run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

Index of Functions, Variables, & Directives 165

Index of Functions, Variables, & Directives

$

$% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 $(%D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(%F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(*D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(*F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(?D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 $(?F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 $(@D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(@F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(^D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(^F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 $(<D) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $(<F) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 $* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 $*, and static pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 $? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 $@ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 $^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 $< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

%

% (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 %D (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 %F (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

*

* (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 *D (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 *F (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

.

.DEFAULT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 123 .DEFAULT, and empty commands . . . . . . . . . . . . . . . . . . . 63 .EXPORT ALL VARIABLES . . . . . . . . . . . . . . . . . . . . . . . . 41, 59 .IGNORE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 54 .PHONY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38, 40 .PRECIOUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 55 .SILENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 51 .SUFFIXES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 124

/

/usr/gnu/include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 /usr/include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 /usr/local/include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

?

? (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 ?D (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 ?F (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

                                @

@ (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 @D (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 @F (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

                                ^

^(automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 ^D (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ^F (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

                                <

< (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 <D (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 <F (automatic variable) . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

A

addprefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 addsuffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 AR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 ARFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 AS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 ASFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

B

basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

C

CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 CFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 CO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 COFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 CPP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 CTANGLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 CWEAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 CXX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

D

define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

E

else. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 endef. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 endif. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

F

FC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 FFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 filter-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 findstring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 firstword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 foreach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

G

GET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 GFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 GNUmakefile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

I

ifdef. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 ifeq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 ifndef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 ifneq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

J

join. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

L

LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 LEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 LFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

M

MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 68 MAKE COMMAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 MAKEFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 60 MAKEFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 MAKEINFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 MAKELEVEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 68 MAKEOVERRIDES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 MFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

N

notdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

O

origin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 OUTPUT OPTION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 override . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

P

patsubst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 86 PC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 PFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

R

RFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 RM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

S

shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 SHELL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 SHELL (command execution) . . . . . . . . . . . . . . . . . . . . . . . 52 sort. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 strip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

Index of Functions, Variables, & Directives 167

subst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42, 86 suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 SUFFIXES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

T

TANGLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 TEXI2DVI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

U

unexport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

V

vpath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 34 VPATH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

W

WEAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 wildcard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 91 word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 words. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

Y

YACC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 YACCR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 YFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

168 GNU make

Short Contents

GNU GENERAL PUBLIC LICENSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Overview of make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2 An Introduction to Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3 Writing Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4 Writing Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5 Writing the Commands in Rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 6 How to Use Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7 Conditional Parts of Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 8 Functions for Transforming Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 9 How to Run make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 10 Using Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 11 Using make to Update Archive Files . . . . . . . . . . . . . . . . . . . . . . . . . 127 12 Features of GNU make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 13 Incompatibilities and Missing Features. . . . . . . . . . . . . . . . . . . . . . . . 135 14 Makefile Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Appendix A Quick Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 Appendix B Complex Makefile Example . . . . . . . . . . . . . . . . . . . . . . . . 151 Index of Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 Index of Functions, Variables, & Directives . . . . . . . . . . . . . . . . . . . . . . . . 165

ii GNU make

Table of Contents

GNU GENERAL PUBLIC LICENSE . . . . . . . . . . . . . . . . . . 1

Preamble. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND

MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 How to Apply These Terms to Your New Programs . . . . . . . . . . . . . . . . . . . . . 7

1 Overview of make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

1.1 How to Read This Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.2 Problems and Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2 An Introduction to Makefiles . . . . . . . . . . . . . . . . . . . . . . 13

2.1 What a Rule Looks Like . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.2 A Simple Makefile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.3 How make Processes a Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.4 Variables Make Makefiles Simpler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.5 Letting make Deduce the Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.6 Another Style of Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.7 Rules for Cleaning the Directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

3 Writing Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.1 What Makefiles Contain. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.2 What Name to Give Your Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.3 Including Other Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.4 The Variable MAKEFILES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.5 How Makefiles Are Remade . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6 Overriding Part of Another Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

4 Writing Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

4.1 Rule Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.2 Using Wildcard Characters in File Names . . . . . . . . . . . . . . . . . . . . . . . . 30

4.2.1 Wildcard Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.2.2 Pitfalls of Using Wildcards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.2.3 The Function wildcard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

4.3 Searching Directories for Dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.3.1 VPATH: Search Path for All Dependencies . . . . . . . . . . . . . . . . 33 4.3.2 The vpath Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.3.3 Writing Shell Commands with Directory Search . . . . . . . . . 36 4.3.4 Directory Search and Implicit Rules . . . . . . . . . . . . . . . . . . . . . 36 4.3.5 Directory Search for Link Libraries . . . . . . . . . . . . . . . . . . . . . . 37

4.4 Phony Targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.5 Rules without Commands or Dependencies . . . . . . . . . . . . . . . . . . . . . . . 39 4.6 Empty Target Files to Record Events. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.7 Special Built-in Target Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.8 Multiple Targets in a Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.9 Multiple Rules for One Target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.10 Static Pattern Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

4.10.1 Syntax of Static Pattern Rules . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.10.2 Static Pattern Rules versus Implicit Rules . . . . . . . . . . . . . . 46

4.11 Double-Colon Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.12 Generating Dependencies Automatically . . . . . . . . . . . . . . . . . . . . . . . . . 47

5 Writing the Commands in Rules. . . . . . . . . . . . . . . . . . . 51

5.1 Command Echoing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5.2 Command Execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 5.3 Parallel Execution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 5.4 Errors in Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.5 Interrupting or Killing make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.6 Recursive Use of make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

5.6.1 How the MAKE Variable Works . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.6.2 Communicating Variables to a Sub-make. . . . . . . . . . . . . . . . . 57 5.6.3 Communicating Options to a Sub-make . . . . . . . . . . . . . . . . . . 60 5.6.4 The ‘--print-directory’ Option . . . . . . . . . . . . . . . . . . . . . . . 61

5.7 Defining Canned Command Sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.8 Using Empty Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

6 How to Use Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.1 Basics of Variable References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.2 The Two Flavors of Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 6.3 Advanced Features for Reference to Variables . . . . . . . . . . . . . . . . . . . . . 69

6.3.1 Substitution References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.3.2 Computed Variable Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

6.4 How Variables Get Their Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 6.5 Setting Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 6.6 Appending More Text to Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 6.7 The override Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.8 Defining Variables Verbatim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 6.9 Variables from the Environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

7 Conditional Parts of Makefiles . . . . . . . . . . . . . . . . . . . . . 79

7.1 Example of a Conditional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.2 Syntax of Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 7.3 Conditionals that Test Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

8 Functions for Transforming Text. . . . . . . . . . . . . . . . . . . 85

8.1 Function Call Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 8.2 Functions for String Substitution and Analysis. . . . . . . . . . . . . . . . . . . . 86 8.3 Functions for File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 8.4 The foreach Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 8.5 The origin Function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 8.6 The shell Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

9 How to Run make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

9.1 Arguments to Specify the Makefile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 9.2 Arguments to Specify the Goals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 9.3 Instead of Executing the Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 9.4 Avoiding Recompilation of Some Files . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 9.5 Overriding Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 9.6 Testing the Compilation of a Program . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 9.7 Summary of Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

10 Using Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

10.1 Using Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 10.2 Catalogue of Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 10.3 Variables Used by Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 10.4 Chains of Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 10.5 Defining and Redefining Pattern Rules . . . . . . . . . . . . . . . . . . . . . . . . . 115

10.5.1 Introduction to Pattern Rules . . . . . . . . . . . . . . . . . . . . . . . . . 116 10.5.2 Pattern Rule Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 10.5.3 Automatic Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 10.5.4 How Patterns Match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 10.5.5 Match-Anything Pattern Rules. . . . . . . . . . . . . . . . . . . . . . . . 120 10.5.6 Canceling Implicit Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

10.6 Defining Last-Resort Default Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 10.7 Old-Fashioned Suffix Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 10.8 Implicit Rule Search Algorithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 vi GNU make

11 Using make to Update Archive Files . . . . . . . . . . . . . 127

11.1 Archive Members as Targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 11.2 Implicit Rule for Archive Member Targets . . . . . . . . . . . . . . . . . . . . . . 127

11.2.1 Updating Archive Symbol Directories . . . . . . . . . . . . . . . . . 128 11.3 Suffix Rules for Archive Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

12 Features of GNU make . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 13 Incompatibilities and Missing Features . . . . . . . . . 135 14 Makefile Conventions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

14.1 General Conventions for Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 14.2 Utilities in Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 14.3 Standard Targets for Users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 14.4 Variables for Specifying Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 14.5 Variables for Installation Directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

Appendix A Quick Reference . . . . . . . . . . . . . . . . . . . . . . . 145 Appendix B Complex Makefile Example . . . . . . . . . . . 151 Index of Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 Index of Functions, Variables, & Directives . . . . . . . . . 165